### v1

Abstract
For a relational structure ${\mathbb{X}}$ we investigate the partial order $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$, where ${\mathbb{P}}({\mathbb{X}}):=\{f[X]:f\in\mathop{\rm Emb}\nolimits({\mathbb{X}})\}$. A previous analysis shows that ${\mathfrak{h}}=\omega_{1}$ implies that for each countable ordinal we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin})$. But in ZFC we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))\cong\prod_{i=0}^{n}((\mathop{\rm rp}\nolimits^{r_{i}}(P(\omega^{\gamma_{i}})/{\mathcal{I}}_{\omega^{\gamma_{i}}}))^{+})^{s_{i}}$, where $\sum_{i=n}^{0}\omega^{\gamma_{i}+r_{i}}s_{i}+k$ is the presentation of $\alpha$ in the Cantor normal form and $\mathop{\rm rp}\nolimits({\mathbb{B}})$ denotes the reduced power of a Boolean algebra ${\mathbb{B}}$ modulo finite: ${\mathbb{B}}^{\omega}/\mathop{\rm Fin}$. Consequently, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin})^{+}\ast\pi)$, where $\pi$ is an $P(\omega)/\mathop{\rm Fin}$-name for an $\sigma$-closed separative atomless poset.
Here we consider uncountable ordinals. Since $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is isomorphic to the direct product $\prod_{i=1}^{n}(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta_{i}}))^{s_{i}}$, where $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is the Cantor normal form for $\alpha$, the analysis is reduced to the investigation of the posets of the form ${\mathbb{P}}(\omega^{\delta})$. It turns out that, in ZFC, either the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed and completely embeds $P(\omega)/\mathop{\rm Fin}$ and, hence, preserves $\omega_{1}$ and forces $|{\mathfrak{c}}|=|{\mathfrak{h}}|$, or, otherwise, completely embeds the algebra $P(\lambda)/[\lambda]^{<\lambda}$, for some regular $\omega<\lambda\leq\mathop{\rm cf}\nolimits(\delta)$, and collapses $\omega_{2}$ to $\omega$. Regarding the Cantor normal form, the first case appears iff for each $i\leq n$ we have $\mathop{\rm cf}\nolimits(\delta_{i})\leq\omega$, or $\delta_{i}=\theta_{i}+\mathop{\rm cf}\nolimits(\delta_{i})$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta_{i}\geq\mathop{\rm cf}\nolimits(\delta_{i})>\mathop{\rm cf}\nolimits(\theta_{i})=\omega$ and $\theta_{i}=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\mathop{\rm cf}\nolimits(\delta_{i})$, for all $n\in\omega$.
2020 MSC: 06A05, 06A10, 03E40, 03E35.
Keywords: uncountable ordinal, poset of copies, $\sigma$-closed poset, cardinal collapse, forcing.
FORCING WITH COPIES OF UNCOUNTABLE ORDINALS
Miloš S. Kurilić [^1] Department of Mathematics and Informatics, Faculty of Sciences, University of Novi Sad, Trg Dositeja Obradovića 4, 21000 Novi Sad, Serbia. e-mail: milos@dmi.uns.ac.rs
1 Introduction
In this paper we continue the investigation of posets of copies of ordinals initiated in [14], where countable ordinals were considered. Generally speaking, if ${\mathbb{X}}$ is a relational structure and $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is the set of its self-embeddings, $f:{\mathbb{X}}\hookrightarrow{\mathbb{X}}$, then the partial order $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$, where ${\mathbb{P}}({\mathbb{X}}):=\{f[X]:f\in\mathop{\rm Emb}\nolimits({\mathbb{X}})\}$, is called the poset of copies of ${\mathbb{X}}$. Denoting by $\mathop{\rm sq}\nolimits({\mathbb{P}})$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))$ the separative quotient and the Boolean completion of a partial order ${\mathbb{P}}$, the implications
${\mathbb{X}}\cong{\mathbb{Y}}\Rightarrow{\mathbb{P}}({\mathbb{X}})\cong{\mathbb{P}}({\mathbb{Y}})\Rightarrow\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})\Rightarrow\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})$
express the relationship between different “similarity relations” in the class of relational structures and each of them provides a classification of structures [12, 15]. The last one, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})$, gives the coarsest classification of structures, it is equivalent to the forcing equivalence of posets of copies, ${\mathbb{P}}({\mathbb{X}})\equiv_{forc}{\mathbb{P}}({\mathbb{Y}})$, and, hence, it can be explored using forcing-theoretic methods and known facts from set theory. The related results can be regarded as the statements belonging to the theory of relations, set theory or theory of Boolean algebras.
It turns out that, when the posets of copies of countable linear orders are in question, the main roles are played by the poset ${\mathbb{S}}$ of perfect subsets of the real line ordered by inclusion (the Sacks forcing) and the algebra $P(\omega)/\mathop{\rm Fin}\nolimits$. Namely, if ${\mathbb{X}}$ is a non-scattered linear order (i.e, if ${\mathbb{Q}}\hookrightarrow{\mathbb{X}}$), then $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}}\ast\pi)$, where $\pi$ is an ${\mathbb{S}}$-name for a $\sigma$-closed poset and, consistently, the name for $P(\omega)/\mathop{\rm Fin}\nolimits$ in the Sacks extension [17]. On the other hand, if ${\mathbb{X}}$ is a scattered linear order, then the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})$ is $\sigma$-closed, atomless and, consistently, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ [13]. Well orders were investigated in [14] and the main results will be presented later in this section. We note that for the real line ${\mathbb{R}}$ we have $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{R}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}})$ [16] and that the Sacks forcing appears as a factor in the presentation of posets of copies of several ultrahomogeneous structures, [8, 17, 18, 19, 20, 21].
The results obtained in this area can be applied in investigation of transformation monoids. If ${\mathbb{X}}$ is a relational structure and $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),\circ,\mathop{\rm id}\nolimits_{X}\rangle$ is the monoid of its self-embeddings, first we note that the set $\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})$ of automorphisms of ${\mathbb{X}}$ together with a choice of an embedding $f_{A}\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$ such that $f_{A}[X]=A$, for each $A\in{\mathbb{P}}({\mathbb{X}})$, completely determine the set of embeddings: $\mathop{\rm Emb}\nolimits({\mathbb{X}})=\{f_{A}\circ a:A\in{\mathbb{P}}({\mathbb{X}})\mbox{ and }a\in\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})\}$. Second, we recall that right Green’s preorder on $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is defined by $f\preccurlyeq^{R}g$ iff $f\circ h=g$, for some $h\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$, and note that, defining $f\thickapprox^{R}g$ iff $f\preccurlyeq^{R}g$ and $g\preccurlyeq^{R}f$, we have
$f\preccurlyeq^{R}g\Leftrightarrow g[X]\subset f[X]\;\;\mbox{ and }\;\;f\thickapprox^{R}g\Leftrightarrow g[X]=f[X].$ (1)
Thus the poset of copies $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$ is exactly the antisymmetric quotient of the preorder $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),(\preccurlyeq^{R})^{-1}\rangle$.
Now we turn to the particular case considered in this paper, when ${\mathbb{X}}$ is a well order, say an ordinal $\alpha$. Then it is a rigid structure ($\mathop{\mathrm{Aut}}\nolimits(\alpha)=\{\mathop{\rm id}\nolimits_{\alpha}\}$) and for each copy $A\in{\mathbb{P}}(\alpha)$ there is a unique $f_{A}\in\mathop{\rm Emb}\nolimits(\alpha)$ such that $f_{A}[\alpha]=A$. Thus we have a simple reconstruction:
$\mathop{\rm Emb}\nolimits(\alpha)=\{f_{A}:A\in{\mathbb{P}}(\alpha)\}.$
Moreover, defining the binary operation $\ast$ on ${\mathbb{P}}(\alpha)$ by $A\ast B=f_{A}[B]$ and the mapping $F:{\mathbb{P}}(\alpha)\rightarrow\mathop{\rm Emb}\nolimits(\alpha)$ by $F(A)=f_{A}$, we obtain an isomorphism of monoids:
$\langle{\mathbb{P}}(\alpha),\ast,\alpha\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\circ,\mathop{\rm id}\nolimits_{\alpha}\rangle.$ (2)
In addition, the preorder $\preccurlyeq^{R}$ on $\mathop{\rm Emb}\nolimits(\alpha)$ becomes a partial order and the same mapping $F$ witnesses that
$\langle{\mathbb{P}}(\alpha),\supset\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\preccurlyeq^{R}\rangle.$ (3)
For convenience we consider the poset $\langle{\mathbb{P}}(\alpha),\subset\rangle$, shortly denoted by ${\mathbb{P}}(\alpha)$.
In [14] the posets of copies of countable ordinals $\alpha\geq\omega$ were investigated and, in order to present the main results, we explain some notation. If ${\mathbb{B}}$ is Boolean lattice, $\mathop{\rm rp}\nolimits({\mathbb{B}})$ denotes the reduced power ${\mathbb{B}}^{\omega}/\mathop{\rm Fin}\nolimits$, and the iterated reduced powers $\mathop{\rm rp}\nolimits^{n}({\mathbb{B}})$, for $n\in\omega$, are defined by $\mathop{\rm rp}\nolimits^{0}({\mathbb{B}})={\mathbb{B}}$ and $\mathop{\rm rp}\nolimits^{n+1}({\mathbb{B}})=\mathop{\rm rp}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))$. If $\delta>0$ is an ordinal, then ${\mathcal{I}}_{\omega^{\delta}}:=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}\subset P(\omega^{\delta})$ is an ideal and the Boolean algebra $P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$ is the corresponding quotient. The first result shows that the separative quotient of the poset ${\mathbb{P}}(\alpha)$ is isomorphic to the forcing (free) product of finitely many Boolean algebras.
Theorem 1.1
If $\alpha$ is a countable ordinal presented in the Cantor normal form,
$\alpha=\omega^{\gamma_{n}+r_{n}}s_{n}+\dots+\omega^{\gamma_{0}+r_{0}}s_{0}+k,$ (4)
where $k\in\omega$, $r_{i}\in\omega$, $s_{i}\in{\mathbb{N}}$, $\gamma_{i}\in\mathop{\rm Lim}\nolimits\cup\{1\}$ and $\gamma_{n}+r_{n}>\dots>\gamma_{0}+r_{0}$, then
$\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))\cong\prod_{i=0}^{n}\Big{(}\Big{(}\mathop{\rm rp}\nolimits^{r_{i}}(P(\omega^{\gamma_{i}})/{\mathcal{I}}_{\omega^{\gamma_{i}}})\Big{)}^{+}\Big{)}^{s_{i}}.$ (5)
So, inductively, for $\alpha=\omega$ from (5) we obtain $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$; further we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega s)=((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})^{s}$, for $s\in{\mathbb{N}}$; $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega\omega)=(\mathop{\rm rp}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits))^{+}$; and, generally, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{r}))=(\mathop{\rm rp}\nolimits^{r-1}(P(\omega)/\mathop{\rm Fin}\nolimits))^{+}$, for $r\in{\mathbb{N}}$. Further for a limit $\gamma$ and $r\in\omega$ we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\gamma+r}))=(\mathop{\rm rp}\nolimits^{r}(P(\omega^{\gamma})/{\mathcal{I}}_{\omega^{\gamma}}))^{+}$; and, finally, by (5) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is the free product of the posets $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha_{i}))$, where $\alpha_{i}=\omega^{\gamma_{i}+r_{i}}$, for $i\leq n$, are the infinite summands from (4).
The second result is related to the aforementioned coarse classification of structures (corresponding to the similarity $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\beta)))$) and, by (3), a classification of their embedding monoids. We recall that the distributivity number ${\mathfrak{h}}({\mathbb{P}})$ of a partial order ${\mathbb{P}}$ is the minimal size of a family ${\mathcal{D}}$ of open dense subsets of ${\mathbb{P}}$ such that $\bigcap{\mathcal{D}}$ is not dense, and that $\omega_{1}\leq{\mathfrak{h}}:={\mathfrak{h}}((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})\leq{\mathfrak{c}}$.
Theorem 1.2
If $\alpha\geq\omega+\omega$ is a countable ordinal, then we have
  • (a)
    ${\mathbb{P}}(\alpha)\equiv_{forc}(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi$, where $[\omega]\Vdash$$\pi$ is an $\sigma$-closed separative atomless poset”; and, consequently,
    $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi);$ (6)
  • (b)
    If ${\mathfrak{h}}=\omega_{1}$, then ${\mathbb{P}}(\alpha)\equiv_{forc}(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$ and, hence,
    $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+}).$ (7)
[…]
The following example shows that the statement that (7) holds for all countable ordinals is independent of ZFC. So the coarse classification depends on the model of ZFC in which we live. Defining ${\mathfrak{h}}_{n}:={\mathfrak{h}}(((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})^{n})$, for $n\in{\mathbb{N}}$, we have ${\mathfrak{h}}_{1}\geq{\mathfrak{h}}_{2}\geq{\mathfrak{h}}_{3}\geq\dots\geq\omega_{1}$ and, by Theorem 1.1, ${\mathfrak{h}}(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega n))={\mathfrak{h}}_{n}$. Shelah and Spinas proved in [24] that for each $n\in{\mathbb{N}}$ there is a model of ZFC in which ${\mathfrak{h}}_{n+1}<{\mathfrak{h}}_{n}$, which implies that ${\mathbb{P}}(\omega n)\not\equiv_{forc}{\mathbb{P}}(\omega(n+1))$ and, hence, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega n)))\not\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega(n+1))))$.
In this paper we investigate the posets ${\mathbb{P}}(\alpha)$ for uncountable ordinals $\alpha$. Section 2 contains definitions and known facts used in the paper.
In Section 3 we show that the separative quotient $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is isomorphic to the direct product of $\sum_{i=1}^{n}s_{i}$-many factors $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta_{i}})$, where $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is the Cantor normal form for $\alpha$, and reduce the analysis to the investigation of these factors.
So, in Section 4 we regard the posets of the form ${\mathbb{P}}(\omega^{\delta})$, for ordinals $\delta>0$. It turns out that the behavior of ${\mathbb{P}}(\omega^{\delta})$ depends on the cofinality of $\delta$. If $\mathop{\rm cf}\nolimits(\delta)\in\{1,\omega\}$ (Case (A)), then the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed, completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$ and, hence, collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$. Otherwise we have $\mathop{\rm cf}\nolimits(\delta):=\kappa\geq\omega_{1}$ and, presenting $\delta$ in its Cantor normal form in the base $\kappa$, we detect additional four cases, (B)–(E). If (B) holds, then again, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. In Cases (D) and (E) the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds ${\mathcal{P}}_{\kappa}^{+}$, where ${\mathcal{P}}_{\kappa}$ denotes the algebra $P(\kappa)/[\kappa]^{<\kappa}$, and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ (and, hence, $\omega_{2}$) to $\omega$. Finally, in Case (C) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds ${\mathcal{P}}_{\lambda}^{+}$, where $\omega_{1}<\lambda<\kappa$, and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})$ (and, again, $\omega_{2}$) to $\omega$.
Thus, by the (ZFC) results of Section 4, the poset ${\mathbb{P}}(\omega^{\delta})$ either preserves $\omega_{1}$ and forces $|{\mathfrak{c}}|=|{\mathfrak{h}}|$, or collapses $\omega_{2}$ to $\omega$. So, it is natural to ask whether ${\mathbb{P}}(\omega^{\delta})$ collapses more cardinals (which, of course, depends on $|\delta|$, $\mathop{\rm cc}\nolimits({\mathbb{P}}(\omega^{\delta}))$, etc.) and a more ambitious task is to determine $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))$ (using, for example, some known characterizations of collapsing algebras). Here we note that, generally, for a poset ${\mathbb{P}}$ we have
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\cong\mathop{\rm Col}\nolimits(\lambda,\kappa)\Rightarrow\mathop{\rm Col}\nolimits(\lambda,\kappa)\hookrightarrow_{c}\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\Rightarrow 1_{\mathbb{P}}\Vdash|\check{\kappa}|=|\check{\lambda}|$
and by Kripke’s theorem (see [10], p. 227) for a Boolean algebra ${\mathbb{B}}$ of size $\leq\kappa$ we have $\mathop{\rm ro}\nolimits({\mathbb{B}}^{+}\times\mathop{\rm Col}\nolimits(\omega,\kappa)^{+})\cong\mathop{\rm Col}\nolimits(\omega,\kappa)$; thus the first iteration is “absorbed” by $\mathop{\rm Col}\nolimits(\omega,\kappa)$. Since $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|=|P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}|\leq 2^{|\delta|}$, the extreme case is when ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega_{1}$, in cases (A) and (B), and, in other cases, collapses $2^{|\delta|}$ to $\omega$. A more ambitious desire is to have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits(P(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{k},2^{|\delta|})$, for $k=1$ in the first case and for $k=0$ in the second. So, in Section 5 we consider that phenomenon of “maximal collapse”. For example, concerning Case (A), we show that if ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$, then ${\mathfrak{h}}(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))=\omega_{1}$ (in ZFC) and that, under some additional conditions (for example, if ${\mathfrak{h}}=\omega_{1}$ or if $0^{\sharp}$ does not exist and ${\mathfrak{b}}={\mathfrak{d}}$, e.g. if MA is true), we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits(P(\omega^{\delta+1})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. Regarding Cases (D) and (E) we show that $|\delta|=\mathop{\rm cf}\nolimits(\delta)=\kappa>\omega$ and $2^{\kappa}=\kappa^{+}$ implies that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$. A simple example shows that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$ is consistent. In Section 6 on the basis of the results from Sections 4 and 5 we analyze the general case, when $\alpha\geq\omega_{1}$.
2 Preliminaries
Partial orders and forcing
If ${\mathbb{P}}=\langle P,\leq\rangle$ is a preorder, the elements $p$ and $q$ of $P$ are incompatible, we write $p\perp q$, iff there is no $r\in P$ such that $r\leq p,q$. A set $A\subset P$ is an antichain iff $p\perp q$, for different $p,q\in A$. An element $p$ of $P$ is an atom iff each two elements $q,r\leq p$ are compatible. ${\mathbb{P}}$ is called: atomless iff it has no atoms; homogeneous iff it has a largest element and ${\mathbb{P}}\cong p\!\downarrow$, for each $p\in{\mathbb{P}}$, where $p\!\downarrow=\{q\in{\mathbb{P}}:q\leq p\}$. A set $D\subset P$ is called dense iff for each $p\in P$ there is $q\in D$ such that $q\leq p$; $D$ is called open iff $q\leq p\in D$ implies $q\in D$. A set $\Phi\subset P$ is a filter in ${\mathbb{P}}$ iff $\Phi\ni p\leq q$ implies $q\in\Phi$ and for each $p,q\in\Phi$ there is $r\in\Phi$ such that $r\leq p,q$. If $\kappa$ is a cardinal, ${\mathbb{P}}$ is called $\kappa$-closed iff for each $\gamma<\kappa$ each sequence $\langle p_{\alpha}:\alpha<\gamma\rangle$ in $P$, such that $\alpha<\beta\Rightarrow p_{\beta}\leq p_{\alpha}$, has a lower bound in $P$. $\omega_{1}$-closed preorders are called $\sigma$-closed. By ${\mathfrak{h}}({\mathbb{P}})$ we denote the distributivity number of a poset ${\mathbb{P}}$.
Fact 2.1
If ${\mathbb{P}}_{i}$, $i\in I$, are $\kappa$-closed preorders, then $\prod_{i\in I}{\mathbb{P}}_{i}$ is $\kappa$-closed.
A partial order ${\mathbb{P}}=\langle P,\leq\rangle$ is called separative iff for each $p,q\in P$ satisfying $p\not\leq q$ there is $r\leq p$ such that $r\perp q$. The separative modification of ${\mathbb{P}}$ is the separative preorder $\mathop{\rm sm}\nolimits({\mathbb{P}})=\langle P,\leq^{*}\rangle$, where
$p\leq^{*}q\Leftrightarrow\forall r\leq p\;\exists s\leq r\;s\leq q.$ (8)
The separative quotient of ${\mathbb{P}}$ is the separative partial order $\mathop{\rm sq}\nolimits({\mathbb{P}})=\langle P/\!\!=^{*},\trianglelefteq\rangle$, where $p=^{*}q\Leftrightarrow p\leq^{*}q\land q\leq^{*}p\;$ and $\;[p]\trianglelefteq[q]\Leftrightarrow p\leq^{*}q$. Two preorders ${\mathbb{P}}$ and ${\mathbb{Q}}$ are called forcing equivalent iff they produce the same generic extensions.
Fact 2.2
Let ${\mathbb{P}},{\mathbb{Q}}$ and ${\mathbb{P}}_{i}$, $i\in I$, be partial orderings. Then
(a) ${\mathbb{P}}$, $\mathop{\rm sm}\nolimits({\mathbb{P}})$ and $\mathop{\rm sq}\nolimits({\mathbb{P}})$ are forcing equivalent forcing notions;
(b) $\mathop{\rm sm}\nolimits({\mathbb{P}})$ is $\kappa$-closed iff $\mathop{\rm sq}\nolimits({\mathbb{P}})$ is $\kappa$-closed;
(c) If $p_{0},p_{1},\dots p_{n}\in{\mathbb{P}}$, where $p_{n}\leq^{*}p_{n-1}\leq^{*}\dots\leq^{*}p_{0}$, then there is $q\in{\mathbb{P}}$ such that $q\leq p_{k}$, for all $k\leq n$.
(d) ${\mathbb{P}}\cong{\mathbb{Q}}$ implies that $\mathop{\rm sm}\nolimits{\mathbb{P}}\cong\mathop{\rm sm}\nolimits{\mathbb{Q}}$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}\cong\mathop{\rm sq}\nolimits{\mathbb{Q}}$;
(e) $\mathop{\rm sm}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})=\prod_{i\in I}\mathop{\rm sm}\nolimits{\mathbb{P}}_{i}$ and $\mathop{\rm sq}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})\cong\prod_{i\in I}\mathop{\rm sq}\nolimits{\mathbb{P}}_{i}$.
(f) If $X$ is an infinite set, ${\mathcal{I}}\subset P(X)$ an ideal containing $[X]^{<\omega}$ and ${\mathcal{I}}^{+}=P(X)\setminus{\mathcal{I}}$ the corresponding family of ${\mathcal{I}}$-positive sets, then sm+,normal-$\mathop{\rm sm}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle$ $=\langle{\mathcal{I}}^{+},\subset_{\mathcal{I}}\rangle$, where $A\subset_{\mathcal{I}}B\Leftrightarrow A\setminus B\in{\mathcal{I}}$, for $A,B\in{\mathcal{I}}^{+}$. Also $\mathop{\rm sq}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle=(P(X)/{\mathcal{I}})^{+}$.
(g) $A\not\perp_{\mathbb{P}}B$ iff $A\cap B\in{\mathcal{I}}^{+}$.
Proof. All the statements are folklore except, maybe, (c). For a proof of (c), by recursion we define the sequence $\langle q_{k}:k\leq m\rangle$ such that (i) $q_{0}=p_{n}$ and (ii) $q_{k}\leq q_{k-1},p_{n-k}$, for $0<k\leq n$. Then $q_{n}\leq p_{k}$, for all $k\leq n$. $\Box$
We recall that, if $\langle{\mathbb{P}}\leq_{\mathbb{P}},1_{\mathbb{P}}\rangle$ and $\langle{\mathbb{Q}}\leq_{\mathbb{Q}},1_{\mathbb{Q}}\rangle$ are preorders, then a mapping $f:{\mathbb{P}}\rightarrow{\mathbb{Q}}$ is a complete embedding, in notation $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$ iff
(ce1) $p_{1}\leq_{\mathbb{P}}p_{2}\Rightarrow f(p_{1})\leq_{\mathbb{Q}}f(p_{2})$,
(ce2) $p_{1}\perp_{\mathbb{P}}p_{2}\Leftrightarrow f(p_{1})\perp_{\mathbb{Q}}f(p_{2})$,
(ce3) $\forall q\in{\mathbb{Q}}\;\exists p\in{\mathbb{P}}\;\forall p^{\prime}\leq_{\mathbb{P}}p\;\;f(p^{\prime})\not\perp_{\mathbb{Q}}q$.
Then, for $q\in{\mathbb{Q}}$, $\mathop{\rm red}\nolimits_{\mathbb{P}}(q)=\{p\in{\mathbb{P}}:\forall p^{\prime}\leq_{\mathbb{P}}p\;\;\exists q^{\prime}\in{\mathbb{Q}}\;\;q^{\prime}\leq f(p^{\prime}),q\}$ is the set of reductions of $q$ to ${\mathbb{P}}$. The following fact is folklore (see [11]).
Fact 2.3
If $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$, then ${\mathbb{Q}}$ is forcing equivalent to the two step iteration ${\mathbb{P}}\ast\langle\pi,\leq_{\pi},\check{1_{\mathbb{Q}}}\rangle$, where $1_{\mathbb{P}}\Vdash_{\mathbb{P}}\pi\subset\check{{\mathbb{Q}}}$ and for each $p\in{\mathbb{P}}$ and $q,q_{1},q_{2}\in{\mathbb{Q}}$
(a) $p\Vdash\check{q}\in\pi$ iff $p\in\mathop{\rm red}\nolimits(q)$;
(b) $p\Vdash\check{q_{1}}\leq_{\pi}\check{q_{2}}$ iff $q_{1}\leq_{\mathbb{Q}}q_{2}$ and $p\in\mathop{\rm red}\nolimits(q_{1})$.
Fact 2.4
If $V\models\mathrm{GCH}$ and, in $V$, $\kappa>\omega$ is a regular cardinal and ${\mathbb{P}}=\mathop{\mathrm{Fn}}\nolimits(\kappa,2)$, then in $V[G]$ we have ${\mathfrak{c}}=\kappa$. If, in $V$, $\mu$ and $\theta$ are infinite cardinals, then (see [11], p. 246)
$(\theta^{\mu})^{V[G]}=(\max(\kappa,\theta)^{\mu})^{V}$ (9)
Collapsing algebras and the algebras $P(\kappa)/[\kappa]^{<\kappa}$
Let $\kappa\geq 2$ and $\lambda\geq\omega$ be cardinals. The collapsing algebra $\mathop{\rm Col}\nolimits(\lambda,\kappa)$ is the Boolean completion $\mathop{\rm ro}\nolimits({}^{<\lambda}\kappa)$ of the reversed tree $\langle{}^{<\lambda}\kappa,\supset\rangle$. A Boolean algebra ${\mathbb{B}}$ is $(\lambda,\cdot,\kappa)$-nowhere distributive iff there are maximal antichains $A_{\alpha}$, $\alpha<\lambda$, such that each $b\in B^{+}$ is compatible with $\geq\kappa$ elements of some $A_{\alpha}$. (Then $\mathop{\rm cc}\nolimits({\mathbb{B}})>\kappa$.) Clearly, the algebra $\mathop{\rm Col}\nolimits(\lambda,\kappa)$ is $(\lambda,\cdot,\kappa)$-nowhere distributive and we have the following characterization.
Fact 2.5 (Balcar, Simon, [3], p. 338)
Let ${\mathbb{B}}$ be a homogeneous Boolean algebra, $\lambda\geq\omega$ a regular cardinal and $\kappa\geq 2$ a cardinal. If ${\mathbb{B}}$ is $(\lambda,\cdot,\kappa)$-nowhere distributive and contains a $\lambda$-closed dense subset, then $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,\kappa)\Leftrightarrow\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,\kappa^{<\lambda})\Leftrightarrow\pi({\mathbb{B}})=\kappa^{<\lambda}$.
Fact 2.6
(Folklore) Let ${\mathbb{P}}$ be a separative atomless preorder of size $\kappa$ and $\lambda$ a regular cardinal. If ${\mathbb{P}}$ is $\lambda$-closed, $1_{\mathbb{P}}\Vdash|\check{\kappa}|=\check{\lambda}$ and $\kappa^{<\lambda}=\kappa$, then ${\mathbb{P}}\equiv_{forc}\mathop{\rm Col}\nolimits(\lambda,\kappa)$. In particular, if $1_{\mathbb{P}}\Vdash|\check{\kappa}|=\check{\omega}$, then ${\mathbb{P}}\equiv_{forc}\mathop{\rm Col}\nolimits(\omega,\kappa)$.
Let $\kappa\geq\omega$ be a cardinal and $\subset^{*}$ the preorder on the set $[\kappa]^{\kappa}$ defined by $A\subset^{*}B$ iff $|A\setminus B|<\kappa$. By ${\mathcal{P}}_{\kappa}$ we denote the quotient algebra $P(\kappa)/[\kappa]^{<\kappa}$.
Fact 2.7
If $\kappa$ is an infinite cardinal, then
(a) The algebra ${\mathcal{P}}_{\kappa}$ is atomless, homogeneous, of size $2^{\kappa}$ and $\kappa^{++}\leq\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})\leq(2^{\kappa})^{+}$ (see [3], p. 372);
(b) (Balcar, Simon, [2]) If $\kappa>\mathop{\rm cf}\nolimits(\kappa)>\omega$, then the algebra ${\mathcal{P}}_{\kappa}$ is $(\omega,\cdot,\kappa^{+})$-nowhere distributive.
Fact 2.8
If $\kappa$ is an infinite cardinal, then
(a) $\mathop{\rm sm}\nolimits\langle[\kappa]^{\kappa},\subset\rangle=\langle[\kappa]^{\kappa},\subset^{*}\rangle$ and $\mathop{\rm sq}\nolimits\langle[\kappa]^{\kappa},\subset\rangle={\mathcal{P}}_{\kappa}^{+}$;
(b) (Shelah, [23]) If $\kappa>2^{\mathop{\rm cf}\nolimits(\kappa)}>\mathop{\rm cf}\nolimits(\kappa)>\omega$, then $\mathop{\rm Col}\nolimits(\omega,\kappa^{+})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$.
(c) (Shelah, [23]) If $\kappa=\mathop{\rm cf}\nolimits(\kappa)>\omega$, then forcing by ${\mathcal{P}}_{\kappa}^{+}$ collapses each $\lambda<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ to $\omega$;
(d) (Baumgartner, [5]) It is consistent with ZFC that $2^{\omega_{1}}$ is arbitrarily large and that $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}$;
(e) (Balcar, Pelant and Simon, [1]) Forcing by ${\mathcal{P}}_{\omega}^{+}$ collapses $2^{\omega}$ to ${\mathfrak{h}}$;
(f) (Kojman, Shelah [9]) If $\kappa>\mathop{\rm cf}\nolimits(\kappa)=\omega$, then $\mathop{\rm Col}\nolimits(\omega_{1},\kappa^{\omega})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$ (see also [4]).
Fact 2.9
Let $\kappa\in\mathop{\mathrm{Card}}\nolimits$, $2^{\kappa}=\kappa^{+}$ and $2^{\mathop{\rm cf}\nolimits(\kappa)}=\mathop{\rm cf}\nolimits(\kappa)^{+}$. Then (see [3], p. 380)
(a) If $\mathop{\rm cf}\nolimits(\kappa)>\omega$, then $\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})\cong\mathop{\rm Col}\nolimits(\omega\,,2^{\kappa})$;
(b) If $\mathop{\rm cf}\nolimits(\kappa)=\omega$, then $\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\kappa})$.
Linear orders. Indecomposable ordinals and ordinal ideals
A linear order $\langle L,<\rangle$ is called dense iff for each $x,y\in L$ satisfying $x<y$ there is $z\in L$ such that $x<z<y$. $\langle L,<\rangle$ is said to be scattered iff it does not contain a dense suborder iff ↪̸${\mathbb{Q}}\not\hookrightarrow L$.
Fact 2.10
A linear order $L$ satisfying $L+L\hookrightarrow L$ is not scattered ([22], p. 176).
Fact 2.11
For a limit ordinal $\alpha$ the following conditions are equivalent:
(a) $\beta+\gamma<\alpha$, for each $\beta,\gamma<\alpha$, (that is, $\alpha$ is indecomposable),
(b) $\beta+\gamma=\alpha\land\gamma>0\Rightarrow\gamma=\alpha$ (that is, $\alpha$ is right indecomposable),
(c) $\,\beta+\alpha=\alpha$, for each $\beta<\alpha$,
(d) $\alpha=\omega^{\delta}$, for some ordinal $\delta>0$,
(e) $\alpha$ is an indivisible structure ($\alpha\hookrightarrow A$ or $\alpha\hookrightarrow B$, whenever $\alpha=A\;\dot{\cup}\;B$),
(f) ${\mathcal{I}}_{\alpha}=\{I\subset\alpha:\alpha\not\hookrightarrow I\}$ is an ideal in $P(\alpha)$.
Proof. For (a) $\Leftrightarrow$ (d) $\Leftrightarrow$ (c) see [25], p. 282, 323. For (b) $\Leftrightarrow$ (d) see [22], p. 176. The equivalence (a) $\Leftrightarrow$ (e) is 6.8.1 of [7]. (e) $\Leftrightarrow$ (f) is evident. $\Box$
Fact 2.12
(a) $\omega^{\delta}\geq\delta$, for each ordinal $\delta$;
(b) $|\omega^{\delta}|=\omega|\delta|$, for each $\delta\geq 1$;
(c) If $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is an uncountable ordinal, then $|\alpha|=|\delta_{n}|$.
Proof. (a) First we have $\omega^{0}=1\geq 0$. Suppose that $\omega^{\delta}\geq\delta$, for each ordinal $\delta<\theta$. If $\theta=\delta+1$, then $\omega^{\theta}=\omega^{\delta}\omega>\omega^{\delta}1=\omega^{\delta}\geq\delta$ and, hence, $\omega^{\theta}\geq\delta+1=\theta$. If $\theta$ is a limit ordinal, then $\omega^{\theta}=\sup\{\omega^{\delta}:\delta<\theta\}\geq\sup\{\delta:\delta<\theta\}=\theta$.
(b) If $\delta\geq 1$, then $|\omega^{\delta}|\geq\omega$ and, by (a), $|\omega^{\delta}|\geq|\delta|$, so $|\omega^{\delta}|\geq\omega|\delta|$. Let
$\forall\delta<\theta\;\;|\omega^{\delta}|=\omega|\delta|.$ (10)
If $\theta=\delta+1$, then $|\omega^{\theta}|=|\omega^{\delta}\omega|=|\omega^{\delta}|\omega=|\delta|\omega\omega=|\delta+1|\omega=|\theta|\omega$.
If $\theta$ is a limit ordinal, then $\omega^{\theta}=\sup\{\omega^{\delta}:\delta<\theta\}$ and by (10), for each $\delta<\theta$ we have $|\omega^{\delta}|=\omega|\delta|\leq\omega|\theta|=|\theta|$. So the ordinal $\omega^{\theta}$ is an union of $|\theta|$-many ordinals of size $\leq|\theta|$ and, hence, $|\omega^{\theta}|=|\theta|=\omega|\theta|$.
(c) $\alpha$ is a finite sum of summands of the form $\omega^{\delta_{i}}$ and by (b) we have $|\omega^{\delta_{i}}|=\omega|\delta_{i}|$. So $\alpha\geq\omega_{1}$ implies that the largest summand, $\omega^{\delta_{n}}$ is uncountable and, hence, $|\delta_{n}|\geq\omega_{1}$. Since $i<n$ implies $\delta_{i}<\delta_{n}$ we have $\omega^{\delta_{i}}<\omega^{\delta_{n}}$, so $|\omega^{\delta_{i}}|\leq|\omega^{\delta_{n}}|$. Thus by (b) we have $|\alpha|=|\omega^{\delta_{n}}|=\omega|\delta_{n}|=|\delta_{n}|$. $\Box$
3 Factorization of ${\mathbb{P}}(\alpha)$
Here we start with analysis of the posets of the form ${\mathbb{P}}(\alpha)$, where $\alpha$ is an uncountable ordinal. Preliminarily we note that these posets are homogeneous (clearly, ${\mathbb{P}}({\mathbb{X}})\upharpoonright A\cong{\mathbb{P}}({\mathbb{X}})$, for each $A\in{\mathbb{P}}({\mathbb{X}})$, whenever ${\mathbb{X}}$ is a relational structure) and atomless. Although the last claim follows from Theorems 3.2 and 4.10 we sketch a simple proof. By the next lemma we can consider the case when $\alpha$ is a limit ordinal. Then each $\xi<\alpha$ is of the form $\xi=\gamma+n$, where $\gamma\in\mathop{\rm Lim}\nolimits\cup\{0\}$ and $n\in\omega$, and the sets $E$ and $O$ of even and odd elements of $\alpha$ are disjoint copies of $\alpha$ (the corresponding embeddings are given by $f(\gamma+n)=\gamma+2n$ and $g(\gamma+n)=\gamma+2n+1$).
Lemma 3.1
$\langle{\mathbb{P}}(\gamma+m),\subset\rangle\cong\langle{\mathbb{P}}(\gamma),\subset\rangle$, for each limit ordinal $\gamma$ and each $m\in{\mathbb{N}}$.
Proof. First we prove ${\mathbb{P}}(\gamma+m)=\{C\cup\{\gamma,\gamma+1,\dots,\gamma+m-1\}:C\in{\mathbb{P}}(\gamma)\}$. The inclusion $\supset$” is evident. If $A\in{\mathbb{P}}(\gamma+m)$ and $f:\gamma+m\hookrightarrow\gamma+m$, where $A=f[\gamma+m]$, then, since $f$ is an increasing function, we have $f(\beta)\geq\beta$, for each $\beta\in\gamma+m$, which implies $f(\gamma+i)=\gamma+i$, for $i<m$, and, hence, $C=f[\gamma]\in{\mathbb{P}}(\gamma)$ and $A=C\cup\{\gamma,\gamma+1,\dots,\gamma+m-1\}$.
Now it is easy to show that the mapping $F:\langle{\mathbb{P}}(\gamma),\subset\rangle\rightarrow\langle{\mathbb{P}}(\gamma+m),\subset\rangle$, given by $F(C)=C\cup\{\gamma,\gamma+1,\dots,\gamma+m-1\}$, is an isomorphism. $\Box$
Theorem 3.2
[…]
Proof. Since $\delta_{1},s_{1}>0$, by Lemma 3.1 we have ${\mathbb{P}}(\alpha)\cong{\mathbb{P}}(\gamma)$, where, defining $k:=\sum_{i=1}^{n}s_{i}$
$\gamma=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}=\sum_{j=1}^{k}{\mathbb{L}}_{j},$ (11)
and by induction on $k\in{\mathbb{N}}$ we prove that ${\mathbb{P}}(\gamma)\cong\prod_{i=1}^{n}\big{(}{\mathbb{P}}(\omega^{\delta_{i}})\big{)}^{s_{i}}(\cong\prod_{j=1}^{k}{\mathbb{P}}({\mathbb{L}}_{j})$), which will follow from the equality
${\mathbb{P}}(\gamma)=\big{\{}\bigcup_{j=1}^{k}C_{j}:\langle C_{1},\dots,C_{k}\rangle\in\prod_{j=1}^{k}{\mathbb{P}}({\mathbb{L}}_{j})\big{\}}.$ (12)
For $k=1$ we have $n=s_{1}=1$, so $\gamma=\omega^{\delta_{1}}={\mathbb{L}}_{1}$, ${\mathbb{P}}(\gamma)={\mathbb{P}}({\mathbb{L}}_{1})$ and (12) is true.
[…]
Let $C\in{\mathbb{P}}(\gamma^{\prime})$ and $f\in\mathop{\rm Emb}\nolimits(\gamma^{\prime})$, where $f[\gamma^{\prime}]=C$. We show that $C_{1}:=f[L_{1}]$ is a cofinal subset of ${\mathbb{L}}_{1}$; then we will have $f[\gamma]\subset\gamma$ and, hence, $f[\gamma]\in{\mathbb{P}}(\gamma)$; so, by the induction hypothesis (12) will be true for $k+1$.
Assuming that $F:=f[L_{1}]\cap\gamma\neq\emptyset$ the order $F$ would be isomorphic to a final part of ${\mathbb{L}}_{1}\cong\omega^{\delta_{n}}$ and, by Fact 2.11(b) we would have $F\cong\omega^{\delta_{n}}$. So we would have $\gamma^{\prime}\cong F+f[\gamma]\subset\gamma\subset\gamma^{\prime}$ and, hence $\gamma^{\prime}\cong\gamma=\omega^{\delta_{n}}(s_{n}-1)+\dots+\omega^{\delta_{1}}s_{1}$, which is impossible since each ordinal has a unique representation in the Cantor normal form. Thus $f[L_{1}]\subset L_{1}$ and $f[L_{1}]\cong L_{1}\cong\omega^{\delta_{n}}$, which implies that $C_{1}:=f[L_{1}]$ is a cofinal subset of ${\mathbb{L}}_{1}$ and $C_{1}\in{\mathbb{P}}({\mathbb{L}}_{1})$. Consequently, $f[\gamma]\subset\gamma$ and, hence, $f[\gamma]\in{\mathbb{P}}(\gamma)$; so, by the induction hypothesis $f[\gamma]=\bigcup_{j=2}^{k+1}C_{j}$, where $\langle C_{2},\dots,C_{k}\rangle\in\prod_{j=2}^{k+1}{\mathbb{P}}({\mathbb{L}}_{j})$ and $f[\gamma^{\prime}]=\bigcup_{j=1}^{k+1}C_{j}$.
Thus, ${\mathbb{P}}(\alpha)\cong\prod_{i=1}^{n}({\mathbb{P}}(\omega^{\delta_{i}}))^{s_{i}}$ and the rest follows from Fact 2.2(e). $\Box$
4 Factorization of factors ${\mathbb{P}}(\omega^{\delta})$
Classification of exponents $\delta$
If $\delta$ is an ordinal, $\mathop{\rm cf}\nolimits(\delta)=\kappa\geq\omega$, $\langle\delta_{\xi}:\xi\in\kappa\rangle\in{}^{\kappa}\delta$ is a strictly increasing sequence of ordinals and $\delta=\sup\{\delta_{\xi}:\xi\in\kappa\}$, then we will write $\delta=\lim_{\xi\rightarrow\kappa}\delta_{\xi}$. Let $\mathop{\rm Lim}\nolimits_{0}:=\{0\}\cup\mathop{\rm Lim}\nolimits$.
Lemma 4.1
If $\delta>0$ is an ordinal and $\mathop{\rm cf}\nolimits(\delta)=\kappa$, then exactly one of the following conditions is satisfied.
  • (A)
    $\delta$ is a successor ordinal or $\mathop{\rm cf}\nolimits(\delta)=\omega$ (that is, $\kappa\leq\omega$);
  • (B)
    $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=\omega$,
    and $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$;
  • (C)
    $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)>\omega$,
    and $\theta=\lim_{\xi\rightarrow\,\lambda}\delta_{\xi}$, where $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi\in\lambda:=\mathop{\rm cf}\nolimits(\theta)$;
  • (D)
    $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\mathop{\rm cf}\nolimits(\theta)\geq\kappa>\omega$ or $\theta=0$;
  • (E)
    $\delta=\lim_{\xi\rightarrow\kappa}\delta_{\xi}$, where $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa>\omega$, for all $\xi\in\kappa$.
Proof. Let $\kappa>\omega$. Since $\delta\geq\mathop{\rm cf}\nolimits(\delta)=\kappa$, the ordinal $\delta$ can be presented in the Cantor normal form in the base $\kappa$
$\delta=\kappa^{\xi_{n}}\zeta_{n}+\kappa^{\xi_{n-1}}\zeta_{n-1}+\dots+\kappa^{\xi_{1}}\zeta_{1}+\beta,$ (13)
where $n\in{\mathbb{N}}$, $1\leq\zeta_{k}<\kappa$, for $k\leq n$, $0\leq\beta<\kappa$ and $1\leq\xi_{1}<\xi_{2}<\dots<\xi_{n}$. First, $\beta>0$ would mean that $\beta$, and, hence, $\delta$, is a successor ordinal or a limit ordinal of cofinality $<\kappa$; so $\beta=0$.
Second, since $\zeta_{1}<\kappa$, $\zeta_{1}\in\mathop{\rm Lim}\nolimits$ would imply that $\mathop{\rm cf}\nolimits(\zeta_{1})<\kappa$ and the final part $\kappa^{\xi_{1}}\zeta_{1}$ of $\delta$, which is a $\zeta_{1}$-sum of copies of $\kappa^{\xi_{1}}$ would be of cofinality $<\kappa$; this would imply that $\mathop{\rm cf}\nolimits(\delta)<\kappa$ as well. Thus $\kappa^{\xi_{1}}\zeta_{1}=\kappa^{\xi_{1}}(\gamma+m)=\kappa^{\xi_{1}}\gamma+\kappa^{\xi_{1}}+\dots+\kappa^{\xi_{1}}$, where $\gamma\in\mathop{\rm Lim}\nolimits_{0}\;\cap\;\kappa$ and $m\in{\mathbb{N}}$, so $\delta$ has a final part isomorphic to $\kappa^{\xi_{1}}$.
Third, $\omega\leq\mathop{\rm cf}\nolimits(\xi_{1})<\kappa$ would imply that $\mathop{\rm cf}\nolimits(\kappa^{\xi_{1}})<\kappa$ and, hence, $\mathop{\rm cf}\nolimits(\delta)<\kappa$. So either $\mathop{\rm cf}\nolimits(\xi_{1})=\kappa$ or $\xi_{1}=\gamma^{\prime}+m^{\prime}$, where $\gamma^{\prime}\in\mathop{\rm Lim}\nolimits_{0}$ and $m^{\prime}\in{\mathbb{N}}$. Thus defining
$\theta=\kappa^{\xi_{n}}\zeta_{n}+\dots+\kappa^{\xi_{2}}\zeta_{2}+\kappa^{\xi_{1}}\gamma+\kappa^{\xi_{1}}(m-1),$ (14)
where $\mbox{where }\;\zeta_{1}=\gamma+m\;\;\mbox{ and }\;\;\gamma\in\mathop{\rm Lim}\nolimits_{0}\;\cap\;\kappa\;\;\mbox{ and }\;\;m\in{\mathbb{N}},$ (15)
we have
$\delta$ $=$ $\theta+\kappa^{\xi_{1}},\mbox{ where }\mathop{\rm cf}\nolimits(\xi_{1})=\kappa,\mbox{ or }$ (16)
$\delta$ $=$ $\theta+\kappa^{\gamma^{\prime}+m^{\prime}},\mbox{ where }\gamma^{\prime}\in\mathop{\rm Lim}\nolimits_{0}\mbox{ and }m^{\prime}\in{\mathbb{N}}.$ (17)
If (16) holds, let $\langle\alpha_{\xi}:\xi<\kappa\rangle$ be an increasing sequence cofinal in $\xi_{1}$ and let us define $\delta_{\xi}:=\theta+\kappa^{\alpha_{\xi}+1}$, for $\xi<\kappa$. Then for $\xi<\xi^{\prime}<\kappa$ we have $\delta_{\xi}<\delta_{\xi^{\prime}}<\delta$ and $\sup\{\delta_{\xi}:\xi<\kappa\}=\theta+\sup\{\kappa^{\alpha_{\xi}+1}:\xi<\kappa\}=\theta+\kappa^{\xi_{1}}=\delta$. So the sequence $\langle\delta_{\xi}:\xi\in\kappa\rangle$ is increasing, cofinal in $\delta$ and $\mathop{\rm cf}\nolimits(\delta_{\xi})=\mathop{\rm cf}\nolimits(\kappa^{\alpha_{\xi}}\kappa)=\kappa$, because $\kappa^{\alpha_{\xi}}\kappa$ is a $\kappa$-sum. So (E) is true.
If (17) holds and $\gamma^{\prime}+m^{\prime}\geq 2$, then we have $\alpha:=\gamma^{\prime}+m^{\prime}-1\geq 1$ and $\delta=\theta+\kappa^{\alpha}\kappa$. Thus the final part $\kappa^{\alpha}\kappa$ of $\delta$ is a $\kappa$-sum of copies of the ordinal $\kappa^{\alpha}\geq\kappa$. Now defining the ordinals $\delta_{\xi}:=\theta+\kappa^{\alpha}\xi+\kappa$, for $\xi<\kappa$, we have
$\kappa^{\alpha}\xi+\kappa\leq\kappa^{\alpha}\xi+\kappa^{\alpha}=\kappa^{\alpha}(\xi+1)<\kappa^{\alpha}(\xi+1)+\kappa\leq\kappa^{\alpha}(\xi+2)<\kappa^{\alpha}\kappa$
and, hence, $\delta_{\xi}<\delta$. Clearly we have $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi<\kappa$, and $\langle\delta_{\xi}:\xi<\kappa\rangle$ is an increasing sequence cofinal in $\delta$. So (E) is true again.
If (17) holds and $\gamma^{\prime}+m^{\prime}<2$, then $\gamma^{\prime}=0$, $m^{\prime}=1$, $\xi_{1}=1$ and $\delta=\theta+\kappa$, where by (14) and (15) we have
$\theta=(\kappa^{\xi_{n}}\zeta_{n}+\dots+\kappa^{\xi_{2}}\zeta_{2})+\kappa(\gamma+m-1)=\theta^{\prime}+\kappa\alpha,$ (18)
where $\theta^{\prime}=0$, if $n=1$, and $\alpha:=\gamma+m-1$ is an ordinal satisfying $0\leq\alpha<\kappa$. If $\theta=0$, then we have (D).
If $\theta>0$, then $\alpha>0$ or $\theta^{\prime}>0$ (and, hence, $\xi_{2}>1$ and $\zeta_{2}\geq 1$), which gives $\theta\geq\kappa$. In addition, the equality (18) presents the ordinal $\theta$ in the Cantor normal form and, because of its uniqueness, $\theta$ can not be presented in the form $\theta_{1}+\beta$, where $0<\beta<\kappa$; thus, $\theta$ is a limit ordinal. First, if $\mathop{\rm cf}\nolimits(\theta)\geq\kappa$, we have (D).
Second, if $\lambda:=\mathop{\rm cf}\nolimits(\theta)<\kappa$, let $\langle\alpha_{\xi}:\xi<\lambda\rangle$ be an increasing sequence cofinal in $\theta$. Assuming that $\alpha_{\xi}+\kappa\geq\theta$, for some $\xi<\lambda$, since $\mathop{\rm cf}\nolimits(\theta)<\kappa=\mathop{\rm cf}\nolimits(\alpha_{\xi}+\kappa)$ we would have $\alpha_{\xi}+\kappa>\theta$, and, hence, $\theta=\alpha_{\xi}+\beta$, where $0<\beta<\kappa$, which is false. So $\alpha_{\xi}+\kappa<\theta$, for all $\xi<\lambda$, and $\theta=\sup\{\alpha_{\xi}:\xi<\lambda\}\leq\sup\{\alpha_{\xi}+\kappa:\xi<\lambda\}\leq\theta$. Defining $\delta_{\xi}=\alpha_{\xi}+\kappa$, for $\xi<\lambda$, we have $\theta=\sup\{\delta_{\xi}:\xi<\lambda\}$ and $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi<\lambda$. Now, if $\lambda=\omega$ we have (B) and if $\lambda>\omega$ we have (C) ((C) is possible if $\kappa>\omega_{1}$).
Finally we show that each $\delta$ satisfies exactly one of conditions (A)–(E). If (A) is true, then $\kappa\leq\omega$ and (B)–(E) are false, because (B)–(E) imply that $\kappa>\omega$. If (E) holds, then (B)–(D) are false, since (B)–(D) imply that $\delta=\theta+\kappa$ and for an ordinal $\beta\in(\theta,\theta+\kappa)$ we have $\mathop{\rm cf}\nolimits(\beta)<\kappa$.
Suppose that $\delta$ satisfies (B) and (C). Then $\delta=I+F=I^{\prime}+F^{\prime}$, where $F\cong F^{\prime}\cong\kappa$, $I\cong\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=\omega$, and $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$; and $I^{\prime}\cong\theta^{\prime}\geq\kappa>\mathop{\rm cf}\nolimits(\theta^{\prime})=\lambda>\omega$, and $\theta^{\prime}=\lim_{\xi\rightarrow\,\lambda}\delta_{\xi}$, where $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi\in\lambda$. Since $\mathop{\rm cf}\nolimits(\theta)<\mathop{\rm cf}\nolimits(\theta^{\prime})$ we have $I\neq I^{\prime}$; so either $I\varsubsetneq I^{\prime}$ or $I^{\prime}\varsubsetneq I$. If $I\varsubsetneq I^{\prime}$, then there is $\xi\in\lambda$ such that $\delta_{\xi}\in I^{\prime}\setminus I$; thus $\delta_{\xi}\in(\theta,\theta+\kappa)$ and, as above, $\mathop{\rm cf}\nolimits(\delta_{\xi})<\kappa$, which is false because $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$. If $I^{\prime}\varsubsetneq I$, then there is $n\in\omega$ such that $\delta_{n}\in I\setminus I^{\prime}$; thus $\delta_{n}\in(\theta^{\prime},\theta^{\prime}+\kappa)$ and, hence, $\mathop{\rm cf}\nolimits(\delta_{n})<\kappa$, which is false since $\,\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$. Thus, conditions (B) and (C) are disjoint and in the same way we show that the pairs (B),(D) and (C),(D) are disjoint as well. $\Box$
Example 4.2
Regarding the classification given in Lemma 4.1 and the ordinals $\delta$ of, say, size $\omega_{2}$, (A) holds for $\omega_{2}+1$ and $\omega_{2}+\omega$, (B) for $\omega_{2}\omega+\omega_{2}$, (C) for $\omega_{2}\omega_{1}+\omega_{2}$, (D) for $\omega_{2}$ and $\omega_{2}+\omega_{2}$, and (E) holds for $\omega_{2}\omega_{2}$.
The separative modification of ${\mathbb{P}}(\omega^{\delta})$
We note that if $A\in{\mathbb{P}}({\mathbb{X}})$, then ${\mathbb{P}}({\mathbb{A}})=\{C\subset A:C\in{\mathbb{P}}({\mathbb{X}})\}$.
Proposition 4.3
Let $\delta>0$ be an ordinal and $A,B\in{\mathbb{P}}(\omega^{\delta})$. Then we have
(a) ${\mathbb{P}}(\omega^{\delta})=\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle\equiv_{forc}\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset_{{\mathcal{I}}_{\omega^{\delta}}}\rangle\equiv_{forc}(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$;
(b) $A\subset_{{\mathcal{I}}_{\omega^{\delta}}}B$ iff $\forall C\in{\mathbb{P}}({\mathbb{A}})\;C\cap B\in{\mathbb{P}}(\omega^{\delta})$;
(c) $A\not\perp_{\subset}B$ iff $A\not\perp_{\subset_{{\mathcal{I}}_{\omega^{\delta}}}}B$ iff $A\cap B\in{\mathbb{P}}(\omega^{\delta})$.
Proof. By Fact 2.11(f), ${\mathcal{I}}_{\omega^{\delta}}=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}$ is an ideal in $P(\omega^{\delta})$. Clearly we have ${\mathbb{P}}(\omega^{\delta})\subset{\mathcal{I}}_{\omega^{\delta}}^{+}$. Conversely, if $A\in{\mathcal{I}}_{\omega^{\delta}}^{+}$ and $C\in{\mathbb{P}}(\omega^{\delta})$, where $C\subset A$, then $\omega^{\delta}\hookrightarrow A\hookrightarrow\omega^{\delta}$, which implies that ${\mathbb{A}}\cong\omega^{\delta}$ and, hence, $A\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(\omega^{\delta})={\mathcal{I}}_{\omega^{\delta}}^{+}$. By Fact 2.2(f) we have $\mathop{\rm sq}\nolimits\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle=(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$ and by Fact 2.2(a), $\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle\equiv_{forc}(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$.
(b) By (8), $A\subset_{{\mathcal{I}}_{\omega^{\delta}}}B$ iff for each $C\in{\mathbb{P}}(\omega^{\delta})$ satisfying $C\subset A$ the set $C\cap B$ contains a copy of $\omega^{\delta}$, which by (a) means that $C\cap B\in{\mathbb{P}}(\omega^{\delta})$.
(c) follows from (a) and Fact 2.2(g). $\Box$
Lemma 4.4
Let $\delta$ be an ordinal. Then
(a) If $\mathop{\rm cf}\nolimits(\delta)=\kappa\geq\omega$ and $\delta=\lim_{\alpha\rightarrow\kappa}\delta_{\alpha}$, then $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}:=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, for $\alpha<\kappa$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, for $\alpha\neq\beta$.
(b) If $\delta=\delta^{\prime}+1$, then $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}\cong\sum_{\omega}\omega^{\delta^{\prime}}=\sum_{\alpha<\omega}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta^{\prime}}$, for $\alpha<\omega$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, for $\alpha\neq\beta$.
Proof. (a) By the assumptions $\delta$ is a limit ordinal; so, $\omega^{\delta}=\sup\{\omega^{\eta}:\eta<\delta\}$, and for $\eta<\delta$ there is $\alpha<\kappa$ such that $\eta<\delta_{\alpha}$, which implies $\omega^{\eta}<\omega^{\delta_{\alpha}}<\omega^{\delta}$. Thus $\langle\omega^{\delta_{\alpha}}:\alpha<\kappa\rangle$ is an increasing cofinal sequence in $\omega^{\delta}$ and, hence, $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$.
Let $\theta:=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$ and w.l.o.g. suppose that $\theta=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, $\alpha<\kappa$, are disjoint. Clearly we have $\theta\geq\omega^{\delta_{\alpha}}$, for all $\alpha<\kappa$; so, $\theta\geq\sup\{\omega^{\delta_{\alpha}}:\alpha<\kappa\}=\omega^{\delta}$. Suppose that $\omega^{\delta}<\theta$. Then we would have $\omega^{\delta}\in L_{\alpha_{0}}$, for some $\alpha_{0}<\kappa$; let $\xi=\min L_{\alpha_{0}}$. By Fact 2.11(b), $\omega^{\delta}>\xi$ would imply that $\omega^{\delta}\hookrightarrow L_{\alpha_{0}}\cong\omega^{\delta_{\alpha_{0}}}$ which is impossible because $\delta_{\alpha_{0}}<\delta$ and, hence, $\omega^{\delta_{\alpha_{0}}}<\omega^{\delta}$. So $\omega^{\delta}=\xi$ and, hence, $\omega^{\delta}=\sum_{\alpha<\alpha_{0}}\omega^{\delta_{\alpha}}$. Now, if $\alpha_{0}=\alpha+1$ we would have $\omega^{\delta}\hookrightarrow L_{\alpha}\cong\omega^{\delta_{\alpha}}$, which is impossible, and if $\alpha_{0}$ is a limit ordinal we would have $\mathop{\rm cf}\nolimits(\omega^{\delta})\leq\mathop{\rm cf}\nolimits(\alpha_{0})\leq|\alpha_{0}|<\kappa$, which is impossible too. So, $\omega^{\delta}=\theta=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$.
(b) If $\delta=\delta^{\prime}+1$, then $\omega^{\delta}=\omega^{\delta^{\prime}+1}=\omega^{\delta^{\prime}}\omega=\sum_{\omega}\omega^{\delta^{\prime}}$ and $\langle\omega^{\delta^{\prime}}n:n\in\omega\rangle$ is an increasing and cofinal sequence in $\omega^{\delta}$, which gives $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$. $\Box$
Fact 4.5
If $\delta>0$ is an ordinal and $\omega^{\delta}\hookrightarrow L=\sum_{i\in I}L_{i}$, where $I$ and $L_{i}$, $i\in I$, are linear orders and $|I|<\mathop{\rm cf}\nolimits(\omega^{\delta})$, then $\omega^{\delta}\hookrightarrow L_{i}$, for some $i\in I$.
Proof. Let $\kappa=\mathop{\rm cf}\nolimits(\omega^{\delta})$ and let $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ be an increasing cofinal sequence in $\omega^{\delta}$. Then $\omega^{\delta}=\sup\{\beta_{\alpha}:\alpha<\kappa\}$ and for $\alpha<\kappa$ we have $g(\beta_{\alpha})\in\bigcup_{i\in I}L_{i}$, which implies that there is a unique $i_{\alpha}\in I$ such that $g(\beta_{\alpha})\in L_{i_{\alpha}}$. Clearly, the function $h:\kappa\rightarrow I$ defined by $h(\alpha)=i_{\alpha}$ is non-decreasing and, since $|I|<\kappa=\bigcup_{i\in I}h^{-1}[\{i\}]$, by the regularity of $\kappa$ there is $i^{*}\in I$ such that $h^{-1}[\{i^{*}\}]\in[\kappa]^{\kappa}$. Let $h^{-1}[\{i^{*}\}]=\{\alpha_{\xi}:\xi<\ \kappa\}$ be an increasing enumeration. Now, if $\alpha\in\kappa\setminus\alpha_{0}$ and $\xi\in\kappa$ where $\alpha<\alpha_{\xi}$, then $i^{*}=h(\alpha_{0})\leq h(\alpha)\leq h(\alpha_{\xi})=i^{*}$ thus for each $\alpha\in[\alpha_{0},\kappa)$ we have $h(\alpha)=i^{*}$, that is $g(\beta_{\alpha})\in L_{i^{*}}$. So, if $\beta_{\alpha_{0}}\leq\theta<\omega^{\delta}$, then, by the cofinality of the sequence $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ in $\omega^{\delta}$, there is $\alpha\in\kappa$ such that $\theta\leq\beta_{\alpha}$, which implies that $g(\beta_{\alpha_{0}})\leq g(\theta)\leq g(\beta_{\alpha})$. So, since $g(\beta_{\alpha_{0}}),g(\beta_{\alpha})\in L_{i^{*}}$ and $L_{i^{*}}$ is a convex suborder of $L$, we have $g(\theta)\in L_{i^{*}}$, for all $\theta\in[\beta_{\alpha_{0}},\omega^{\delta})$. By Fact 2.11(b) we have $\omega^{\delta}\cong[\beta_{\alpha_{0}},\omega^{\delta})\cong g[[\beta_{\alpha_{0}},\omega^{\delta})]\subset L_{i^{*}}$, that is $\omega^{\delta}\hookrightarrow L_{i^{*}}$. $\Box$
Proposition 4.6
Let $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$, where $\delta_{\alpha}\leq\delta_{\beta}$, whenever $\alpha<\beta<\kappa$. If $L:=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, for $\alpha<\kappa$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, for $\alpha\neq\beta$, then
(a) Each $C\in{\mathbb{P}}(L)$ intersects $\kappa$-many summands $L_{\alpha}$;
(b) If $C\subset L$, then $C\in{\mathbb{P}}(L)$ iff
$\forall\alpha,\eta<\kappa\;\;\exists K\in[\kappa\setminus\eta]^{<\kappa}\;\;L_{\alpha}\hookrightarrow\bigcup_{\beta\in K}L_{\beta}\cap C;$ (19)
(c) If, in addition, $\mathop{\rm cf}\nolimits(\omega^{\delta_{\alpha}})\geq\kappa$, for all $\alpha<\kappa$,
$C\in{\mathbb{P}}(L)\Leftrightarrow\forall\alpha<\kappa\;\;S^{\alpha}_{C}:=\{\beta<\kappa:\omega^{\delta_{\alpha}}\hookrightarrow L_{\beta}\cap C\}\in[\kappa]^{\kappa}.$ (20)
Proof. (a) Let $C\in{\mathbb{P}}(L)$ and suppose that $C\subset\bigcup_{\alpha<\eta}L_{\alpha}$, for some $\eta<\kappa$. By Fact 2.11(b) we have $L\cong\bigcup_{\alpha\geq\eta}L_{\alpha}$ and, hence, $L+L\hookrightarrow L$, which is, by Fact 2.10(a), impossible.
(b) Let $C\in{\mathbb{P}}(L)$, let $f:L\hookrightarrow L$, where $C=f[L]=\sum_{\alpha\in\kappa}f[L_{\alpha}]$ and let $\alpha_{0},\eta_{0}\in\kappa$. By (a) there are $\alpha_{1}<\kappa$ and $\eta_{1}\geq\eta_{0}$ such that $f[L_{\alpha_{1}}]\cap L_{\eta_{1}}\neq\emptyset$ and, hence, $C_{1}:=\bigcup_{\alpha>\alpha_{1}}f[L_{\alpha}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. Since $C_{1}$ is a final part of $C$ and $C\cong L$, by Fact 2.11(b) there is $g:L\hookrightarrow\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$ and, hence, $g[L_{\alpha_{0}}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. If $\beta_{0}\geq\eta_{0}$, where $g[L_{\alpha_{0}+1}]\cap L_{\beta_{0}}\neq\emptyset$, then, since $g[L_{\alpha_{0}}]<g[L_{\alpha_{0}+1}]$, we have $g[L_{\alpha_{0}}]\subset\bigcup_{\eta_{0}\leq\beta\leq\beta_{0}}L_{\beta}\cap C$. Thus $L_{\alpha_{0}}\hookrightarrow\bigcup_{\beta\in K}L_{\beta}\cap C$, where $K=[\eta_{0},\beta_{0}]\in[\kappa\setminus\eta_{0}]^{<\kappa}$, and (19) is true.
Assuming (19) by recursion we define sequences $\langle K_{\alpha}:\alpha\in\kappa\rangle$ and $\langle f_{\alpha}:\alpha\in\kappa\rangle$ such that for each $\alpha,\alpha^{\prime}\in\kappa$ we have:
(i) $K_{\alpha}\in[\kappa]^{<\kappa}$,
(ii) $\alpha<\alpha^{\prime}\Rightarrow K_{\alpha}<K_{\alpha^{\prime}}$ and
(iii) $f_{\alpha}:L_{\alpha}\hookrightarrow\bigcup_{\beta\in K_{\alpha}}L_{\beta}\cap C$.
By (19) for $\alpha=\eta=0$ there are $K_{0}\in[\kappa]^{<\kappa}$ and $f_{0}:L_{0}\hookrightarrow\bigcup_{\beta\in K_{0}}L_{\beta}\cap C$.
Let $\alpha^{\prime}<\kappa$ and let $K_{\alpha}$ and $f_{\alpha}$ be defined for $\alpha<\alpha^{\prime}$ such that (i)-(iii) is true. By (i) and the regularity of $\kappa$ we have $\eta:=\sup(\bigcup_{\alpha<\alpha^{\prime}}K_{\alpha})+1<\kappa$ and, by (19), there are $K_{\alpha^{\prime}}\in[\kappa\setminus\eta]^{<\kappa}$ and $f_{\alpha^{\prime}}:L_{\alpha^{\prime}}\hookrightarrow\bigcup_{\beta\in K_{\alpha^{\prime}}}L_{\beta}\cap C$; so, the recursion works.
Let $f=\bigcup_{\alpha\in\kappa}f_{\alpha}$. By (ii) and (iii), $\alpha<\alpha^{\prime}$ implies $f_{\alpha}[L_{\alpha}]<f_{\alpha^{\prime}}[L_{\alpha^{\prime}}]$ and, hence, $f:L\hookrightarrow C$. Thus $C^{\prime}=f[L]\in{\mathbb{P}}(L)$, $C^{\prime}\subset C$, and by Proposition 4.3(a) we have $C\in{\mathbb{P}}(L)$.
(c) By (b) and Fact 4.5 we have $C\in{\mathbb{P}}(L)$ iff for each $\alpha<\kappa$, for each $\eta<\kappa$ there is $\beta\geq\eta$ such that $\omega^{\delta_{\alpha}}\hookrightarrow L_{\beta}\cap C$. $\Box$
Complete embeddings of ${\mathcal{P}}_{\kappa}$
Theorem 4.7
If $\delta$ is an ordinal, $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa\geq\omega$, $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$, where $\delta_{\xi}\leq\delta_{\zeta}$, whenever $\xi<\zeta<\kappa$, and $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})\geq\kappa$, for all $\xi<\kappa$, then $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Proof. Let ${\mathbb{L}}_{\xi}\cong\omega^{\delta_{\xi}}$, for $\xi<\kappa$, where $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Then $\omega^{\delta}\cong{\mathbb{L}}:=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$ and, by (d) and (f) of Fact 2.2 and Proposition 4.3(a), $\langle P(\omega^{\delta})\setminus{\mathcal{I}}_{\omega^{\delta}},\subset_{\omega^{\delta}}\rangle=\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}({\mathbb{L}})=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$. So, we show that $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$. For $A\subset L$ let $\mathop{\rm supp}\nolimits(A)=\{\xi<\kappa:L_{\xi}\cap A\neq\emptyset\}$. By Proposition 4.6(c) we have
$A\in P(L)\setminus{\mathcal{I}}\Leftrightarrow\forall\zeta<\kappa\;\;S^{\zeta}_{A}:=\{\xi<\kappa:\omega^{\delta_{\zeta}}\hookrightarrow L_{\xi}\cap A\}\in[\kappa]^{\kappa},$ (21)
and, since $S^{\zeta}_{A}\subset\mathop{\rm supp}\nolimits(A)$, for $\zeta<\kappa$,
$|\mathop{\rm supp}\nolimits(A)|<\kappa\Rightarrow A\in{\mathcal{I}}.$ (22)
We show that $f:\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where $f(S)=\bigcup_{\xi\in S}L_{\xi}$, for $S\in[\kappa]^{\kappa}$. First, if $S\in[\kappa]^{\kappa}$ and $\zeta<\kappa$, then $|S\setminus\zeta|=\kappa$ and for $\xi\in S\setminus\zeta$ we have $\delta_{\zeta}\leq\delta_{\xi}$, which gives $\omega^{\delta_{\zeta}}\hookrightarrow\omega^{\delta_{\xi}}\cong L_{\xi}=L_{\xi}\cap f(S)$ and, hence, $\xi\in S^{\zeta}_{f(S)}$. Thus $[\kappa]^{\kappa}\ni S\setminus\zeta\subset S^{\zeta}_{f(S)}$, for each $\zeta<\kappa$, and by (21) we have $f(S)\in P(L)\setminus{\mathcal{I}}$. So, $f$ is a mapping from $[\kappa]^{\kappa}$ to $P(L)\setminus{\mathcal{I}}$ indeed. Let $S,T\in[\kappa]^{\kappa}$.
(ce1) If $S\subset^{*}T$, then $|\mathop{\rm supp}\nolimits(f(S)\setminus f(T))|=|\mathop{\rm supp}\nolimits(\bigcup_{\xi\in S\setminus T}L_{\xi})|=|S\setminus T|<\kappa$ and by (22) $f(S)\setminus f(T)\in{\mathcal{I}}$, that is $f(S)\subset_{\mathcal{I}}f(T)$.
(ce2) If $S\perp T$, then $|\mathop{\rm supp}\nolimits(f(S)\cap f(T))|=|\mathop{\rm supp}\nolimits(\bigcup_{\xi\in S\cap T}L_{\xi})|=|S\cap T|<\kappa$ and by (22) $f(S)\cap f(T)\in{\mathcal{I}}$, that is $f(S)\perp_{\subset_{\mathcal{I}}}f(T)$. If $S\not\perp T$, then $S\cap T\in[\kappa]^{\kappa}$ and $f(S)\cap f(T)=\bigcup_{\xi\in S\cap T}L_{\xi}=f(S\cap T)\in P(L)\setminus{\mathcal{I}}$ and, hence, $f(S)\not\perp_{\subset_{\mathcal{I}}}f(T)$.
(ce3) First we show that a set $S\in[\kappa]^{\kappa}$ is a reduction of $A\in P(L)\setminus{\mathcal{I}}$ to $\langle[\kappa]^{\kappa},\subset^{*}\rangle$, if $S$ is a pseudointersection of the family $\{S^{\zeta}_{A}:\zeta<\kappa\}$, that is
$(\forall\zeta<\kappa\;\;S\subset^{*}S^{\zeta}_{A})\;\Rightarrow\;\forall T\in[\kappa]^{\kappa}\;\;(T\subset^{*}S\Rightarrow f(T)\cap A\in P(L)\setminus{\mathcal{I}}).$ (23)
So, let $[\kappa]^{\kappa}\ni S\subset^{*}S^{\zeta}_{A}$, for all $\zeta<\kappa$, and $[\kappa]^{\kappa}\ni T\subset^{*}S$. Then for each $\zeta<\kappa$ we have $T\subset^{*}S^{\zeta}_{A}$, which implies that
$\forall\zeta<\kappa\;\;|T\cap S^{\zeta}_{A}|=\kappa.$ (24)
By recursion we construct a sequence $\langle\alpha_{\xi}:\xi<\kappa\rangle$ such that for each $\xi<\kappa$ we have
(i) $\alpha_{\xi}\in T$,
(ii) $\xi^{\prime}<\xi\Rightarrow\alpha_{\xi^{\prime}}<\alpha_{\xi}$ and
(iii) $\omega^{\delta_{\xi}}\hookrightarrow L_{\alpha_{\xi}}\cap A$.
By (24) there is $\alpha_{0}\in T\cap S^{0}_{A}$ and, hence, $\omega^{\delta_{0}}\hookrightarrow L_{\alpha_{0}}\cap A$. Let $\zeta<\kappa$ and let $\langle\alpha_{\xi}:\xi<\zeta\rangle$ be a sequence satisfying (i)–(iii). By (24) we have $|T\cap S^{\zeta}_{A}|=\kappa$ so there is $\alpha_{\zeta}=\min(T\cap S^{\zeta}_{A}\setminus(\sup\{\alpha_{\xi}:\xi<\zeta\}+1))$. Now we have $\alpha_{\zeta}\in T$ and $\alpha_{\zeta}>\alpha_{\xi}$, for $\xi<\zeta$, so (i) and (ii) are true. Since $\alpha_{\zeta}\in S^{\zeta}_{A}$ we have $\omega^{\delta_{\zeta}}\hookrightarrow L_{\alpha_{\zeta}}\cap A$ so (iii) is true and the recursion works.
Clearly $B:=\bigcup_{\xi<\kappa}L_{\alpha_{\xi}}\cap A\subset A$ and by (i) we have $B\subset\bigcup_{\xi<\kappa}L_{\alpha_{\xi}}\subset\bigcup_{\zeta\in T}L_{\zeta}=f(T)$. If $\zeta<\kappa$, then for $\xi\geq\zeta$ we have $L_{\alpha_{\xi}}\cap B=L_{\alpha_{\xi}}\cap A$ and, by (iii) and since $\delta_{\zeta}\leq\delta_{\xi}$, we have $\omega^{\delta_{\zeta}}\hookrightarrow\omega^{\delta_{\xi}}\hookrightarrow L_{\alpha_{\xi}}\cap A=L_{\alpha_{\xi}}\cap B$, which implies that $\alpha_{\xi}\in S^{\zeta}_{B}$. So for each $\zeta<\kappa$ we have $\{\alpha_{\xi}:\xi\geq\zeta\}\subset S^{\zeta}_{B}$, which by (ii) implies that $S^{\zeta}_{B}\in[\kappa]^{\kappa}$. Thus, by (21), $B\in P(L)\setminus{\mathcal{I}}$ and, since $B\subset f(T)\cap A$, we have $f(T)\cap A\in P(L)\setminus{\mathcal{I}}$. So (23) is true.
Finally we prove (ce3). If $A\in P(L)\setminus{\mathcal{I}}$, then by (21) and since $\delta_{0}\leq\delta_{1}\leq\dots$ we have
$S^{0}_{A}\supset S^{1}_{A}\supset\dots\supset S^{\zeta}_{A}\supset\dots\;\;\mbox{ and }\;\;\forall\zeta<\kappa\;\;|S^{\zeta}_{A}|=\kappa.$ (25)
So we can take $\alpha_{0}=\min S^{0}_{A}$ and $\alpha_{\zeta}=\min(S^{\zeta}_{A}\setminus\{\alpha_{\xi}:\xi<\zeta\})$, for $\zeta<\kappa$. Now we have $S:=\{\alpha_{\zeta}:\zeta<\kappa\}\in[\kappa]^{\kappa}$ and, by (25), $\{\alpha_{\xi}:\xi\geq\zeta\}\subset S^{\zeta}_{A}$, which implies that $S\subset^{*}S^{\zeta}_{A}$, for all $\zeta<\kappa$. By (23), $S$ is a reduction of $A$ to $\langle[\kappa]^{\kappa},\subset^{*}\rangle$. $\Box$
Theorem 4.8
If $\delta=\theta+\kappa$, where $\theta$ is an ordinal and $\theta\geq\kappa=\mathop{\rm cf}\nolimits(\delta)>\mathop{\rm cf}\nolimits(\theta)=:\lambda\geq\omega$, then $\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Proof. By Lemma 4.1, $\theta=\lim_{\nu\rightarrow\lambda}\delta_{\nu}$, where $\mathop{\rm cf}\nolimits(\delta_{\nu})=\kappa$, for all $\nu<\lambda$. Clearly we have $\omega^{\delta}=\omega^{\theta}\omega^{\kappa}=\sup\{\omega^{\delta_{\nu}}:\nu<\lambda\}\kappa$. Let ${\mathbb{L}}=\langle L,<\rangle=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$, where ${\mathbb{L}}_{\xi}=\langle L_{\xi},<_{\xi}\rangle\cong\omega^{\theta}$, for $\xi<\kappa$, and $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Let =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$ and, for $A\subset L$ and $\nu<\lambda$, let
$S^{\nu}_{A}=\{\xi<\kappa:\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap A\}.$
Claim 4.9
For $A,B\subset L$ we have:
(a) $A\in P(L)\setminus{\mathcal{I}}\;\Leftrightarrow\;\forall\nu<\lambda\;\;|S^{\nu}_{A}|=\kappa$;
(b) $A\subset_{{\mathcal{I}}}B\;\Leftrightarrow\;A\setminus B\in{\mathcal{I}}\;\Leftrightarrow\;\exists\nu<\lambda\;\;|S^{\nu}_{A\setminus B}|<\kappa$.
Proof. (a) ($\Rightarrow$) Since ${\mathbb{L}}\cong\omega^{\theta}\kappa$, by Proposition 4.6(b) we have: $A\in P(L)\setminus{\mathcal{I}}$ iff
$\forall\eta<\kappa\;\;\exists K\in[\kappa\setminus\eta]^{<\kappa}\;\;\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$ (26)
and we prove
$\forall\nu<\lambda\;\;\forall\eta<\kappa\;\;\exists\xi\in\kappa\setminus\eta\;\;\xi\in S^{\nu}_{A}.$ (27)
If $\nu<\lambda$ and $\eta<\kappa$, then by (26) there is $K\in[\kappa\setminus\eta]^{<\kappa}$ such that $\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$ which implies that $\omega^{\delta_{\nu}}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$ and, since $|K|<\kappa=\mathop{\rm cf}\nolimits(\delta_{\nu})=\mathop{\rm cf}\nolimits(\omega^{\delta_{\nu}})$, by Fact 4.5 there is $\xi\in K$ such that $\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap A$ and, hence, $\xi\in S^{\nu}_{A}$. Since $K\subset\kappa\setminus\eta$ we have $\xi\geq\eta$.
($\Leftarrow$) Assuming (27) we prove (26). Let $\eta<\kappa$. By recursion we construct a sequence $\langle\xi_{\nu}:\nu<\lambda\rangle$ such that for all $\nu,\nu^{\prime}<\lambda$ we have:
(i) $\nu<\nu^{\prime}\Rightarrow\eta\leq\xi_{\nu}<\xi_{\nu^{\prime}}<\kappa$,
(ii) $\omega^{\delta_{\nu}}\hookrightarrow A\cap L_{\xi_{\nu}}$.
First, by (27), for $\nu=0$ there is $\xi_{0}\geq\eta$ such that $\xi_{0}\in S^{0}_{A}$ that is $\omega^{\delta_{0}}\hookrightarrow A\cap L_{\xi_{0}}$. If $\nu<\lambda$ and the sequence $\langle\xi_{\nu^{\prime}}:\nu^{\prime}<\nu\rangle$ satisfies (i) and (ii), then by (i) and since $\lambda<\kappa$ we have $\eta_{\nu}:=\sup\{\xi_{\nu^{\prime}}:\nu^{\prime}<\nu\}+1<\kappa$ and by (27) there is $\xi_{\nu}\geq\eta_{\nu}$ such that $\xi_{\nu}\in S^{\nu}_{A}$ that is $\omega^{\delta_{\nu}}\hookrightarrow A\cap L_{\xi_{\nu}}$. Thus the recursion works.
Now $K:=\{\xi_{\nu}:\nu<\lambda\}\in[\kappa\setminus\eta]^{<\kappa}$ and $\sum_{\nu<\lambda}\omega^{\delta_{\nu}}\hookrightarrow\bigcup_{\nu<\lambda}L_{\xi_{\nu}}\cap A$, that is $\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$. So (26) is true. (b) follows from (a). $\Box$
Let $\{I_{\nu}:\nu<\lambda\}\subset[\kappa]^{\kappa}$ be a partition of $\kappa$. For $\nu<\lambda$ and $\xi\in I_{\nu}$, let $L^{\prime}_{\xi}\subset L_{\xi}$ be the initial part of ${\mathbb{L}}_{\xi}$ such that ${\mathbb{L}}_{\xi}^{\prime}\cong\omega^{\delta_{\nu}}$ and let $C=\bigcup_{\nu<\lambda}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$. Then for each $\nu<\lambda$ the set $S^{\nu}_{C}=\{\xi<\kappa:\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap C\}=\bigcup_{\mu\geq\nu}I_{\mu}$ is of size $\kappa$ and, by Claim 4.9(a), $C\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(C)\cong{\mathbb{P}}(\omega^{\delta})$ and we prove that $f:\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(C)$, where for $S\in[\lambda]^{\lambda}$
$f(S)=C_{S}:=\bigcup_{\nu\in S}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}.$
First, since $S^{\nu}_{C_{S}}=\bigcup_{\mu\in S\setminus\nu}I_{\mu}$, for $\nu<\lambda$, by Claim 4.9(a) we have $C_{S}\in{\mathbb{P}}(\omega^{\delta})$ and, hence, $C_{S}\in{\mathbb{P}}(C)$ as well. Let $S,T\in[\lambda]^{\lambda}$.
(ce1) If $S\subset^{*}T$, then, since $f(S)\setminus f(T)=\bigcup_{\nu\in S\setminus T}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$, $|S\setminus T|<\lambda$ and $\lambda$ is a regular cardinal, for $\nu>\sup(S\setminus T)$ we have $S^{\nu}_{f(S)\setminus f(T)}=\emptyset$, which by Claim 4.9(b) gives $f(S)\subset_{\mathcal{I}}f(T)$.
(ce2) Since $f(S)\cap f(T)=\bigcup_{\nu\in S\cap T}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$, by Claim 4.9(a) we have $S\perp T$ iff $|S\cap T|<\lambda$ iff $f(S)\cap f(T)\in{\mathcal{I}}$ iff $f(S)\perp_{\mathcal{I}}f(T)$.
(ce3) For $A\in{\mathbb{P}}(C)$ we will find a set $S\in[\lambda]^{\lambda}$ such that $f(T)\cap A\in{\mathbb{P}}(C)$, whenever $[\lambda]^{\lambda}\ni T\subset^{*}S$. By recursion we construct a sequence $\langle\nu_{\alpha}:\alpha<\lambda\rangle$ in $\lambda$ such that
(i) $\nu_{0}=\min\{\nu\in\lambda:|S^{0}_{A}\cap I_{\nu}|=\kappa\}$,
(ii) $\nu_{\alpha+1}=\min\{\nu\in\lambda\setminus(\nu_{\alpha}+1):|S^{\nu_{\alpha}+1}_{A}\cap I_{\nu}|=\kappa\}$, for $\alpha<\lambda$,
(iii) $\nu_{\beta}=\sup\{\nu_{\alpha}:\alpha<\beta\}$, for $\beta\in\mathop{\rm Lim}\nolimits\cap\lambda$.
Since $A\in{\mathbb{P}}({\mathbb{L}})$, by Claim 4.9(a) we have $[\kappa]^{\kappa}\ni S^{0}_{A}\subset\bigcup_{\nu<\lambda}I_{\nu}$ and, since $\kappa$ is a regular cardinal and $\lambda<\kappa$, the ordinal $\nu_{0}$ is well defined by (i). Let $\beta<\lambda$ and let $\langle\nu_{\alpha}:\alpha<\beta\rangle$ be a sequence satisfying (i)–(iii). If $\beta=\alpha+1$, then for $\nu\leq\nu_{\alpha}$ and $\xi\in I_{\nu}$ we have $\mathop{\mathrm{tp}}\nolimits(L_{\xi}^{\prime})\leq\omega^{\delta_{\nu_{\alpha}}}$ and, since $A\subset C$, we have $\omega^{\delta_{\nu_{\alpha}+1}}\not\hookrightarrow L_{\xi}^{\prime}\cap A=L_{\xi}\cap A$. Thus $S^{\nu_{\alpha}+1}_{A}\subset\bigcup_{\nu\geq\nu_{\alpha}+1}I_{\nu}$ and, as above, there is $\nu\geq\nu_{\alpha}+1$ such that $|S^{\nu_{\alpha}+1}_{A}\cap I_{\nu}|=\kappa$. So $\nu_{\alpha+1}$ is well defined by (ii). If $\beta$ is a limit ordinal, then $\nu_{\beta}$ is well defined by (iii), because $\lambda$ is a regular cardinal and $\nu_{\alpha}<\lambda$, for all $\alpha<\beta$. Thus, the recursion works and by (ii) and (iii) we have $S:=\{\nu_{\alpha}:\alpha<\lambda\}\in[\lambda]^{\lambda}$.
Now, if $T\in[\lambda]^{\lambda}$ and $T\subset^{*}S$, then $|T\setminus S|<\lambda$ and, hence, there is $\beta<\lambda$ such that $\{\nu_{\alpha}:\alpha\geq\beta\}\subset T$. So, we have $\bigcup_{\alpha\geq\beta}\bigcup_{\xi\in I_{\nu_{\alpha}}}L_{\xi}^{\prime}\subset f(T)$, which gives $B:=\bigcup_{\alpha\geq\beta}\bigcup_{\xi\in I_{\nu_{\alpha}}}L_{\xi}^{\prime}\cap A\subset f(T)\cap A$. For $\alpha\geq\beta$ by (ii) we have $|S^{\nu_{\alpha}+1}_{A}\cap I_{\nu_{\alpha+1}}|=\kappa$, which means that there are $\kappa$-many $\xi\in I_{\nu_{\alpha+1}}$ such that $\omega^{\delta_{\nu_{\alpha}+1}}\hookrightarrow L_{\xi}\cap A=L_{\xi}\cap B$. Since the sequence $\langle\nu_{\alpha}+1:\alpha<\lambda\rangle$ is cofinal in $\lambda$, by Claim 4.9(a) we have $B\in{\mathbb{P}}(C)$ and, consequently, $f(T)\cap A\in{\mathbb{P}}(C)$. $\Box$
Factorization of factors ${\mathbb{P}}(\omega^{\delta})$
Theorem 4.10
If $\delta>0$ is an ordinal and $\mathop{\rm cf}\nolimits(\delta)=\kappa$, then regarding ${\mathbb{P}}(\omega^{\delta})$ we have
  • (A)
    If $\delta$ is a successor ordinal or $\mathop{\rm cf}\nolimits(\delta)=\omega$ (that is, $\kappa\leq\omega$),
    then $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed and $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
  • (B)
    If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=\omega$
    and $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$,
    then $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed and $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
  • (C)
    If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=:\lambda>\omega$
    and $\theta=\lim_{\xi\rightarrow\,\lambda}\delta_{\xi}$, where $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi\in\lambda$,
    then $\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
  • (D)
    If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\mathop{\rm cf}\nolimits(\theta)\geq\kappa>\omega$ or $\theta=0$,
    then $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$;
  • (E)
    If $\delta=\lim_{\xi\rightarrow\kappa}\delta_{\xi}$, where $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa>\omega$, for all $\xi<\kappa$,
    then $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
In cases (C), (D) and (E) forcing ${\mathbb{P}}(\omega^{\delta})$ collapses (at least) $\omega_{2}$ to $\omega$.
Proof. Case (A) (a) Let $\delta=\theta+1$; for $\theta=0$ we have $\omega^{\delta}=\omega$ and $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega)=\langle[\omega]^{\omega},\subset^{*}\rangle$. If $\theta\geq 1$, then by Lemma 4.4(b) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{\omega}\omega^{\theta}$. Since $\theta\geq 1$ we have $\mathop{\rm cf}\nolimits(\omega^{\theta})\geq\omega$; so, by Theorem 4.7 we have $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
If $\mathop{\rm cf}\nolimits\delta=\omega$, let $\delta=\sup\{\delta_{i}:i<\omega\}$, where $\langle\delta_{i}:i<\omega\rangle$ is an increasing sequence of non-zero ordinals. Then by Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{i<\omega}\omega^{\delta_{i}}$. For $i<\omega$ we have $\delta_{i}>0$ and, hence, $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$. Thus, by Theorem 4.7 we have $\langle[\omega]^{\omega},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
(b) By Lemma 4.4 we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}\cong\sum_{i<\omega}\omega^{\delta_{i}}=\sum_{i<\omega}L_{i}=:L$. Thus $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}(L)=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where :=${\mathcal{I}}:=\{A\subset L:L\not\hookrightarrow A\}$. For $\delta>1$ as in (a) we conclude that $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$, for all $i<\omega$, and, by Proposition 4.6(c), for $A\subset L$ we have
$A\in P(L)\setminus{\mathcal{I}}\;\;\Leftrightarrow\;\forall i\in\omega\;\;S^{i}_{A}:=\{j\in\omega:L_{i}\hookrightarrow L_{j}\cap A\}\in[\omega]^{\omega}.$ (28)
For $A_{n}\in P(L)\setminus{\mathcal{I}}$, $n\in\omega$, where $\dots\subset_{\mathcal{I}}A_{1}\subset_{\mathcal{I}}A_{0}$, we will construct $A\in P(L)\setminus{\mathcal{I}}$ such that $A\subset_{\mathcal{I}}A_{n}$, for all $n\in\omega$. Since ${\mathcal{I}}$ is an ideal we have $P(L)\setminus{\mathcal{I}}\ni B_{i}:=\bigcap_{n\leq i}A_{n}\subset A_{i}$, so we can assume that $A_{0}\supset A_{1}\supset A_{2}\supset\dots$. By recursion we define the sequences $\langle j_{i}:i\in\omega\rangle$ and $\langle f_{i}:i\in\omega\rangle$ such that for each $i\in\omega$ we have
(i) $j_{i}\in\omega$,
(ii) $j_{i}<j_{i+1}$,
(iii) $f_{i}:L_{i}\hookrightarrow L_{j_{i}}\cap A_{i}$.
Since $A_{0}\in P(L)\setminus{\mathcal{I}}$, by (28) for $i=0$ there are $j_{0}\in\omega$ and $f_{0}:L_{0}\hookrightarrow L_{j_{0}}\cap C_{0}$.
Let $j_{0},\dots,j_{i}$ and $f_{0},\dots,f_{i}$ be sequences satisfying (i)–(iii). Since $A_{i+1}\in P(L)\setminus{\mathcal{I}}$ by (28) for $i+1$ there is $j_{i+1}\in S^{i+1}_{A_{i+1}}\setminus(j_{i}+1)$ so we have $j_{i}<j_{i+1}$ and there is $f_{i+1}:L_{i+1}\hookrightarrow L_{j_{i+1}}\cap A_{i+1}$. The recursion works.
By (ii) $i_{1}<i_{2}$ implies $j_{i_{1}}<j_{i_{2}}$, and, hence, $L_{j_{i_{1}}}<L_{j_{i_{2}}}$, which by (iii) implies $f_{i_{1}}[L_{i_{1}}]<f_{i_{2}}[L_{i_{2}}]$. Thus $f=\bigcup_{i\in\omega}f_{i}:L\hookrightarrow L$ and
$A:=f[L]=\bigcup_{i\in\omega}f_{i}[L_{i}]\in P(L)\setminus{\mathcal{I}}.$ (29)
For $n^{*}\in\omega$ we show that $A\subset_{{\mathcal{I}}}A_{n^{*}}$, that is $A\setminus A_{n^{*}}\in{\mathcal{I}}$. By (iii) we have $A\subset\bigcup_{i\in\omega}L_{j_{i}}\cap A_{i}$ and, since $A_{i}\subset A_{n^{*}}$, for $i\geq n^{*}$, we have $A\setminus A_{n^{*}}\subset\bigcup_{i\in\omega}L_{j_{i}}\cap A_{i}\setminus A_{n^{*}}=\bigcup_{i<n^{*}}L_{j_{i}}\cap A_{i}\setminus A_{n^{*}}\subset\bigcup_{i<n^{*}}L_{j_{i}}$. Thus $S^{0}_{A\setminus A_{n^{*}}}:=\{j\in\omega:L_{0}\hookrightarrow L_{j}\cap A\setminus A_{n^{*}}\}\subset\{j_{i}:i<n^{*}\}$ and by (28) we have $A\setminus A_{n^{*}}\in{\mathcal{I}}$.
Case (B) (a) follows from Theorem 4.8, for $\lambda=\omega$.
(b) By Lemma 4.1, $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$. Clearly we have $\omega^{\delta}=\omega^{\theta}\omega^{\kappa}=\sup\{\omega^{\delta_{n}}:n\in\omega\}\kappa$. Let ${\mathbb{L}}=\langle L,<\rangle=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$, where ${\mathbb{L}}_{\xi}=\langle L_{\xi},<_{\xi}\rangle\cong\omega^{\theta}$, for $\xi<\kappa$, and $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Let =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$ and, for $A\subset L$ and $n\in\omega$, let
$S^{n}_{A}=\{\xi<\kappa:\omega^{\delta_{n}}\hookrightarrow L_{\xi}\cap A\}.$
Then by Claim 4.9 (where $\lambda=\omega$) for $A,B\subset L$ we have:
$A\in P(L)\setminus{\mathcal{I}}$$\Leftrightarrow$ $\forall n\in\omega\;\;|S^{n}_{A}|=\kappa,$ (30)
$A\subset_{{\mathcal{I}}}B$$\Leftrightarrow$ $\exists n\in\omega\;\;|S^{n}_{A\setminus B}|<\kappa.$ (31)
We prove that $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed. Let $A_{k}\in{\mathbb{P}}(L)$, for $k\in\omega$, where $\dots\subset_{\mathcal{I}}A_{1}\subset_{\mathcal{I}}A_{0}$. W.l.o.g. we can suppose that
$A_{0}\supset A_{1}\supset\dots,$ (32)
because ${\mathcal{I}}$ is an ideal and, hence, $B_{k}:=\bigcap_{i\leq k}A_{k}\in{\mathbb{P}}(L)$ and $B_{0}\supset B_{1}\supset\dots$; so a lower bound for $B_{k}$’s will be a lower bound for $A_{k}$’s.
Also w.l.o.g. we suppose that for each $k\in\omega$ we have $A_{k}\not\subset_{\mathcal{I}}A_{k+1}$, so, by (31), $S^{k}_{A_{k}\setminus A_{k+1}}\in[\kappa]^{\kappa}$. For $\xi\in S^{k}_{A_{k}\setminus A_{k+1}}$ we have $\omega^{\delta_{k}}\hookrightarrow L_{\xi}\cap A_{k}\setminus A_{k+1}$, and we choose sets $A^{k}_{\xi}$ such that
$\omega^{\delta_{k}}\cong A^{k}_{\xi}\subset L_{\xi}\cap A_{k}\setminus A_{k+1},\mbox{ for }\xi\in S^{k}_{A_{k}\setminus A_{k+1}}\mbox{ and }k\in\omega,$ (33)
and define
$A=\bigcup_{k\in\omega}\bigcup_{\xi\in S^{k}_{A_{k}\setminus A_{k+1}}}A^{k}_{\xi}.$ (34)
By (33) and (34), for $\xi\in S^{k}_{A_{k}\setminus A_{k+1}}$ we have $\omega^{\delta_{k}}\hookrightarrow L_{\xi}\cap A$, that is $\xi\in S^{k}_{A}$; so $S^{k}_{A_{k}\setminus A_{k+1}}\subset S^{k}_{A}$ and, hence, $|S^{k}_{A}|=\kappa$, for all $k\in\omega$, which by (30) gives $A\in{\mathbb{P}}(L)$.
We show that $A\subset_{\mathcal{I}}A_{k}$, for $k\in\omega$. By (33) we have $A^{k}_{\xi}\subset A_{k}$, for all $\xi\in S^{k}_{A_{k}\setminus A_{k+1}}$ so, by (32) and (34),
$A\setminus A_{k}\subset\bigcup_{l<k}\bigcup_{\xi\in S^{l}_{A_{l}\setminus A_{l+1}}}A^{l}_{\xi}.$ (35)
Suppose that $\omega^{\delta_{k}}\hookrightarrow L_{\xi_{0}}\cap A\setminus A_{k}$, for some $\xi_{0}<\kappa$. Then by (35) and since $A^{l}_{\xi}\subset L_{\xi}$ we have $\omega^{\delta_{k}}\hookrightarrow\bigcup_{l<k\;\land\;\xi_{0}\in S^{l}_{A_{l}\setminus A_{l+1}}}A^{l}_{\xi_{0}}=:B$. But $B$ is a union of $\leq k$ many summands $A^{l}_{\xi_{0}}\cong\omega^{\delta_{l}}$, where $l\leq k-1$ and, hence, $\mathop{\mbox{type}}\nolimits(B)\leq\omega^{\delta_{k-1}}k<\omega^{\delta_{k-1}}\omega=\omega^{\delta_{k-1}+1}\leq\omega^{\delta_{k}}$ and $\omega^{\delta_{k}}\hookrightarrow B$ is impossible. Thus for all $\xi<\kappa$ we have $\omega^{\delta_{k}}\not\hookrightarrow L_{\xi}\cap A\setminus A_{k}$, that is $S^{k}_{A\setminus A_{k}}=\emptyset$ and, by (31), $A\subset_{\mathcal{I}}A_{k}$.
Case (C) The statement follows from Theorem 4.8, for $\lambda>\omega$.
Case (D) Clearly we have $\mathop{\rm cf}\nolimits(\delta)=\kappa$ so, by Lemma 4.4(a), $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$. In addition we have $\omega^{\delta}=\omega^{\theta+\kappa}=\omega^{\theta}\kappa=\sum_{\kappa}\omega^{\theta}$ and, by Lemma 4.4(a) again, $\mathop{\rm cf}\nolimits(\omega^{\theta})=\mathop{\rm cf}\nolimits(\theta)\geq\kappa$. Thus, by Theorem 4.7, $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Case (E) By Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$. By the assumption and Lemma 4.4(a) again, for each $\xi<\kappa$ we have $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})=\mathop{\rm cf}\nolimits(\delta_{\xi})\geq\kappa$; thus, by Theorem 4.7, $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$. $\Box$
5 The posets ${\mathbb{P}}(\omega^{\delta})$: Maximal collapse
5.1 Case (A)
Successors
Here we consider a subcase of Case (A) of Theorem 4.10: the posets of the form ${\mathbb{P}}(\omega^{\delta+1})$, where $\delta$ is an ordinal. By Fact 4.3(a) we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1}))={\mathcal{P}}_{\delta+1}:=(P(\omega^{\delta+1})/{\mathcal{I}}_{\omega^{\delta+1}})^{+}$ and, since by Fact 2.12(a) we have $|\omega^{\delta+1}|=|\delta|$, it follows that $|{\mathcal{P}}_{\delta+1}|\leq|P(\omega^{\delta+1})|=2^{|\delta|}$. By Theorem 4.10 the forcing ${\mathcal{P}}_{\delta+1}$ is $\omega_{1}$-closed, thus preserves $\omega_{1}$. Since ${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi$ and the first iteration collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$, it is natural to ask whether it collapses more cardinals (which, of course, depends on $\delta$, ${\mathbb{P}}(\omega^{\delta})$, $\mathop{\rm cc}\nolimits({\mathbb{P}}(\omega^{\delta+1}))$, etc.). Here we show that if ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$ (examples of such $\delta$’s can be constructed using (C), (D), and (E) of Theorem 4.10), then ${\mathfrak{h}}(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))=\omega_{1}$ and, moreover, under some additional conditions, (for example, if ${\mathfrak{h}}=\omega_{1}$, which holds under GCH and in all Cohen extensions, or if $0^{\sharp}$ does not exist and ${\mathfrak{b}}={\mathfrak{d}}$, e.g. if MA is true) the collapse by ${\mathbb{P}}(\omega^{\delta+1})$ is maximal possible: $2^{|\delta|}$ is collapsed to $\omega_{1}$ and, moreover,
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|}).$ (36)
Fact 5.1
Let $0^{\sharp}$ does not exist let ${\mathbb{P}}$ be a separative $\omega$-distributive forcing and $V_{\mathbb{P}}[G]$ a generic extension by ${\mathbb{P}}$. Then
(a) If ${\mathfrak{b}}={\mathfrak{d}}$ and ${\mathfrak{b}}^{V_{\mathbb{P}}[G]}=\omega_{1}$, then $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$;
(b) If $\mathop{\rm cf}\nolimits^{V_{\mathbb{P}}[G]}({\mathfrak{h}})=\omega_{1}$, then $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$.
Proof. (a) Under the assumptions, in $V$ there is a dominating family ${\mathcal{D}}=\{f_{\alpha}:\alpha<{\mathfrak{b}}\}$ in ${}^{\omega}\omega$ well ordered by $\leq^{*}$ (a scale) and in $V_{\mathbb{P}}[G]$ there is an unbounded family $\{g_{\xi}:\xi<\omega_{1}\}$. For $\xi<\omega_{1}$ let $\alpha_{\xi}<{\mathfrak{b}}$, where $g_{\xi}\leq^{*}f_{\alpha_{\xi}}$ Assuming that the set $\{a_{\xi}:\xi<\omega_{1}\}$ is not cofinal in ${\mathfrak{b}}$ there would be $\alpha<{\mathfrak{b}}$ such that for each $\xi<\omega_{1}$ we have $a_{\xi}\leq\alpha$ and, hence, $g_{\xi}\leq^{*}f_{\alpha_{\xi}}\leq^{*}f_{\alpha}$, which is false. Thus the set $X:=\{a_{\xi}:\xi<\omega_{1}\}$ is cofinal in ${\mathfrak{b}}$. Since $0^{\sharp}$ does not exist in $V_{\mathbb{P}}[G]$ too, by Jensen’s Covering Lemma applied in $V_{\mathbb{P}}[G]$ there is a set $C\in L^{V_{\mathbb{P}}[G]}=L\subset V$ such that $X\subset C$ and $|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$. Clearly, the set $C\cap{\mathfrak{b}}$ is unbounded in ${\mathfrak{b}}$ and belongs to $V$ and, since ${\mathfrak{b}}$ is a regular cardinal (in $V$) we have $|C\cap{\mathfrak{b}}|^{V}={\mathfrak{b}}$. Hence we have $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}=|C\cap{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}\leq|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$ and, by the $\omega$-distributivity of ${\mathbb{P}}$, $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}\geq\omega_{1}$, which gives $|{\mathfrak{b}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$.
(b) Let $\mathop{\rm cf}\nolimits^{V_{\mathbb{P}}[G]}({\mathfrak{h}})=\omega_{1}$ and, in $V_{\mathbb{P}}[G]$, let $X:=\{a_{\xi}:\xi<\omega_{1}\}$ be a cofinal subset of ${\mathfrak{h}}$. Then, as above, there is a set $C\in L\subset V$ such that $X\subset C$ and $|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$. Again, the set $C\cap{\mathfrak{h}}\in V$ is unbounded in ${\mathfrak{h}}$ and, since ${\mathfrak{h}}$ is a regular cardinal (in $V$) we have $|C\cap{\mathfrak{h}}|^{V}={\mathfrak{h}}$. Hence we have $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}=|C\cap{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}\leq|C|^{V_{\mathbb{P}}[G]}=\omega_{1}$ and, by the $\omega$-distributivity of ${\mathbb{P}}$, $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}\geq\omega_{1}$, which gives $|{\mathfrak{h}}|^{V_{\mathbb{P}}[G]}=\omega_{1}$. $\Box$
Theorem 5.2
If $\delta\geq\omega_{1}$ is an ordinal and forcing by ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$, then we have
(a) $1_{{\mathbb{P}}(\omega^{\delta+1})}\Vdash|((2^{|\delta|})^{V})\check{\;}|=|({\mathfrak{h}}^{V})\check{\;}|$;
(b) ${\mathfrak{h}}(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))=\omega_{1}$;
(c) If $1_{{\mathbb{P}}(\omega^{\delta+1})}\Vdash|({\mathfrak{h}}^{V})\check{\;}|=\check{\omega_{1}}$ (e.g. if ${\mathfrak{h}}=\omega_{1}$), then (36) holds;
(d) If $0^{\sharp}$ does not exist and ${\mathfrak{b}}={\mathfrak{d}}$ (e.g., under MA), then (36) holds.
Proof. (a) Since $\omega^{\delta+1}=\omega^{\delta}\omega\cong L:=\sum_{i\in\omega}L_{i}$, where $L_{i}:=\{i\}\times\omega^{\delta}\cong\omega^{\delta}$, for $i\in\omega$, for convenience we identify the posets ${\mathbb{P}}(\omega^{\delta+1})$ and ${\mathbb{P}}(L)=\langle P,\subset\rangle$. By Proposition 4.6(c)
$A\in P\Leftrightarrow\mathop{\rm supp}\nolimits(A):=\{i\in\omega:\omega^{\delta}\hookrightarrow A\cap L_{i}\}\in[\omega]^{\omega}.$ (37)
If $\mathop{\rm sm}\nolimits{\mathbb{P}}(L)=\langle P,\leq\rangle$ and $A,B\in P$, then by Proposition 4.3(a) we have $A\leq B$ iff $A\setminus B\not\in P$, iff, by (37), $|\mathop{\rm supp}\nolimits(A\setminus B)|<\omega$; thus,
$A\leq B\Leftrightarrow\exists j\in\omega\;\;\forall i\geq j\;\;\omega^{\delta}\not\hookrightarrow A\setminus B\cap L_{i}.$ (38)
We recall that for a preorder ${\mathbb{P}}$ by $\mathop{\rm rp}\nolimits({\mathbb{P}})$ we denote the reduced power ${\mathbb{P}}^{\omega}/\mathop{\rm Fin}\nolimits$. Defining ${\mathcal{P}}_{\delta}:=P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$, by Fact 4.3(a) we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathcal{P}}_{\delta}$ and it is easy to check (see [14] or use (37) and (38) for a direct proof) that $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta+1})\cong\mathop{\rm rp}\nolimits({\mathcal{P}}_{\delta})^{+}$. Since ${\mathcal{P}}_{\delta}$ is a dense in $\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta})$ it is evident that $\mathop{\rm rp}\nolimits({\mathcal{P}}_{\delta})^{+}$ is a dense subset of the poset $\mathop{\rm rp}\nolimits(\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta}))^{+}$; thus
${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}\mathop{\rm rp}\nolimits(\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta}))^{+}.$ (39)
On the other hand, by Fact 2.12(a) we have $|\omega^{\delta}|=|\delta|$ and, hence, $|{\mathcal{P}}_{\delta}|\leq|P(\omega^{\delta})|=2^{|\delta|}$. Since forcing by ${\mathcal{P}}_{\delta}$ collapses $2^{|\delta|}$ to $\omega$ we have $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\delta})\geq(2^{|\delta|})^{+}$, which implies that $\pi({\mathcal{P}}_{\delta})=|{\mathcal{P}}_{\delta}|=2^{|\delta|}$. So, by Fact 2.5, $\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta})\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$, which implies that $\mathop{\rm rp}\nolimits(\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta}))^{+}\cong\mathop{\rm rp}\nolimits(\mathop{\rm Col}\nolimits(\omega,2^{|\delta|}))^{+}$; so, by (39) we have
${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}\mathop{\rm rp}\nolimits(\mathop{\rm Col}\nolimits(\omega,2^{|\delta|}))^{+}.$ (40)
Further, the reversed tree (${}^{<\omega}(2^{|\delta|})$ densely embeds in $\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})^{+}$. Adding a smallest element, say $s$, to the reversed tree (${}^{<\omega}(2^{|\delta|})$ we easily prove that $\mathop{\rm rp}\nolimits(\{s\}\cup{}^{<\omega}(2^{|\delta|}))^{+}$ densely embeds in $\mathop{\rm rp}\nolimits(\mathop{\rm Col}\nolimits(\omega,2^{|\delta|}))^{+}$. So by (40) the posets ${\mathbb{P}}(\omega^{\delta+1})$ and ${\mathbb{Q}}^{\prime}:=\mathop{\rm rp}\nolimits(\{s\}\cup{}^{<\omega}(2^{|\delta|}))^{+}$ are forcing equivalent. Finally, the poset ${\mathbb{Q}}^{\prime}$ is the separative quotient of the preorder ${\mathbb{Q}}=\langle Q,\leq_{\mathbb{Q}}\rangle$, where
$Q=\Big{\{}\psi=\langle\psi_{i}:i\in\omega\rangle\in(\{s\}\cup{}^{<\omega}(2^{|\delta|}))^{\omega}:\;\mathop{\rm supp}\nolimits(\psi):=\{i\in\omega:\psi_{i}\neq s\}\in[\omega]^{\omega}\Big{\}},$ (41)
$\psi\leq_{\mathbb{Q}}\eta\Leftrightarrow\exists j\in\omega\;\;\forall i\in\mathop{\rm supp}\nolimits(\psi)\setminus j\;\;(i\in\mathop{\rm supp}\nolimits(\eta)\land\psi_{i}\supset\eta_{i}).$ (42)
So, by Fact 2.2(a), ${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}{\mathbb{Q}}$ and we consider ${\mathbb{Q}}$. The sets $\Delta_{\xi}:=\{\psi\in Q:\forall i\in\mathop{\rm supp}\nolimits(\psi)\;\;\exists\eta_{i}\in{}^{<\omega}(2^{|\delta|})\;\;\psi_{i}=\eta_{i}{}^{\smallfrown}\xi\}$, where $\xi<2^{|\delta|}$, are dense in ${\mathbb{Q}}$: if $\eta\in Q$, defining $\psi_{i}=\eta_{i}{}^{\smallfrown}\xi$, for $i\in\mathop{\rm supp}\nolimits(\eta)$, and $\psi_{i}=s$, otherwise, we have $\Delta_{\xi}\ni\langle\psi_{i}:i\in\omega\rangle\leq_{\mathbb{Q}}\eta$.
Let $G$ be a ${\mathbb{Q}}$ generic filter over $V$ and let $\psi^{\xi}\in G\cap\Delta_{\xi}$, for $\xi<2^{|\delta|}$. Since the forcing ${\mathbb{P}}(\omega^{\delta+1})$ is $\sigma$-closed, the forcing ${\mathbb{Q}}$ does not reals and, hence, $([\omega]^{\omega})^{V[G]}=([\omega]^{\omega})^{V}$; in addition, for each $S\in([\omega]^{\omega})^{V[G]}$ we have $({}^{S}\omega)^{V[G]}=({}^{S}\omega)^{V}$. Thus, in $V[G]$, for the family ${\mathcal{S}}:=\bigcup_{S\in[\omega]^{\omega}}\{S\}\times{}^{S}\omega$ (belonging to $V$) we have $|{\mathcal{S}}|^{V[G]}=|{\mathfrak{c}}^{V}|^{V[G]}=|{\mathfrak{h}}^{V}|^{V[G]}$, because by Theorem 4.10(A) we have ${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}\langle[\omega]^{\omega},\subset\rangle\ast\pi$ and by Fact 2.8(e) the first iteration collapses ${\mathfrak{c}}^{V}$ to ${\mathfrak{h}}^{V}$.
So, for $\xi<2^{|\delta|}$ we have $\mathop{\rm supp}\nolimits(\psi^{\xi})\in([\omega]^{\omega})^{V}$ and the sequence $\langle|\psi^{\xi}_{i}|:i\in\mathop{\rm supp}\nolimits(\psi^{\xi})\rangle$ belongs to $({}^{\mathop{\rm supp}\nolimits(\psi^{\xi})}\omega)^{V}$; thus, the function $F:2^{|\delta|}\rightarrow{\mathcal{S}}$, given by $F(\xi)=\langle\mathop{\rm supp}\nolimits(\psi^{\xi}),\langle|\psi^{\xi}_{i}|:i\in\mathop{\rm supp}\nolimits(\psi^{\xi})\rangle\rangle$, for $\xi<2^{|\delta|}$, is well defined.
Assuming that $F$ is not an injection there would be $\xi<\zeta<2^{|\delta|}$ and $\langle S,\langle n_{i}:i\in S\rangle\rangle\in{\mathcal{S}}$ such that $F(\xi)=F(\zeta)=\langle S,\langle n_{i}:i\in S\rangle\rangle$, that is, $\mathop{\rm supp}\nolimits(\psi^{\xi})=\mathop{\rm supp}\nolimits(\psi^{\zeta})=S$ and $|\psi^{\xi}_{i}|=|\psi^{\zeta}_{i}|=n_{i}\in{\mathbb{N}}$, for $i\in S$. So, for $i\in S$ we would have $\psi^{\xi}_{i}=\eta_{i}{}^{\smallfrown}\xi$ and $\psi^{\zeta}_{i}=\theta_{i}{}^{\smallfrown}\zeta$, where $\eta_{i}=\psi^{\xi}_{i}\upharpoonright(n_{i}-1)$ and $\theta_{i}=\psi^{\zeta}_{i}\upharpoonright(n_{i}-1)$. Since $\psi^{\xi},\psi^{\zeta}\in G$ there is $\psi\in G$ such that $\psi\leq_{\mathbb{Q}}\psi^{\xi},\psi^{\zeta}$ and by (42) there would be $j\in\omega$ such that for each $i\in S\setminus j$ we have $\psi_{i}\supset\psi^{\xi}_{i},\psi^{\zeta}_{i}$, and, hence, $\xi=\psi^{\xi}_{i}(n_{i}-1)=\psi_{i}(n_{i}-1)=\psi^{\zeta}_{i}(n_{i}-1)=\zeta$, which is false. Thus $F$ is an injection and, hence, $|2^{|\delta|}|^{V[G]}\leq|{\mathcal{S}}|^{V[G]}=|{\mathfrak{h}}^{V}|^{V[G]}$.
(b) In order to prove that ${\mathbb{Q}}$ is not $\omega_{1}$-distributive we first show that the sets $D_{\xi}$, $\xi<2^{|\delta|}$, are dense and open in ${\mathbb{Q}}$, where
$D_{\xi}:=\{\psi\in Q:\exists j\in\omega\;\;\forall i\in\mathop{\rm supp}\nolimits(\psi)\setminus j\;\;\xi\in\mathop{\mathrm{ran}}\nolimits(\psi_{i})\}.$
If $\eta\in Q$, defining $\psi_{i}=\eta_{i}{}^{\smallfrown}\xi$, for $i\in\mathop{\rm supp}\nolimits(\eta)$, and $\psi_{i}=s$, for $i\in\omega\setminus\mathop{\rm supp}\nolimits(\eta)$, we have $D_{\xi}\ni\psi\leq_{\mathbb{Q}}\eta$; so, $D_{\xi}$ is dense in ${\mathbb{Q}}$.
If $\psi\leq_{\mathbb{Q}}\eta\in D_{\xi}$, then there is $j_{1}\in\omega$ such that for each $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j_{1}$ we have $i\in\mathop{\rm supp}\nolimits(\eta)$ and $\psi_{i}\supset\eta_{i}$ and there is $j_{2}\in\omega$ such that $\xi\in\mathop{\mathrm{ran}}\nolimits(\eta_{i})$, for all $i\in\mathop{\rm supp}\nolimits(\eta)\setminus j_{2}$; Thus, if $j=\max\{j_{1},j_{2}\}$ then for $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j$ we have $i\in\mathop{\rm supp}\nolimits(\eta)\setminus j_{2}$ and $\psi_{i}\supset\eta_{i}$, which implies $\xi\in\mathop{\mathrm{ran}}\nolimits(\eta_{i})\subset\mathop{\mathrm{ran}}\nolimits(\psi_{i})$ and, hence $\psi\in D_{\xi}$. So, $D_{\xi}$ is open in ${\mathbb{Q}}$.
Suppose that $\psi\in\bigcap_{\xi<\omega_{1}}D_{\xi}$. Then for each $\xi<\omega_{1}$ there is $j_{\xi}\in\omega$ such that for each $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j_{\xi}$ we have $\xi\in\mathop{\mathrm{ran}}\nolimits(\psi_{i})$ and, hence, there are $j^{*}\in\omega$ and $S\in[\omega_{1}]^{\omega_{1}}$ such that $j_{\xi}=j^{*}$, for all $\xi\in S$. Thus, for each $\xi\in S$ and each $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j^{*}$ we have $\xi\in\mathop{\mathrm{ran}}\nolimits(\psi_{i})$ and, hence, for an $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j^{*}$ we have $S\subset\mathop{\mathrm{ran}}\nolimits(\psi_{i})$, which is impossible, because $|\mathop{\mathrm{ran}}\nolimits(\psi_{i})|<\omega$. Thus $\bigcap_{\xi<\omega_{1}}D_{\xi}=\emptyset$ and ${\mathbb{Q}}$ is not $\omega_{1}$-distributive.
(c) Let $1_{{\mathbb{P}}(\omega^{\delta+1})}\Vdash|({\mathfrak{h}}^{V})\check{\;}|=\check{\omega_{1}}$. By Fact 4.3(a) we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1}))={\mathcal{P}}_{\delta+1}:=(P(\omega^{\delta+1})/{\mathcal{I}}_{\omega^{\delta+1}})^{+}$. By Fact 2.12(a) we have $|\omega^{\delta+1}|=|\delta|$ and, hence, $|{\mathcal{P}}_{\delta+1}|\leq|P(\omega^{\delta+1})|=2^{|\delta|}$. By Theorem 4.10 the forcing ${\mathcal{P}}_{\delta+1}$ is $\omega_{1}$-closed, thus preserves $\omega_{1}$. Since ${\mathcal{P}}_{\delta+1}$ collapses ${\mathfrak{h}}$ to $\omega_{1}$, by (a) it collapses $2^{|\delta|}$ to $\omega_{1}$ and, hence, it is $(\omega_{1},\cdot,2^{|\delta|})$-nowhere distributive. Thus we have $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\delta+1})\geq(2^{|\delta|})^{+}$, which implies that $\pi({\mathcal{P}}_{\delta+1})=|{\mathcal{P}}_{\delta+1}|=2^{|\delta|}=(2^{|\delta|})^{<\omega_{1}}$. So, by Fact 2.5, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))=\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta+1})\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$.
(d) Let $V_{\mathbb{Q}}[G]$ be a generic extension; by (b) it is sufficient to prove that $|{\mathfrak{h}}|^{V_{\mathbb{Q}}[G]}=\omega_{1}$. Suppose that $|{\mathfrak{h}}|^{V_{\mathbb{Q}}[G]}>\omega_{1}$. Then, since ${\mathfrak{h}}\leq{\mathfrak{b}}$, by Fact 5.1 we have $\mathop{\rm cf}\nolimits^{V_{\mathbb{Q}}[G]}({\mathfrak{h}})>\omega_{1}$ and ${\mathfrak{b}}^{V_{\mathbb{Q}}[G]}>\omega_{1}$. It is evident that the sets
$D_{\xi}$ ::=:= $\{\psi\in Q:\forall i\in\mathop{\rm supp}\nolimits(\psi)\;\;\xi\in\mathop{\mathrm{ran}}\nolimits(\psi_{i})\},\mbox{ for }\xi<2^{|\delta|},$
$E_{S}$ ::=:= $\{\psi\in Q:\mathop{\rm supp}\nolimits(\psi)\subset S\lor\mathop{\rm supp}\nolimits(\psi)\subset\omega\setminus S\},\mbox{ for }S\subset\omega,$
$F_{f}$ ::=:= $\{\psi\in Q:\forall i\in\mathop{\rm supp}\nolimits(\psi)\;|\psi_{i}|\geq f(i)\},\mbox{ for }f:\omega\rightarrow\omega,$
are dense in ${\mathbb{Q}}$. (E.g., if $\xi<2^{|\delta|}$ and $\eta\in Q$, defining $\psi_{i}=\eta_{i}{}^{\smallfrown}\xi$, for $i\in\mathop{\rm supp}\nolimits(\eta)$, and $\psi_{i}=s$, for $i\in\omega\setminus\mathop{\rm supp}\nolimits(\eta)$, we have $D_{\xi}\ni\psi\leq_{\mathbb{Q}}\eta$.)
By Theorem 4.10 we have ${\mathbb{P}}(\omega^{\delta+1})\equiv_{forc}\langle[\omega]^{\omega},\subset^{*}\rangle\ast\pi$; so, (see [3], p. 356) the generic filter in the first iteration is a selective ultrafilter ${\mathcal{U}}_{G}$ in $P(\omega)$ having a tower of length ${\mathfrak{h}}$ as its base, say $\{U_{\alpha}:\alpha<{\mathfrak{h}}\}$. It is evident that ${\mathcal{U}}_{G}=\{\mathop{\rm supp}\nolimits(\psi):\psi\in G\}$.
Let $\eta^{\xi}\in G\cap D_{\xi}$, for $\xi<\omega_{1}$. Then for each $\xi<\omega_{1}$ there is $\alpha_{\xi}\in{\mathfrak{h}}$ such that $U_{\alpha_{\xi}}\subset^{*}\mathop{\rm supp}\nolimits(\eta^{\xi})$. Since $\mathop{\rm cf}\nolimits^{V_{\mathbb{Q}}[G]}({\mathfrak{h}})>\omega_{1}$ we have $\beta:=\sup\{\alpha_{\xi}:\xi<\omega_{1}\}<{\mathfrak{h}}$ and $U_{\beta}\subset^{*}U_{\alpha_{\xi}}\subset^{*}\mathop{\rm supp}\nolimits(\eta^{\xi})$, for $\xi<\omega_{1}$.
Let $\eta^{U_{\beta}}\in G\cap E_{U_{\beta}}$; clearly we have $\mathop{\rm supp}\nolimits(\eta^{U_{\beta}})\subset U_{\beta}$. For $\xi<\omega_{1}$ let the function $\varphi^{\xi}:U_{\beta}\rightarrow\omega$ be defined by
$\varphi^{\xi}(i)=\left\{\begin{array}[]{cl}|\eta^{\xi}_{i}|,&\mbox{ if }i\in U_{\beta}\cap\mathop{\rm supp}\nolimits(\eta^{\xi}),\\ \emptyset,&\mbox{ if }i\in U_{\beta}\setminus\mathop{\rm supp}\nolimits(\eta^{\xi}).\end{array}\right.$
Since ${\mathfrak{b}}^{V_{\mathbb{Q}}[G]}>\omega_{1}$ there is $\varphi:U_{\beta}\rightarrow\omega$ such that $\varphi^{\xi}\leq^{*}\varphi$, for all $\xi<\omega_{1}$. Let $\varphi\subset f:\omega\rightarrow\omega$ and $\eta^{f}\in G\cap F_{f}$.
Let $\xi<\omega_{1}$. Since $\eta^{U_{\beta}},\eta^{f},\eta^{\xi}\in G$ there is $\psi\in G$ such that $\psi\leq_{Q}\eta^{U_{\beta}},\eta^{f},\eta^{\xi}$ and by (42) there is $j_{1}\in\omega$ such that
$\forall i\in\mathop{\rm supp}\nolimits(\psi)\setminus j_{1}\;\;\Big{(}i\in\mathop{\rm supp}\nolimits(\eta^{U_{\beta}})\cap\mathop{\rm supp}\nolimits(\eta^{f})\cap\mathop{\rm supp}\nolimits(\eta^{\xi})\land\eta^{U_{\beta}}_{i},\eta^{f}_{i},\eta^{\xi}_{i}\subset\psi_{i}\Big{)}.$ (43)
Since $\mathop{\rm supp}\nolimits(\psi)\subset^{*}\mathop{\rm supp}\nolimits(\eta^{U_{\beta}})\subset U_{\beta}$ and on the set $U_{\beta}$ we have $\varphi^{\xi}\leq^{*}\varphi=f\upharpoonright U_{\beta}$, there is $j_{2}\geq j_{1}$ such that
$\forall i\in\mathop{\rm supp}\nolimits(\psi)\setminus j_{2}\;\;|\eta^{\xi}_{i}|=\varphi^{\xi}(i)\leq\varphi(i)=f(i).$ (44)
Since $\eta^{f}\in F_{f}$ for each $i\in\mathop{\rm supp}\nolimits(\eta^{f})$ we have $f(i)\leq|\eta^{f}_{i}|$. Thus, by (44) for each $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j_{2}$ we have $|\eta^{\xi}_{i}|\leq|\eta^{f}_{i}|$ and, since by (43) we have $\eta^{f}_{i},\eta^{\xi}_{i}\subset\psi_{i}$, it follows that $\eta^{\xi}_{i}\subset\eta^{f}_{i}\subset\psi_{i}$. Since $\eta^{\xi}\in D_{\xi}$ there is $j_{3}\geq j_{2}$ such that for all $i\in\mathop{\rm supp}\nolimits(\psi)\setminus j_{3}$ we have $\xi\in\mathop{\mathrm{ran}}\nolimits(\eta^{\xi}_{i})\subset\mathop{\mathrm{ran}}\nolimits(\psi_{i})$. Thus $\omega_{1}\subset\bigcup_{i\in\mathop{\rm supp}\nolimits(\psi)}\mathop{\mathrm{ran}}\nolimits(\psi_{i})=:R_{\psi}$, which is impossible because the set $R_{\psi}$ is countable (and $\omega_{1}$ is not collapsed). $\Box$
Corollary 5.3
If $\kappa>\omega$ is a regular cardinal and $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}$ (in particular, if $2^{\kappa}=\kappa^{+}$), then $1_{{\mathbb{P}}(\omega^{\kappa+1})}\Vdash|((2^{\kappa})^{V})\check{\;}|=|({\mathfrak{h}}^{V})\check{\;}|$.
Proof. Under the assumptions we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))=\mathop{\rm sq}\nolimits({\mathbb{P}}(\kappa))={\mathcal{P}}_{\kappa}$ and by Fact 2.8(c) forcing by ${\mathcal{P}}_{\kappa}$ collapses $2^{\kappa}$ to $\omega$. So, by Theorem 5.2(a) we have $1_{{\mathbb{P}}(\omega^{\kappa+1})}\Vdash|((2^{\kappa})^{V})\check{\;}|=|({\mathfrak{h}}^{V})\check{\;}|$. $\Box$
Example 5.4
It is consistent that ${\mathbb{P}}(\omega^{\kappa+1})$ collapses $2^{\kappa}$ to $\omega_{1}$, although $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})<(2^{\kappa})^{+}$ and ${\mathbb{P}}(\omega^{\kappa})$ does not collapse $2^{\kappa}$ to $\omega$. First, concerning claim (c) of Theorem 5.2(c) we note that the equality ${\mathfrak{h}}=\omega_{1}$ holds in Cohen extensions of models of GCH; then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+1})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. If $V^{\prime}$ is a Cohen model for ${\mathfrak{c}}=\omega_{5}$ obtained from a model $V$ of GCH, then in $V^{\prime}$ we have ${\mathfrak{h}}=\omega_{1}$ and, by (9), $2^{\omega_{1}}=(\max(\omega_{5},2)^{\omega_{1}})^{V}=(\omega_{5}^{\omega_{1}})^{V}=\omega_{5}$. By Theorem 4.10(A) and Fact 2.8(e) we have $1_{{\mathbb{P}}(\omega^{\omega_{1}+1})}\Vdash|2^{\omega_{1}}|=\omega_{1}$ although (see [5]) $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}<\omega_{6}=(2^{\omega_{1}})^{+}$ and forcing by ${\mathbb{P}}(\omega^{\omega_{1}})$ does not collapse $2^{\omega_{1}}$ to $\omega$.
Countable cofinality
Here we consider the poset ${\mathbb{P}}(\omega^{\delta})$ in the particular case, when $\delta=\mu$ is a cardinal and $\mu>\mathop{\rm cf}\nolimits(\mu)=\omega$. As in the previous subsection we are interested when $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$ (and, hence, the collapse by ${\mathbb{P}}(\omega^{\delta})$ is maximal) and we analyze the relationship between the forcing ${\mathbb{P}}(\omega^{\delta})$ and its first iteration, equivalent to $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. By Fact 2.12 we have $\omega^{\delta}=\omega^{\mu}=\mu$; thus ${\mathbb{P}}(\omega^{\delta})={\mathbb{P}}(\mu)=[\mu]^{\mu}$ and by Fact 2.2(f) we have
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))=\mathop{\rm ro}\nolimits({\mathcal{P}}_{\mu}).$ (45)
Regarding the assumptions of Fact 2.6, the poset ${\mathcal{P}}_{\mu}$ is separative, atomless, $\omega_{1}$-closed and $|{\mathcal{P}}_{\mu}|=2^{\mu}$ (Fact 2.7(a) and Theorem 4.10) we have $1_{{\mathcal{P}}_{\mu}}\Vdash|\mu^{\omega}|=\omega_{1}$ (Fact 2.8(f)), and $(2^{\mu})^{<\omega_{1}}=2^{\mu\omega}=2^{\mu}$. Thus, since in this context forcing equivalence of posets implies isomorphism of their Boolean completions, by (45) we have
$\mu^{\omega}=2^{\mu}\Rightarrow\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu}).$ (46)
We note that, by König’s theorem,
$\mu^{+}\leq\mu^{\omega}=\mu^{\mathop{\rm cf}\nolimits(\mu)}\leq 2^{\mu},$ (47)
so, ${\mathbb{P}}(\omega^{\delta})$ collapses $\mu^{+}$ to $\omega_{1}$. Since $\omega_{1}\leq{\mathfrak{h}}\leq{\mathfrak{c}}=2^{\omega}\leq\mu^{\omega}$ and $1_{{\mathcal{P}}_{\mu}}\Vdash|\mu^{\omega}|=\omega_{1}$, ${\mathbb{P}}(\omega^{\delta})$ collapses ${\mathfrak{c}}$ (and, hence, ${\mathfrak{h}}$) to $\omega_{1}$ as well.
Theorem 5.5
If $\delta=\mu$ is a cardinal and $\mu>\mathop{\rm cf}\nolimits(\mu)=\omega$, then each of the following conditions implies that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$
1. $2^{\mu}=\mu^{+}$
2. $2^{<\mu}=\mu$
3. $\mathrm{MA}+\mu<{\mathfrak{c}}$
4. $V$ is a Cohen model for ${\mathfrak{c}}=\kappa$ obtained from a model of GCH in which $\kappa$ is regular and $\mu<\kappa$.
In case 4 we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ while this is false in cases 1 if e.g., ${\mathfrak{c}}=\omega_{1}$, and in 3.
Proof. 1. If $2^{\mu}=\mu^{+}$, then by (47) we have $\mu^{\omega}=2^{\mu}$ and if, for example, ${\mathfrak{c}}=\omega_{1}$, then by (46), $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})\neq\mathop{\rm Col}\nolimits(\omega_{1},\omega_{1})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$.
2. Since for a singular cardinal $\mu$ we have $2^{\mu}\in\{2^{<\mu}\mu^{\mathop{\rm cf}\nolimits(\mu)},(2^{<\mu})^{\mathop{\rm cf}\nolimits(2^{<\mu})}\}$, if $\mu$ is a strong limit cardinal, that is, $2^{<\mu}=\mu$, then $\mu^{\omega}=2^{\mu}$ and by (46) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$.
3. If $V\models\mathrm{MA}+{\mathfrak{c}}>\mu$, then, since $\mu<{\mathfrak{c}}$ by (47) we have ${\mathfrak{c}}=2^{\omega}\leq\mu^{\omega}\leq 2^{\mu}={\mathfrak{c}}$; so, $\mu^{\omega}=2^{\mu}$ and by (46) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$. Here the first iteration $\langle[\omega]^{\omega},\subset^{*}\rangle$ does not collapse cardinals (because ${\mathfrak{h}}={\mathfrak{c}}$) and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})\neq\mathop{\rm Col}\nolimits({\mathfrak{c}},{\mathfrak{c}})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$.
4. Let $V$ be a model of GCH and, in $V$, let $\kappa>\mu$ be a regular cardinal. If $V^{\prime}=V[G]$ is a Cohen model for ${\mathfrak{c}}=\kappa$, obtained by ${\mathbb{P}}=\mathop{\mathrm{Fn}}\nolimits(\kappa,2)$, then by (9) in $V^{\prime}$ we have $2^{\mu}=(\max(\kappa,2)^{\mu})^{V}=(\kappa^{\mu})^{V}=\kappa$ and $\mu^{\omega}=(\max(\kappa,\mu)^{\omega})^{V}=(\kappa^{\omega})^{V}=\kappa$. So, by (46) in $V^{\prime}$ we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$. Since in $V^{\prime}$ we have ${\mathfrak{h}}=\omega_{1}$ (see [11], p. 256) and, by Fact 2.8(e), forcing by $\langle[\omega]^{\omega},\subset^{*}\rangle$ collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$, and ${\mathfrak{c}}^{<\omega_{1}}={\mathfrak{c}}^{\omega}={\mathfrak{c}}$, by Fact 2.6 we have $\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)=\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits\langle[\omega]^{\omega},\subset^{*}\rangle)\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$. So, the two-step iteration is forcing equivalent to $P(\omega)/\mathop{\rm Fin}\nolimits$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})=\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$. $\Box$
Remark 5.6
By Fact 2.8(f) we have $\mathop{\rm Col}\nolimits(\omega_{1},\mu^{\omega})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\mu})\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))$; but the equality $\mathop{\rm Col}\nolimits(\omega_{1},\mu^{\omega})\cong\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta}))$ is not a theorem of ZFC. Baumgartner [5] constructed a model of ZFC in which ${\mathcal{P}}_{\mu}$ contains an adf of size $2^{\mu}>\mu^{\omega}$ and, hence $\pi(\mathop{\rm ro}\nolimits({\mathcal{P}}_{\mu}))=2^{\mu}>\mu^{\omega}=\pi(\mathop{\rm Col}\nolimits(\omega_{1},\mu^{\omega}))$.
Note that in 1 and 2 of Theorem 5.5 we have ${\mathfrak{c}}<\mu<\mu^{\omega}=2^{\mu}$ and that in 3 and 4 we have $\mu<{\mathfrak{c}}=\mu^{\omega}=2^{\mu}$. Generally, if $\mu<{\mathfrak{c}}$, then $\mu^{\omega}\leq{\mathfrak{c}}^{\omega}=2^{\omega}\leq\mu^{\omega}$ and, hence, $\mu^{\omega}={\mathfrak{c}}$. But $\mu<{\mathfrak{c}}=\mu^{\omega}<2^{\mu}$ is consistent too. Let, in $V$, GCH hold let $S:=\{\aleph_{n}:n<\omega\}$, let $E:S\rightarrow\mathop{\mathrm{Card}}\nolimits$ be the Easton index function defined by $E(\aleph_{n})=\aleph_{\omega+n+2}$, and let ${\mathbb{P}}_{E}$ be the corresponding Easton forcing. Then (see [11], p. 265) in $V[G]$ we have $2^{\aleph_{n}}=\aleph_{\omega+n+2}$, for $n\in\omega$; and $2^{\aleph_{\omega}}=\aleph_{\omega+\omega+1}$. Thus taking $\mu=\aleph_{\omega}$ we have ${\mathfrak{c}}=\aleph_{\omega+2}=\mu^{\omega}<2^{\mu}=\aleph_{\omega+\omega+1}$. Thus $\mu^{+}<\mu^{\mathop{\rm cf}\nolimits(\mu)}<2^{\mu}$.
5.2 Cases (D) and (E)
Using Fact 2.5 and Theorem 4.10 it is easy to prove
Theorem 5.7
If (D) or (E) holds, $|\delta|=\mathop{\rm cf}\nolimits(\delta)=\kappa>\omega$ and $2^{\kappa}=\kappa^{+}$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$ and, in addition,
  • (a)
    If (D) holds, then $\delta$ satisfies one of the following conditions
    • (i)
      $\delta=\kappa$
    • (ii)
      $\delta=\theta+\kappa(\gamma+m)$,
      where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma<\kappa$ and $m\geq 2$ and ($\theta=0$ or $\kappa^{2}\leq\theta<\kappa^{+}$);
    • (iii)
      $\delta=\theta+\kappa^{\gamma_{2}+m_{2}}(\gamma_{1}+m_{1})+\kappa$,
      where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma_{1}<\kappa$ and $\mathop{\rm Lim}\nolimits_{0}\ni\gamma_{2}<\kappa^{+}$ and $m_{1},m_{2}\in{\mathbb{N}}$ and ($\theta=0$ or $\kappa^{\gamma_{2}+m_{2}+1}\leq\theta<\kappa^{+}$);
    • (iv)
      $\delta=\theta+\kappa^{\gamma_{2}}(\gamma_{1}+m_{1})+\kappa$,
      where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma_{1}<\kappa$ and $m_{1}\in{\mathbb{N}}$ and $\mathop{\rm cf}\nolimits(\gamma_{2})=\kappa\leq\gamma_{2}<\kappa^{+}$ and ($\theta=0$ or $\kappa^{\gamma_{2}+1}\leq\theta<\kappa^{+}$).
  • (a)
    If (E) holds, then $\delta$ satisfies one of the following conditions
    • (i)
      $\delta=\theta+\kappa^{\gamma+m}$,
      where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma<\kappa^{+}$ and $m\in{\mathbb{N}}$ and $\gamma+m\leq 2$ and ($\theta=0$ or $\kappa^{\gamma+m}\leq\theta<\kappa^{+}$);
    • (ii)
      $\delta=\theta+\kappa^{\gamma}$,
      where $\kappa=\mathop{\rm cf}\nolimits(\gamma)\leq\gamma<\kappa^{+}$ and ($\theta=0$ or $\kappa^{\gamma}\leq\theta<\kappa^{+}$);
If $\delta=\kappa>\omega$ is a regular cardinal we have case (D) and $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))={\mathcal{P}}_{\kappa}$. If $2^{\kappa}=\kappa^{+}$, then by Theorem 5.7 $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$.
Example 5.8
It is consistent that $\kappa\in\mathop{\mathrm{Reg}}\nolimits$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$. Baumgartner [5] constructed a model of ZFC in which, for $\kappa=\omega_{1}$, $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=\kappa^{++}<2^{\kappa}$. Thus ${\mathcal{P}}_{\kappa}$ collapses $\kappa^{+}=\omega_{2}$ to $\omega$ and the cardinals $\geq\kappa^{++}$ remain cardinals in extensions by ${\mathcal{P}}_{\kappa}$. So, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$.
In fact, if $\omega_{1}\leq\kappa\in\mathop{\mathrm{Reg}}\nolimits$, then by Fact 2.8(c),
$\forall\lambda\in[\omega,\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa}))\;\;1_{{\mathcal{P}}_{\kappa}}\Vdash|\check{\lambda}|=\check{\omega}\;\mbox{ and }\;|((\kappa^{+})^{V})\check{\;}|=\check{\omega}\;\mbox{ and }\;\omega_{1}=(\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})^{V})\check{\;}.$ (48)
By (48) Fact 2.7(a), and Fact 2.6 we have
$\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}\Leftrightarrow 1_{{\mathcal{P}}_{\kappa}}\Vdash|((2^{\kappa})^{V})\check{\;}|=\check{\omega}\Leftrightarrow\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa}).$ (49)
Singular cardinals
Let $\mathop{\mathrm{Card}}\nolimits\ni\delta>\kappa=\mathop{\rm cf}\nolimits(\delta)>\omega$ (the smallest such $\delta$ is $\aleph_{\omega_{1}}$). Since $\delta$ is right indecomposable and $\kappa\in\mathop{\mathrm{Reg}}\nolimits$, cases (B), (C) and (D) would give $\delta=\kappa$, which is false. So this is Case (E) and by Theorem 4.10 we have $\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle[\delta]^{\delta},\subset^{*}\rangle$, which gives ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}{\mathcal{P}}_{\delta}\cong\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$. By Fact 2.7(b) ${\mathcal{P}}_{\delta}$ is $(\omega,\cdot,\delta^{+})$-nowhere distributive and by Fact 2.8(b)
$2^{\kappa}<\delta\Rightarrow\mathop{\rm Col}\nolimits(\omega,\delta^{+})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta}).$ (50)
If $2^{\delta}=\delta^{+}$ and $2^{\kappa}=\kappa^{+}$, then by Fact 2.9(a) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))=\mathop{\rm ro}\nolimits({\mathcal{P}}_{\delta})\cong\mathop{\rm Col}\nolimits(\omega,2^{\delta})$.
6 Ordinals of size $\omega_{1}$
Theorem 6.1
If $|\alpha|=\omega_{1}$, then $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed or $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$. More precisely, if $\alpha$ is presented in the Cantor normal form
$\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m,$ (51)
where $n,s_{1},\dots,s_{n}\in{\mathbb{N}}$, $1\leq\delta_{1}<\dots<\delta_{n}<\omega_{2}\,$ and $\,m\in\omega$, then
either $\forall i\leq n\;\Big{(}\mathop{\rm cf}\nolimits(\delta_{i})\leq\,\omega\,\;\lor\;\exists\theta\geq\omega_{1}\;(\mathop{\rm cf}\nolimits(\theta)=\omega\land\delta_{i}=\theta+\omega_{1})\Big{)},$ (52)
or $\exists i\leq n\;\Big{(}\mathop{\rm cf}\nolimits(\delta_{i})=\omega_{1}\;\land\;\Big{(}\exists\theta\geq\omega_{1}\;(\delta_{i}=\theta+\omega_{1})\Rightarrow\mathop{\rm cf}\nolimits(\theta)=\omega_{1}\Big{)}\Big{)}.$ (53)
(a) $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed and $((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})^{\sum_{i=1}^{n}s_{i}}\hookrightarrow\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ iff (52) holds;
(b) $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$ iff (53) holds;
(c) If (53) holds and $\mathop{\rm cc}\nolimits(P(\omega_{1})/[\omega_{1}]^{<\omega_{1}})=(2^{\omega_{1}})^{+}$, then ${\mathbb{P}}(\alpha)\equiv\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$.
Proof. By Lemma 4.1, either (52) or (53) is true.
By Lemma 3.1, w.l.o.g. we suppose that $m=0$. Then defining $k:=\sum_{i\leq n}s_{i}$ we have $\alpha\cong\sum_{1\leq j\leq k}L_{j}:=L$, where $L_{j}\cong\omega^{\delta^{\prime}_{j}}$ and
$\langle\delta^{\prime}_{j}:1\leq j\leq k\rangle=\langle\delta_{n},\dots,\delta_{n},\delta_{n-1},\dots,\delta_{n-1},\dots,\delta_{1},\dots,\delta_{1}\rangle.$
($\delta_{i}$ appears $s_{i}$ many times). By Theorem 3.2 ${\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ and by Fact 2.2(d) and (e) $\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$.
(a) If (52) holds, then by Theorem 4.10 for each $j\in\{1,\dots,k\}$ the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ is $\sigma$-closed and $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\hookrightarrow\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$. By Facts 2.2(e) and 2.1, the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ is $\sigma$-closed.
(b) If (53) holds and $i_{0}<n$ witnesses that, then, by Theorem 4.10, $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta_{i_{0}}})\hookrightarrow_{c}\prod_{1\leq j\leq k}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$.
(c) By (b) we have $\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)$, which by Fact 2.3 implies that $\mathop{\rm sm}\nolimits{\mathbb{P}}(\alpha)\equiv\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\ast\pi$. So, since $\mathop{\rm cc}\nolimits(P(\omega_{1})/[\omega_{1}]^{<\omega_{1}})=(2^{\omega_{1}})^{+}$, by Facts 2.2 and 2.8(a) and (c) the forcing $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ collapses $2^{\omega_{1}}$ to $\omega$ and, hence, $|\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)|=2^{\omega_{1}}$. Now, by Fact 2.6, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\equiv\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$ and by Fact 2.2(a) we have ${\mathbb{P}}(\alpha)\equiv\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$. $\Box$
Example 6.2
If ${\mathfrak{h}}=\omega_{1}<\omega_{2}=2^{\omega}=2^{\omega_{1}}$ and $|\delta|=\omega_{1}$, then
${\mathbb{P}}(\omega^{\delta})\equiv\left\{\begin{array}[]{cc}\mathop{\rm Col}\nolimits(\omega_{1},\omega_{2})\equiv(P(\omega)/\mathop{\rm Fin}\nolimits)^{+},&\mbox{ if (A) or (B);}\\[2.84526pt] \mathop{\rm Col}\nolimits(\omega,\omega_{2}),&\mbox{ if (D) or (E).}\end{array}\right.$
By Lemma 4.1 we have the following cases ((C) is impossible):
Case (A): $\mathop{\rm cf}\nolimits\delta\leq\omega$. Then by Theorem 4.10 and Fact 2.3 we have ${\mathbb{P}}(\omega^{\delta})\equiv\langle[\omega]^{\omega},\subset^{*}\rangle\ast\pi$. So, by Fact 2.8(e), ${\mathbb{P}}(\omega^{\delta})$ and, hence, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses $\omega_{2}$ to $\omega_{1}$. By Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\omega_{1}$-closed and, by our assumption, $\omega_{2}^{<\omega_{1}}=\omega_{2}$ so, by Fact 2.6, ${\mathbb{P}}(\omega^{\delta})\equiv\mathop{\rm Col}\nolimits(\omega_{1},\omega_{2})\equiv(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$.
Case (B): $\delta=\theta+\omega_{1}$, where $\mathop{\rm cf}\nolimits\theta=\omega<\omega_{1}<\theta<\omega_{2}$. Then by Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\omega_{1}$-closed and ${\mathbb{P}}(\omega^{\delta})\equiv\langle[\omega]^{\omega},\subset^{*}\rangle\ast\pi$. So, as in Case (A) we have ${\mathbb{P}}(\omega^{\delta})\equiv\mathop{\rm Col}\nolimits(\omega_{1},\omega_{2})\equiv(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$.
Case (D) or (E). Then from Theorem 4.10 and Fact 2.3 it follows that ${\mathbb{P}}(\omega^{\delta})\equiv\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\ast\pi$. So, by Fact 2.8(c), ${\mathbb{P}}(\omega^{\delta})$ and, hence, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses $\omega_{2}$ to $\omega$. By Fact 2.6 we have ${\mathbb{P}}(\omega^{\delta})\equiv\mathop{\rm Col}\nolimits(\omega,\omega_{2})$.
Acknowledgement.
This research was supported by the Science Fund of the Republic of Serbia, Program IDEAS, Grant No. 7750027: Set-theoretic, model-theoretic and Ramsey-theoretic phenomena in mathematical structures: similarity and diversity–SMART.