### v2

Abstract
The set of copies ${\mathbb{P}}(\alpha):=\{f[\alpha]:f\in\mathop{\rm Emb}\nolimits(\alpha)\}$ of an ordinal $\alpha$ reconstructs the monoid $\mathop{\rm Emb}\nolimits(\alpha)$ of self embeddings of $\alpha$ and $\langle{\mathbb{P}}(\alpha),\subset\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\succcurlyeq^{R}\rangle$, where $\preccurlyeq^{R}$ is right Green’s preorder. Regarding the classification of countable ordinals up to isomorphism of the Boolean completions of their posets of copies we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin})^{+}\ast\pi)$, where $\pi$ is a $P(\omega)/\mathop{\rm Fin}$-name for a $\sigma$-closed separative poset; in particular, under ${\mathfrak{h}}=\omega_{1}$ we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin})$, whenever $\alpha<\omega_{1}$.
Here we consider ordinals $\alpha\geq\omega_{1}$ and show that the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is either $\sigma$-closed and completely embeds a finite power of $(P(\omega)/\mathop{\rm Fin})^{+}$, or completely embeds $(P(\kappa)/[\kappa]^{<\kappa})^{+}$, for some regular cardinal $\kappa>\omega$ and collapses $\omega_{2}$ to $\omega$. It is consistent that the algebra $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))$ is isomorphic to $\mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}})$ or to $\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$, for each infinite ordinal $\alpha<{\mathfrak{c}}=\omega_{2}$. More generally, the partial order $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is isomorphic to a forcing product of the algebras of the form $P(\omega^{\delta})/{\mathcal{I}}_{\delta}$ and we detect several classes of ordinals $\delta$ such that the algebra $\mathop{\rm ro}\nolimits(P(\omega^{\delta})/{\mathcal{I}}_{\delta})$ is isomorphic to a collapsing algebra.
2020 MSC: 06A05, 06A10, 03E40, 03E35.
Keywords: uncountable ordinal, poset of copies, $\sigma$-closed poset, cardinal collapse, forcing.
FORCING WITH COPIES OF UNCOUNTABLE ORDINALS
Miloš S. Kurilić [^1] Department of Mathematics and Informatics, Faculty of Sciences, University of Novi Sad, Trg Dositeja Obradovića 4, 21000 Novi Sad, Serbia. e-mail: milos@dmi.uns.ac.rs
1 Introduction
In this paper we continue the investigation of the posets of copies of ordinals initiated in [10]. Generally speaking, if ${\mathbb{X}}$ is a relational structure and $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is the set of its self-embeddings, then the partial order $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$, where ${\mathbb{P}}({\mathbb{X}}):=\{f[X]:f\in\mathop{\rm Emb}\nolimits({\mathbb{X}})\}$, is called the poset of copies of ${\mathbb{X}}$. Denoting by $\mathop{\rm sq}\nolimits({\mathbb{P}})$ and $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))$ the separative quotient and the Boolean completion of a partial order ${\mathbb{P}}$, the implications
${\mathbb{X}}\cong{\mathbb{Y}}\Rightarrow{\mathbb{P}}({\mathbb{X}})\cong{\mathbb{P}}({\mathbb{Y}})\Rightarrow\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})\Rightarrow\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})$
express the relationship between “similarity relations” in the class of relational structures and each of them provides a classification of structures [8, 11]. The last one, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{Y}})$, gives the coarsest classification, it is equivalent to the forcing equivalence of posets of copies, and, hence, it can be explored using forcing-theoretic methods. The related results can be regarded as the statements belonging to the theory of relations, set theory or the theory of Boolean algebras.
In addition, these results can be applied in investigation of the transformation monoids $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),\circ,\mathop{\rm id}\nolimits_{X}\rangle$. First we note that the set $\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})$ of automorphisms of ${\mathbb{X}}$ together with a choice of an embedding $f_{A}\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$ such that $f_{A}[X]=A$, for each $A\in{\mathbb{P}}({\mathbb{X}})$, determine the set of embeddings: $\mathop{\rm Emb}\nolimits({\mathbb{X}})=\{f_{A}\circ a:A\in{\mathbb{P}}({\mathbb{X}})\mbox{ and }a\in\mathop{\mathrm{Aut}}\nolimits({\mathbb{X}})\}$. Second, we recall that right Green’s preorder on $\mathop{\rm Emb}\nolimits({\mathbb{X}})$ is defined by $f\preccurlyeq^{R}g$ iff $f\circ h=g$, for some $h\in\mathop{\rm Emb}\nolimits({\mathbb{X}})$, and note that, defining $f\thickapprox^{R}g$ iff $f\preccurlyeq^{R}g$ and $g\preccurlyeq^{R}f$, we have $f\preccurlyeq^{R}g\Leftrightarrow g[X]\subset f[X]$ and $f\thickapprox^{R}g\Leftrightarrow g[X]=f[X]$. So, $\langle{\mathbb{P}}({\mathbb{X}}),\subset\rangle$ is the antisymmetric quotient of the preorder $\langle\mathop{\rm Emb}\nolimits({\mathbb{X}}),\succcurlyeq^{R}\rangle$. If, in particular, ${\mathbb{X}}$ is an ordinal $\alpha$, it is a rigid structure and for each copy $A\in{\mathbb{P}}(\alpha)$ there is a unique $f_{A}\in\mathop{\rm Emb}\nolimits(\alpha)$ such that $f_{A}[\alpha]=A$. Thus, here we have $\mathop{\rm Emb}\nolimits(\alpha)=\{f_{A}:A\in{\mathbb{P}}(\alpha)\}$ and defining the binary operation $\ast$ on ${\mathbb{P}}(\alpha)$ by $A\ast B=f_{A}[B]$ and the mapping $F:{\mathbb{P}}(\alpha)\rightarrow\mathop{\rm Emb}\nolimits(\alpha)$ by $F(A)=f_{A}$, we obtain an isomorphism of monoids; namely, $F:\langle{\mathbb{P}}(\alpha),\ast,\alpha\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\circ,\mathop{\rm id}\nolimits_{\alpha}\rangle$. In addition, the preorder $\preccurlyeq^{R}$ on $\mathop{\rm Emb}\nolimits(\alpha)$ becomes a partial order and the same mapping $F$ witnesses that $\langle{\mathbb{P}}(\alpha),\subset\rangle\cong\langle\mathop{\rm Emb}\nolimits(\alpha),\succcurlyeq^{R}\rangle$.
It turns out that, regarding countable linear orders, the main roles in the description of the algebra $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})$ are played by the poset ${\mathbb{S}}$ of perfect subsets of the real line ordered by inclusion (the Sacks forcing) and the algebra $P(\omega)/\mathop{\rm Fin}\nolimits$. Namely, if ${\mathbb{X}}$ is non-scattered (i.e., ${\mathbb{Q}}\hookrightarrow{\mathbb{X}}$), then $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}}\ast\pi)$, where $\pi$ is an ${\mathbb{S}}$-name for a $\sigma$-closed poset and, consistently, the name for $P(\omega)/\mathop{\rm Fin}\nolimits$ in the Sacks extension [14]. On the other hand, if ${\mathbb{X}}$ is a scattered linear order, then the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})$ is $\sigma$-closed, atomless and, consistently, $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{X}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ [9]. We note that for the real line ${\mathbb{R}}$ we have $\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits{\mathbb{P}}({\mathbb{R}})\cong\mathop{\rm ro}\nolimits\mathop{\rm sq}\nolimits({\mathbb{S}})$ [12] and that the Sacks forcing appears as a factor in the presentation of posets of copies of several ultrahomogeneous structures, [5, 14, 15, 16, 17, 18].
In [10] the posets of copies of countable ordinals $\alpha\geq\omega$ were investigated and, in order to present the main results, we explain some notation. If ${\mathbb{B}}$ is Boolean lattice, $\mathop{\rm rp}\nolimits({\mathbb{B}})$ denotes the reduced power ${\mathbb{B}}^{\omega}/\mathop{\rm Fin}\nolimits$, and the iterated reduced powers $\mathop{\rm rp}\nolimits^{n}({\mathbb{B}})$, for $n\in\omega$, are defined by $\mathop{\rm rp}\nolimits^{0}({\mathbb{B}})={\mathbb{B}}$ and $\mathop{\rm rp}\nolimits^{n+1}({\mathbb{B}})=\mathop{\rm rp}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))$. If $\delta>0$ is an ordinal, then ${\mathcal{I}}_{\omega^{\delta}}:=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}\subset P(\omega^{\delta})$ is an ideal and the Boolean algebra $P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$ is the corresponding quotient. The first result shows that the separative quotient of the poset ${\mathbb{P}}(\alpha)$ is isomorphic to the forcing product of finitely many Boolean algebras: if $\alpha=\omega^{\gamma_{n}+r_{n}}s_{n}+\dots+\omega^{\gamma_{0}+r_{0}}s_{0}+k$ is a countable ordinal presented in the Cantor normal form, where $k\in\omega$, $r_{i}\in\omega$, $s_{i}\in{\mathbb{N}}$, $\gamma_{i}\in\mathop{\rm Lim}\nolimits\cup\{1\}$ and $\gamma_{n}+r_{n}>\dots>\gamma_{0}+r_{0}$, then
$\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))\cong\prod_{i=0}^{n}\Big{(}\Big{(}\mathop{\rm rp}\nolimits^{r_{i}}(P(\omega^{\gamma_{i}})/{\mathcal{I}}_{\omega^{\gamma_{i}}})\Big{)}^{+}\Big{)}^{s_{i}}.$ (1)
The second result is related to the aforementioned coarse classification of ordinals and their embedding monoids. We recall that the distributivity number ${\mathfrak{h}}({\mathbb{P}})$ of a partial order ${\mathbb{P}}$ is the minimal size of a family ${\mathcal{D}}$ of open dense subsets of ${\mathbb{P}}$ such that $\bigcap{\mathcal{D}}$ is not dense, and that $\omega_{1}\leq{\mathfrak{h}}:={\mathfrak{h}}((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})\leq{\mathfrak{c}}$.
Theorem 1.1
If $\alpha$ is a countable ordinal, then
(a) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+}\ast\pi)$, where $[\omega]\Vdash$$\pi$ is a $\sigma$-closed separative poset”;
(b) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm ro}\nolimits((P(\omega)/\mathop{\rm Fin}\nolimits)^{+})$, if $\,{\mathfrak{h}}=\omega_{1}$.
[…]
In this paper we investigate the posets ${\mathbb{P}}(\alpha)$ for uncountable ordinals $\alpha$. Section 2 contains definitions and known facts used in the paper. In Section 3 we show that the separative quotient $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is isomorphic to the direct product of $\sum_{i=1}^{n}s_{i}$-many factors $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta_{i}})$, where $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is the Cantor normal form for $\alpha$, and reduce the analysis to the investigation of these factors.
So, in Section 4 we regard the posets of the form ${\mathbb{P}}(\omega^{\delta})$. It turns out that the behavior of ${\mathbb{P}}(\omega^{\delta})$ depends on the cofinality $\kappa:=\mathop{\rm cf}\nolimits(\delta)$. If $\kappa\leq\omega$ (Case (A)), then the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed, completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$ and, hence, collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$. Otherwise we have $\kappa\geq\omega_{1}$ and, presenting $\delta$ in its Cantor normal form in the base $\kappa$, we detect additional four cases, (B)–(E). If (B) holds, then again, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and completely embeds $(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. In Cases (D) and (E) the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds the poset ${\mathcal{P}}_{\kappa}:=(P(\kappa)/[\kappa]^{<\kappa})^{+}$ and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ (and, hence, $\omega_{2}$) to $\omega$. Finally, in Case (C) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ completely embeds ${\mathcal{P}}_{\lambda}$, where $\omega<\lambda<\kappa$, and collapses (at least) all cardinals $\mu<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})$ (and, again, $\omega_{2}$) to $\omega$.
Thus, by the (ZFC) results of Section 4, the poset ${\mathbb{P}}(\omega^{\delta})$ either preserves $\omega_{1}$ and forces $|{\mathfrak{c}}|=|{\mathfrak{h}}|$, or collapses (at least) $\omega_{2}$ to $\omega$. So, it is natural to ask whether ${\mathbb{P}}(\omega^{\delta})$ collapses more cardinals and a more ambitious task is to characterize the algebra $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))$. Since $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|=|P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}|\leq 2^{|\delta|}$, the extreme case is when ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega_{1}$, in cases (A) and (B), or collapses $2^{|\delta|}$ to $\omega$, in other cases. So, in Section 5 we consider that phenomenon of “maximal collapse”. For example, in cases (D) and (E) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$, if $2^{\mathop{\rm cf}\nolimits(\delta)}=2^{|\delta|}$ and ($2^{<\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)$ or $2^{\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)^{+}$). Regarding case (A) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$, for all $n\in{\mathbb{N}}$, whenever ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$, or $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and collapses $2^{|\delta|}$ to $\omega_{1}$.
2 Preliminaries
Partial orders and forcing
If ${\mathbb{P}}=\langle P,\leq\rangle$ is a preorder, the elements $p$ and $q$ of $P$ are incompatible, we write $p\perp q$, iff there is no $r\in P$ such that $r\leq p,q$. A set $A\subset P$ is an antichain iff $p\perp q$, for different $p,q\in A$. An element $p$ of $P$ is an atom iff each two elements $q,r\leq p$ are compatible. ${\mathbb{P}}$ is called: atomless iff it has no atoms; homogeneous iff it has a largest element and ${\mathbb{P}}\cong(\cdot,p]$, for each $p\in{\mathbb{P}}$. A set $D\subset P$ is called dense iff for each $p\in P$ there is $q\in D$ such that $q\leq p$; $D$ is called open iff $q\leq p\in D$ implies $q\in D$. If $\kappa$ is a cardinal, ${\mathbb{P}}$ is called $\kappa$-closed iff for each $\gamma<\kappa$ each sequence $\langle p_{\alpha}:\alpha<\gamma\rangle$ in $P$, such that $\alpha<\beta\Rightarrow p_{\beta}\leq p_{\alpha}$, has a lower bound in $P$. $\omega_{1}$-closed preorders are called $\sigma$-closed.
A preorder ${\mathbb{P}}=\langle P,\leq\rangle$ is called separative iff for each $p,q\in P$ satisfying $p\not\leq q$ there is $r\leq p$ such that $r\perp q$. The separative modification of ${\mathbb{P}}$ is the separative preorder $\mathop{\rm sm}\nolimits({\mathbb{P}})=\langle P,\leq^{*}\rangle$, where $p\leq^{*}q\Leftrightarrow\forall r\leq p\;\exists s\leq r\;s\leq q$. The separative quotient of ${\mathbb{P}}$ is the separative partial order $\mathop{\rm sq}\nolimits({\mathbb{P}})=\mathop{\rm asq}\nolimits(\mathop{\rm sm}\nolimits({\mathbb{P}}))$, where $\mathop{\rm asq}\nolimits({\mathbb{P}})$ denotes the antisymmetric quotient of a preorder ${\mathbb{P}}$. Preorders ${\mathbb{P}}$ and ${\mathbb{Q}}$ are called forcing equivalent, in notation ${\mathbb{P}}\equiv_{forc}{\mathbb{Q}}$, iff they produce the same generic extensions.
Fact 2.1
(Folklore) If ${\mathbb{P}}$ and ${\mathbb{P}}_{i}$, for $i\in I$, are preorders, then
(a) ${\mathbb{P}}\equiv_{forc}\mathop{\rm sm}\nolimits({\mathbb{P}})\equiv_{forc}\mathop{\rm sq}\nolimits({\mathbb{P}})\equiv_{forc}\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))$;
(b) If ${\mathbb{P}}_{i}$, $i\in I$, are $\kappa$-closed preorders, then $\prod_{i\in I}{\mathbb{P}}_{i}$ is $\kappa$-closed;
(c) $\mathop{\rm sm}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})=\prod_{i\in I}\mathop{\rm sm}\nolimits{\mathbb{P}}_{i}$ and $\mathop{\rm sq}\nolimits(\prod_{i\in I}{\mathbb{P}}_{i})\cong\prod_{i\in I}\mathop{\rm sq}\nolimits{\mathbb{P}}_{i}$;
(d) If $X$ is an infinite set, ${\mathcal{I}}\subset P(X)$ is an ideal containing $[X]^{<\omega}$, ${\mathcal{I}}^{+}=P(X)\setminus{\mathcal{I}}$ is the corresponding co-ideal and $A\subset_{\mathcal{I}}B$ iff $A\setminus B\in{\mathcal{I}}$, for $A,B\in{\mathcal{I}}^{+}$, then $\mathop{\rm sm}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle$ $=\langle{\mathcal{I}}^{+},\subset_{\mathcal{I}}\rangle$ and $\mathop{\rm sq}\nolimits\langle{\mathcal{I}}^{+},\subset\rangle=(P(X)/{\mathcal{I}})^{+}$.
If $\langle{\mathbb{P}},\leq_{\mathbb{P}},1_{\mathbb{P}}\rangle$ and $\langle{\mathbb{Q}},\leq_{\mathbb{Q}},1_{\mathbb{Q}}\rangle$ are preorders, then a mapping $f:{\mathbb{P}}\rightarrow{\mathbb{Q}}$ is a complete embedding, in notation $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$ iff
(ce1) $p_{1}\leq_{\mathbb{P}}p_{2}\Rightarrow f(p_{1})\leq_{\mathbb{Q}}f(p_{2})$,
(ce2) $p_{1}\perp_{\mathbb{P}}p_{2}\Leftrightarrow f(p_{1})\perp_{\mathbb{Q}}f(p_{2})$,
(ce3) $\forall q\in{\mathbb{Q}}\;\exists p\in{\mathbb{P}}\;\forall p^{\prime}\leq_{\mathbb{P}}p\;\;f(p^{\prime})\not\perp_{\mathbb{Q}}q$.
Then, for $q\in{\mathbb{Q}}$, $\mathop{\rm red}\nolimits_{\mathbb{P}}(q)=\{p\in{\mathbb{P}}:\forall p^{\prime}\leq_{\mathbb{P}}p\;\;\exists q^{\prime}\in{\mathbb{Q}}\;\;q^{\prime}\leq f(p^{\prime}),q\}$ is the set of reductions of $q$ to ${\mathbb{P}}$. The following fact is folklore (see [7]).
Fact 2.2
If $f:{\mathbb{P}}\hookrightarrow_{c}{\mathbb{Q}}$, then ${\mathbb{Q}}$ is forcing equivalent to the two step iteration ${\mathbb{P}}\ast\pi$, where $\pi$ is a ${\mathbb{P}}$-name for a preorder. In addition, $F:\mathop{\rm asq}\nolimits({\mathbb{P}})\hookrightarrow_{c}\mathop{\rm asq}\nolimits({\mathbb{Q}})$, where $F([p]_{\mathop{\rm asq}\nolimits({\mathbb{P}})})=[f(p)]_{\mathop{\rm asq}\nolimits({\mathbb{Q}})}$, for $p\in{\mathbb{P}}$.
Fact 2.3
If ${\mathbb{P}}_{i}\hookrightarrow_{c}{\mathbb{P}}_{i}^{\prime}$, for $i\in I$, then $\prod_{i\in I}{\mathbb{P}}_{i}\hookrightarrow_{c}\prod_{i\in I}{\mathbb{P}}_{i}^{\prime}$.
Fact 2.4
If $V\models\mathrm{GCH}$ and, in $V$, $\kappa>\omega$ is a regular cardinal and ${\mathbb{P}}=\mathop{\mathrm{Fn}}\nolimits(\kappa,2)$ is the Cohen forcing, then in $V[G]$ we have ${\mathfrak{c}}=\kappa$. If, in $V$, $\mu$ and $\theta$ are infinite cardinals, then $(\theta^{\mu})^{V[G]}=(\max(\kappa,\theta)^{\mu})^{V}$ (see [7], p. 246).
Collapsing algebras and the algebras $P(\kappa)/[\kappa]^{<\kappa}$
Let $\kappa\geq 2$ and $\lambda\geq\omega$ be cardinals. The collapsing algebra $\mathop{\rm Col}\nolimits(\lambda,\kappa)$ is the Boolean completion of the reversed tree $\langle{}^{<\lambda}\kappa,\supset\rangle$. The following fact is folklore; see [13], Thm 3.1 for a proof.
Fact 2.5
Let $\lambda\geq\omega$ be a regular cardinal and ${\mathbb{P}}$ a separative $\lambda$-closed preorder of size $\kappa=\kappa^{<\lambda}$.
(a) If $\kappa>\lambda$ and $1_{\mathbb{P}}\Vdash|\check{\kappa}|=\check{\lambda}$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\cong\mathop{\rm Col}\nolimits(\lambda,\kappa)$;
(b) If $\kappa=\lambda$ and ${\mathbb{P}}$ is atomless, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}))\cong\mathop{\rm Col}\nolimits(\lambda,\lambda)$.
For a cardinal $\kappa\geq\omega$ the partial order $(P(\kappa)/[\kappa]^{<\kappa})^{+}$ is denoted by ${\mathcal{P}}_{\kappa}$; thus, ${\mathcal{P}}_{\omega}=(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$. If $\subset^{*}$ is the preorder on the set $[\kappa]^{\kappa}$ given by $A\subset^{*}B$ iff $|A\setminus B|<\kappa$, then, clearly, $\mathop{\rm sm}\nolimits\langle[\kappa]^{\kappa},\subset\rangle=\langle[\kappa]^{\kappa},\subset^{*}\rangle$ and $\mathop{\rm sq}\nolimits\langle[\kappa]^{\kappa},\subset\rangle={\mathcal{P}}_{\kappa}$.
Fact 2.6
If $\kappa$ is an infinite cardinal, then for the partial order ${\mathcal{P}}_{\kappa}$ we have
(a) ${\mathcal{P}}_{\kappa}$ is atomless, homogeneous, of size $2^{\kappa}$ and $\kappa^{++}\leq\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})\leq(2^{\kappa})^{+}$;
(b) Forcing by ${\mathcal{P}}_{\omega}$ collapses ${\mathfrak{c}}$ to ${\mathfrak{h}}$ (Balcar, Pelant and Simon, [1]);
(c) If $\kappa>\mathop{\rm cf}\nolimits(\kappa)=\omega$, then $\mathop{\rm Col}\nolimits(\omega_{1},\kappa^{\omega})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$ (Kojman, Shelah [6]);
(d) If $\kappa>2^{\mathop{\rm cf}\nolimits(\kappa)}>\mathop{\rm cf}\nolimits(\kappa)>\omega$, then $\mathop{\rm Col}\nolimits(\omega,\kappa^{+})\hookrightarrow_{c}\mathop{\rm ro}\nolimits({\mathcal{P}}_{\kappa})$ (Shelah, [20]);
(e) If $\kappa=\mathop{\rm cf}\nolimits(\kappa)>\omega$, then ${\mathcal{P}}_{\kappa}$ collapses each $\lambda<\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})$ to $\omega$ (Shelah, [20]).
Linear orders. Indecomposable ordinals
A linear order $\langle L,<\rangle$ is called dense iff for each $x,y\in L$ satisfying $x<y$ there is $z\in L$ such that $x<z<y$. $\langle L,<\rangle$ is said to be scattered iff it does not contain a dense suborder iff ↪̸${\mathbb{Q}}\not\hookrightarrow L$.
Fact 2.7
A linear order $L$ satisfying $L+L\hookrightarrow L$ is not scattered ([19], p. 176).
Fact 2.8
For a limit ordinal $\alpha$ the following conditions are equivalent:
(a) $\beta+\gamma<\alpha$, for each $\beta,\gamma<\alpha$, (that is, $\alpha$ is indecomposable),
(b) $\beta+\gamma=\alpha\land\gamma>0\Rightarrow\gamma=\alpha$ (that is, $\alpha$ is right indecomposable),
(c) $\,\beta+\alpha=\alpha$, for each $\beta<\alpha$,
(d) $\alpha=\omega^{\delta}$, for some ordinal $\delta>0$,
(e) $\alpha$ is an indivisible structure ($\alpha\hookrightarrow A$ or $\alpha\hookrightarrow B$, whenever $\alpha=A\;\dot{\cup}\;B$),
(f) ${\mathcal{I}}_{\alpha}=\{I\subset\alpha:\alpha\not\hookrightarrow I\}$ is an ideal in $P(\alpha)$.
Proof. For (a) $\Leftrightarrow$ (d) $\Leftrightarrow$ (c) see [21], p. 282, 323. For (b) $\Leftrightarrow$ (d) see [19], p. 176. The equivalence (a) $\Leftrightarrow$ (e) is 6.8.1 of [4]. (e) $\Leftrightarrow$ (f) is evident. $\Box$
Fact 2.9
(a) $\omega^{\delta}\geq\delta$, for each ordinal $\delta$;
(b) $|\omega^{\delta}|=\omega|\delta|$, for each $\delta\geq 1$;
(c) If $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$ is an uncountable ordinal, then $|\alpha|=|\delta_{n}|$.
Proof. (a) First we have $\omega^{0}=1\geq 0$. Suppose that $\omega^{\delta}\geq\delta$, for each ordinal $\delta<\theta$. If $\theta=\delta+1$, then $\omega^{\theta}=\omega^{\delta}\omega>\omega^{\delta}1=\omega^{\delta}\geq\delta$ and, hence, $\omega^{\theta}\geq\delta+1=\theta$. If $\theta$ is a limit ordinal, then $\omega^{\theta}=\sup\{\omega^{\delta}:\delta<\theta\}\geq\sup\{\delta:\delta<\theta\}=\theta$.
(b) If $\delta\geq 1$, then $|\omega^{\delta}|\geq\omega$ and, by (a), $|\omega^{\delta}|\geq|\delta|$, so $|\omega^{\delta}|\geq\omega|\delta|$. Let
$\forall\delta<\theta\;\;|\omega^{\delta}|=\omega|\delta|.$ (2)
If $\theta=\delta+1$, then $|\omega^{\theta}|=|\omega^{\delta}\omega|=|\omega^{\delta}|\omega=|\delta|\omega\omega=|\delta+1|\omega=|\theta|\omega$.
If $\theta$ is a limit ordinal, then $\omega^{\theta}=\sup\{\omega^{\delta}:\delta<\theta\}$ and by (2), for each $\delta<\theta$ we have $|\omega^{\delta}|=\omega|\delta|\leq\omega|\theta|=|\theta|$. So the ordinal $\omega^{\theta}$ is an union of $|\theta|$-many ordinals of size $\leq|\theta|$ and, hence, $|\omega^{\theta}|=|\theta|=\omega|\theta|$.
(c) $\alpha$ is a finite sum of summands of the form $\omega^{\delta_{i}}$ and by (b) we have $|\omega^{\delta_{i}}|=\omega|\delta_{i}|$. So $\alpha\geq\omega_{1}$ implies that the largest summand, $\omega^{\delta_{n}}$ is uncountable and, hence, $|\delta_{n}|\geq\omega_{1}$. Since $i<n$ implies $\delta_{i}<\delta_{n}$ we have $\omega^{\delta_{i}}<\omega^{\delta_{n}}$, so $|\omega^{\delta_{i}}|\leq|\omega^{\delta_{n}}|$. Thus by (b) we have $|\alpha|=|\omega^{\delta_{n}}|=\omega|\delta_{n}|=|\delta_{n}|$. $\Box$
3 Factorization of ${\mathbb{P}}(\alpha)$
Here we start with analysis of the posets of the form ${\mathbb{P}}(\alpha)$, where $\alpha$ is an uncountable ordinal. Clearly, these posets are homogeneous and atomless.
Lemma 3.1
$\langle{\mathbb{P}}(\gamma+m),\subset\rangle\cong\langle{\mathbb{P}}(\gamma),\subset\rangle$, for each limit ordinal $\gamma$ and each $m\in{\mathbb{N}}$.
Proof. First we prove ${\mathbb{P}}(\gamma+m)=\{C\cup\{\gamma,\gamma+1,\dots,\gamma+m-1\}:C\in{\mathbb{P}}(\gamma)\}$. The inclusion $\supset$” is evident. If $A\in{\mathbb{P}}(\gamma+m)$ and $f:\gamma+m\hookrightarrow\gamma+m$, where $A=f[\gamma+m]$, then, since $f$ is an increasing function, we have $f(\beta)\geq\beta$, for each $\beta\in\gamma+m$, which implies $f(\gamma+i)=\gamma+i$, for $i<m$, and, hence, $C=f[\gamma]\in{\mathbb{P}}(\gamma)$ and $A=C\cup\{\gamma,\gamma+1,\dots,\gamma+m-1\}$.
Now it is easy to show that the mapping $F:\langle{\mathbb{P}}(\gamma),\subset\rangle\rightarrow\langle{\mathbb{P}}(\gamma+m),\subset\rangle$, given by $F(C)=C\cup\{\gamma,\gamma+1,\dots,\gamma+m-1\}$, is an isomorphism. $\Box$
Theorem 3.2
[…]
Proof. Since $\delta_{1},s_{1}>0$, by Lemma 3.1 we have ${\mathbb{P}}(\alpha)\cong{\mathbb{P}}(\gamma)$, where, defining $k:=\sum_{i=1}^{n}s_{i}$, we have $\gamma=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}=\sum_{j=1}^{k}{\mathbb{L}}_{j}$. By induction on $k\in{\mathbb{N}}$ we prove that ${\mathbb{P}}(\gamma)\cong\prod_{i=1}^{n}\big{(}{\mathbb{P}}(\omega^{\delta_{i}})\big{)}^{s_{i}}(\cong\prod_{j=1}^{k}{\mathbb{P}}({\mathbb{L}}_{j})$), which will follow from
${\mathbb{P}}(\gamma)=\big{\{}\bigcup_{j=1}^{k}C_{j}:\langle C_{1},\dots,C_{k}\rangle\in\prod_{j=1}^{k}{\mathbb{P}}({\mathbb{L}}_{j})\big{\}}.$ (3)
For $k=1$ we have $n=s_{1}=1$, so $\gamma=\omega^{\delta_{1}}={\mathbb{L}}_{1}$, ${\mathbb{P}}(\gamma)={\mathbb{P}}({\mathbb{L}}_{1})$ and (3) is true.
[…]
Let $C\in{\mathbb{P}}(\gamma^{\prime})$ and $f\in\mathop{\rm Emb}\nolimits(\gamma^{\prime})$, where $f[\gamma^{\prime}]=C$. We show that $C_{1}:=f[L_{1}]$ is a cofinal subset of ${\mathbb{L}}_{1}$; then we will have $f[\gamma]\subset\gamma$ and, hence, $f[\gamma]\in{\mathbb{P}}(\gamma)$; so, by the induction hypothesis (3) will be true for $k+1$.
Assuming that $F:=f[L_{1}]\cap\gamma\neq\emptyset$ the order $F$ would be isomorphic to a final part of ${\mathbb{L}}_{1}\cong\omega^{\delta_{n}}$ and, by Fact 2.8(b) we would have $F\cong\omega^{\delta_{n}}$. So we would have $\gamma^{\prime}\cong F+f[\gamma]\subset\gamma\subset\gamma^{\prime}$ and, hence $\gamma^{\prime}\cong\gamma=\omega^{\delta_{n}}(s_{n}-1)+\dots+\omega^{\delta_{1}}s_{1}$, which is impossible since each ordinal has a unique representation in the Cantor normal form. Thus $f[L_{1}]\subset L_{1}$ and $f[L_{1}]\cong L_{1}\cong\omega^{\delta_{n}}$, which implies that $C_{1}:=f[L_{1}]$ is a cofinal subset of ${\mathbb{L}}_{1}$ and $C_{1}\in{\mathbb{P}}({\mathbb{L}}_{1})$. Consequently, $f[\gamma]\subset\gamma$ and, hence, $f[\gamma]\in{\mathbb{P}}(\gamma)$; so, by the induction hypothesis $f[\gamma]=\bigcup_{j=2}^{k+1}C_{j}$, where $\langle C_{2},\dots,C_{k}\rangle\in\prod_{j=2}^{k+1}{\mathbb{P}}({\mathbb{L}}_{j})$ and $f[\gamma^{\prime}]=\bigcup_{j=1}^{k+1}C_{j}$.
Thus, ${\mathbb{P}}(\alpha)\cong\prod_{i=1}^{n}({\mathbb{P}}(\omega^{\delta_{i}}))^{s_{i}}$ and the rest follows from Fact 2.1(c). $\Box$
4 Factorization of factors ${\mathbb{P}}(\omega^{\delta})$. ZFC results
Classification of exponents $\delta$
If $\delta$ is an ordinal, $\mathop{\rm cf}\nolimits(\delta)=\kappa\geq\omega$, $\langle\delta_{\xi}:\xi\in\kappa\rangle\in{}^{\kappa}\delta$ is a strictly increasing sequence of ordinals and $\delta=\sup\{\delta_{\xi}:\xi\in\kappa\}$, then we will write $\delta=\lim_{\xi\rightarrow\kappa}\delta_{\xi}$. Let $\mathop{\rm Lim}\nolimits_{0}:=\{0\}\cup\mathop{\rm Lim}\nolimits$.
Lemma 4.1
If $\delta>0$ is an ordinal and $\mathop{\rm cf}\nolimits(\delta)=\kappa$, then exactly one of the following conditions is satisfied.
  • (A)
    $\delta$ is a successor ordinal or $\mathop{\rm cf}\nolimits(\delta)=\omega$ (that is, $\kappa\leq\omega$);
  • (B)
    $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=\omega$,
    and $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$;
  • (C)
    $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)>\omega$,
    and $\theta=\lim_{\xi\rightarrow\,\lambda}\delta_{\xi}$, where $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi\in\lambda:=\mathop{\rm cf}\nolimits(\theta)$;
  • (D)
    $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\mathop{\rm cf}\nolimits(\theta)\geq\kappa>\omega$ or $\theta=0$;
  • (E)
    $\delta=\lim_{\xi\rightarrow\kappa}\delta_{\xi}$, where $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa>\omega$, for all $\xi\in\kappa$.
Proof. Let $\kappa>\omega$. Since $\delta\geq\mathop{\rm cf}\nolimits(\delta)=\kappa$, the ordinal $\delta$ can be presented in the Cantor normal form in the base $\kappa$
$\delta=\kappa^{\xi_{n}}\zeta_{n}+\kappa^{\xi_{n-1}}\zeta_{n-1}+\dots+\kappa^{\xi_{1}}\zeta_{1}+\beta,$ (4)
where $n\in{\mathbb{N}}$, $1\leq\zeta_{k}<\kappa$, for $k\leq n$, $0\leq\beta<\kappa$ and $1\leq\xi_{1}<\xi_{2}<\dots<\xi_{n}$. First, $\beta>0$ would mean that $\beta$, and, hence, $\delta$, is a successor ordinal or a limit ordinal of cofinality $<\kappa$; so $\beta=0$.
Second, since $\zeta_{1}<\kappa$, $\zeta_{1}\in\mathop{\rm Lim}\nolimits$ would imply that $\mathop{\rm cf}\nolimits(\zeta_{1})<\kappa$ and the final part $\kappa^{\xi_{1}}\zeta_{1}$ of $\delta$, which is a $\zeta_{1}$-sum of copies of $\kappa^{\xi_{1}}$ would be of cofinality $<\kappa$; this would imply that $\mathop{\rm cf}\nolimits(\delta)<\kappa$. Thus $\kappa^{\xi_{1}}\zeta_{1}=\kappa^{\xi_{1}}(\gamma+m)=\kappa^{\xi_{1}}\gamma+\kappa^{\xi_{1}}+\dots+\kappa^{\xi_{1}}$, where $\gamma\in\mathop{\rm Lim}\nolimits_{0}\;\cap\;\kappa$ and $m\in{\mathbb{N}}$, so $\delta$ has a final part isomorphic to $\kappa^{\xi_{1}}$.
Third, $\omega\leq\mathop{\rm cf}\nolimits(\xi_{1})<\kappa$ would imply that $\mathop{\rm cf}\nolimits(\kappa^{\xi_{1}})<\kappa$ and, hence, $\mathop{\rm cf}\nolimits(\delta)<\kappa$. So either $\mathop{\rm cf}\nolimits(\xi_{1})=\kappa$ or $\xi_{1}=\gamma^{\prime}+m^{\prime}$, where $\gamma^{\prime}\in\mathop{\rm Lim}\nolimits_{0}$ and $m^{\prime}\in{\mathbb{N}}$. Thus defining
$\theta=\kappa^{\xi_{n}}\zeta_{n}+\dots+\kappa^{\xi_{2}}\zeta_{2}+\kappa^{\xi_{1}}\gamma+\kappa^{\xi_{1}}(m-1),$ (5)
where $\mbox{where }\;\zeta_{1}=\gamma+m\;\;\mbox{ and }\;\;\gamma\in\mathop{\rm Lim}\nolimits_{0}\;\cap\;\kappa\;\;\mbox{ and }\;\;m\in{\mathbb{N}},$ (6)
we have
$\delta$ $=$ $\theta+\kappa^{\xi_{1}},\mbox{ where }\mathop{\rm cf}\nolimits(\xi_{1})=\kappa,\mbox{ or }$ (7)
$\delta$ $=$ $\theta+\kappa^{\gamma^{\prime}+m^{\prime}},\mbox{ where }\gamma^{\prime}\in\mathop{\rm Lim}\nolimits_{0}\mbox{ and }m^{\prime}\in{\mathbb{N}}.$ (8)
If (7) holds, let $\langle\alpha_{\xi}:\xi<\kappa\rangle$ be an increasing sequence cofinal in $\xi_{1}$ and let us define $\delta_{\xi}:=\theta+\kappa^{\alpha_{\xi}+1}$, for $\xi<\kappa$. Then for $\xi<\xi^{\prime}<\kappa$ we have $\delta_{\xi}<\delta_{\xi^{\prime}}<\delta$ and $\sup\{\delta_{\xi}:\xi<\kappa\}=\theta+\sup\{\kappa^{\alpha_{\xi}+1}:\xi<\kappa\}=\theta+\kappa^{\xi_{1}}=\delta$. So the sequence $\langle\delta_{\xi}:\xi\in\kappa\rangle$ is increasing, cofinal in $\delta$ and $\mathop{\rm cf}\nolimits(\delta_{\xi})=\mathop{\rm cf}\nolimits(\kappa^{\alpha_{\xi}}\kappa)=\kappa$, because $\kappa^{\alpha_{\xi}}\kappa$ is a $\kappa$-sum. So (E) is true.
If (8) holds and $\gamma^{\prime}+m^{\prime}\geq 2$, then we have $\alpha:=\gamma^{\prime}+m^{\prime}-1\geq 1$ and $\delta=\theta+\kappa^{\alpha}\kappa$. Thus the final part $\kappa^{\alpha}\kappa$ of $\delta$ is a $\kappa$-sum of copies of the ordinal $\kappa^{\alpha}\geq\kappa$. Now defining the ordinals $\delta_{\xi}:=\theta+\kappa^{\alpha}\xi+\kappa$, for $\xi<\kappa$, we have
$\kappa^{\alpha}\xi+\kappa\leq\kappa^{\alpha}\xi+\kappa^{\alpha}=\kappa^{\alpha}(\xi+1)<\kappa^{\alpha}(\xi+1)+\kappa\leq\kappa^{\alpha}(\xi+2)<\kappa^{\alpha}\kappa$
and, hence, $\delta_{\xi}<\delta$. Clearly we have $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi<\kappa$, and $\langle\delta_{\xi}:\xi<\kappa\rangle$ is an increasing sequence cofinal in $\delta$. So (E) is true again.
If (8) holds and $\gamma^{\prime}+m^{\prime}<2$, then $\gamma^{\prime}=0$, $m^{\prime}=1$, $\xi_{1}=1$ and $\delta=\theta+\kappa$, where by (5) and (6) we have
$\theta=(\kappa^{\xi_{n}}\zeta_{n}+\dots+\kappa^{\xi_{2}}\zeta_{2})+\kappa(\gamma+m-1)=\theta^{\prime}+\kappa\alpha,$ (9)
where $\theta^{\prime}=0$, if $n=1$, and $\alpha:=\gamma+m-1$ is an ordinal satisfying $0\leq\alpha<\kappa$. If $\theta=0$, then we have (D).
If $\theta>0$, then $\alpha>0$ or $\theta^{\prime}>0$ (and, hence, $\xi_{2}>1$ and $\zeta_{2}\geq 1$), which gives $\theta\geq\kappa$. In addition, the equality (9) presents the ordinal $\theta$ in the Cantor normal form and, because of its uniqueness, $\theta$ can not be presented in the form $\theta_{1}+\beta$, where $0<\beta<\kappa$; thus, $\theta$ is a limit ordinal. First, if $\mathop{\rm cf}\nolimits(\theta)\geq\kappa$, we have (D).
Second, if $\lambda:=\mathop{\rm cf}\nolimits(\theta)<\kappa$, let $\langle\alpha_{\xi}:\xi<\lambda\rangle$ be an increasing sequence cofinal in $\theta$. Assuming that $\alpha_{\xi}+\kappa\geq\theta$, for some $\xi<\lambda$, since $\mathop{\rm cf}\nolimits(\theta)<\kappa=\mathop{\rm cf}\nolimits(\alpha_{\xi}+\kappa)$ we would have $\alpha_{\xi}+\kappa>\theta$, and, hence, $\theta=\alpha_{\xi}+\beta$, where $0<\beta<\kappa$, which is false. So $\alpha_{\xi}+\kappa<\theta$, for $\xi<\lambda$, and $\theta=\sup\{\alpha_{\xi}:\xi<\lambda\}\leq\sup\{\alpha_{\xi}+\kappa:\xi<\lambda\}\leq\theta$. Defining $\delta_{\xi}=\alpha_{\xi}+\kappa$, for $\xi<\lambda$, we have $\theta=\sup\{\delta_{\xi}:\xi<\lambda\}$ and $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi<\lambda$. Now, if $\lambda=\omega$ we have (B) and if $\lambda>\omega$ we have (C).
Finally we show that each $\delta$ satisfies exactly one of conditions (A)–(E). If (A) is true, then $\kappa\leq\omega$ and (B)–(E) are false, because (B)–(E) imply that $\kappa>\omega$. If (E) holds, then (B)–(D) are false, since (B)–(D) imply that $\delta=\theta+\kappa$ and for an ordinal $\beta\in(\theta,\theta+\kappa)$ we have $\mathop{\rm cf}\nolimits(\beta)<\kappa$.
Suppose that $\delta$ satisfies (B) and (C). Then $\delta=I+F=I^{\prime}+F^{\prime}$, where $F\cong F^{\prime}\cong\kappa$, $I\cong\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=\omega$, and $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$; and $I^{\prime}\cong\theta^{\prime}\geq\kappa>\mathop{\rm cf}\nolimits(\theta^{\prime})=\lambda>\omega$, and $\theta^{\prime}=\lim_{\xi\rightarrow\,\lambda}\delta_{\xi}$, where $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi\in\lambda$. Since $\mathop{\rm cf}\nolimits(\theta)<\mathop{\rm cf}\nolimits(\theta^{\prime})$ we have $I\neq I^{\prime}$; so either $I\varsubsetneq I^{\prime}$ or $I^{\prime}\varsubsetneq I$. If $I\varsubsetneq I^{\prime}$, then there is $\xi\in\lambda$ such that $\delta_{\xi}\in I^{\prime}\setminus I$; thus $\delta_{\xi}\in(\theta,\theta+\kappa)$ and, as above, $\mathop{\rm cf}\nolimits(\delta_{\xi})<\kappa$, which is false because $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$. If $I^{\prime}\varsubsetneq I$, then there is $n\in\omega$ such that $\delta_{n}\in I\setminus I^{\prime}$; thus $\delta_{n}\in(\theta^{\prime},\theta^{\prime}+\kappa)$ and, hence, $\mathop{\rm cf}\nolimits(\delta_{n})<\kappa$, which is false since $\,\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$. Thus, conditions (B) and (C) are disjoint and in the same way we show that the pairs (B),(D) and (C),(D) are disjoint as well. $\Box$
Example 4.2
Regarding the classification given in Lemma 4.1 and the ordinals $\delta$ of, say, size $\omega_{2}$, (A) holds for $\omega_{2}+1$ and $\omega_{2}+\omega$, (B) for $\omega_{2}\omega+\omega_{2}$, (C) for $\omega_{2}\omega_{1}+\omega_{2}$, (D) for $\omega_{2}$ and $\omega_{2}+\omega_{2}$, and (E) holds for $\omega_{2}\omega_{2}$.
The separative modification of ${\mathbb{P}}(\omega^{\delta})$
Fact 4.3
If $\delta>0$ is an ordinal, then we have
(a) ${\mathbb{P}}(\omega^{\delta})=\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle$, $\mathop{\rm sm}\nolimits({\mathbb{P}}(\omega^{\delta}))=\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset_{{\mathcal{I}}_{\omega^{\delta}}}\rangle$ and $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))=(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$;
(b) If $\delta\geq\omega$, then $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|\leq 2^{|\delta|}$.
Proof. (a) By Fact 2.8(f), ${\mathcal{I}}_{\omega^{\delta}}=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}$ is an ideal in $P(\omega^{\delta})$. Clearly we have ${\mathbb{P}}(\omega^{\delta})\subset{\mathcal{I}}_{\omega^{\delta}}^{+}$. Conversely, if $A\in{\mathcal{I}}_{\omega^{\delta}}^{+}$ and $C\in{\mathbb{P}}(\omega^{\delta})$, where $C\subset A$, then $\omega^{\delta}\hookrightarrow A\hookrightarrow\omega^{\delta}$, which implies that ${\mathbb{A}}\cong\omega^{\delta}$ and, hence, $A\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(\omega^{\delta})={\mathcal{I}}_{\omega^{\delta}}^{+}$. The rest follows from Fact 2.1(d).
(b) By Fact 2.9(a), $|\omega^{\delta}|=|\delta|$; so, by (a), $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|\leq|P(\omega^{\delta})|=2^{|\delta|}$. $\Box$
Lemma 4.4
Let $\delta$ be an ordinal. Then
(a) If $\mathop{\rm cf}\nolimits(\delta)=\kappa\geq\omega$ and $\delta=\lim_{\alpha\rightarrow\kappa}\delta_{\alpha}$, then $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}:=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, for $\alpha<\kappa$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, if $\alpha\neq\beta$.
(b) If $\delta=\delta^{\prime}+1$, then $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}\cong\sum_{\omega}\omega^{\delta^{\prime}}=\sum_{\alpha<\omega}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta^{\prime}}$, for $\alpha<\omega$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, for $\alpha\neq\beta$.
Proof. (a) By the assumptions $\delta$ is a limit ordinal; so, $\omega^{\delta}=\sup\{\omega^{\eta}:\eta<\delta\}$, and for $\eta<\delta$ there is $\alpha<\kappa$ such that $\eta<\delta_{\alpha}$, which implies $\omega^{\eta}<\omega^{\delta_{\alpha}}<\omega^{\delta}$. Thus $\langle\omega^{\delta_{\alpha}}:\alpha<\kappa\rangle$ is an increasing cofinal sequence in $\omega^{\delta}$ and, hence, $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$.
Let $\theta:=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$ and w.l.o.g. suppose that $\theta=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, $\alpha<\kappa$, are disjoint. Clearly we have $\theta\geq\omega^{\delta_{\alpha}}$, for all $\alpha<\kappa$; so, $\theta\geq\sup\{\omega^{\delta_{\alpha}}:\alpha<\kappa\}=\omega^{\delta}$. Suppose that $\omega^{\delta}<\theta$. Then we would have $\omega^{\delta}\in L_{\alpha_{0}}$, for some $\alpha_{0}<\kappa$; let $\xi=\min L_{\alpha_{0}}$. By Fact 2.8(b), $\omega^{\delta}>\xi$ would imply that $\omega^{\delta}\hookrightarrow L_{\alpha_{0}}\cong\omega^{\delta_{\alpha_{0}}}$ which is impossible because $\delta_{\alpha_{0}}<\delta$ and, hence, $\omega^{\delta_{\alpha_{0}}}<\omega^{\delta}$. So $\omega^{\delta}=\xi$ and, hence, $\omega^{\delta}=\sum_{\alpha<\alpha_{0}}\omega^{\delta_{\alpha}}$. Now, if $\alpha_{0}=\alpha+1$ we would have $\omega^{\delta}\hookrightarrow L_{\alpha}\cong\omega^{\delta_{\alpha}}$, which is impossible, and if $\alpha_{0}$ is a limit ordinal we would have $\mathop{\rm cf}\nolimits(\omega^{\delta})\leq\mathop{\rm cf}\nolimits(\alpha_{0})\leq|\alpha_{0}|<\kappa$, which is impossible too. So, $\omega^{\delta}=\theta=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$.
(b) If $\delta=\delta^{\prime}+1$, then $\omega^{\delta}=\omega^{\delta^{\prime}+1}=\omega^{\delta^{\prime}}\omega=\sum_{\omega}\omega^{\delta^{\prime}}$ and $\langle\omega^{\delta^{\prime}}n:n\in\omega\rangle$ is an increasing and cofinal sequence in $\omega^{\delta}$, which gives $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$. $\Box$
Fact 4.5
If $\delta>0$ is an ordinal and $\omega^{\delta}\hookrightarrow L=\sum_{i\in I}L_{i}$, where $I$ and $L_{i}$, $i\in I$, are linear orders and $|I|<\mathop{\rm cf}\nolimits(\omega^{\delta})$, then $\omega^{\delta}\hookrightarrow L_{i}$, for some $i\in I$.
Proof. Let $\kappa=\mathop{\rm cf}\nolimits(\omega^{\delta})$ and let $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ be an increasing cofinal sequence in $\omega^{\delta}$. Then $\omega^{\delta}=\sup\{\beta_{\alpha}:\alpha<\kappa\}$ and for $\alpha<\kappa$ we have $g(\beta_{\alpha})\in\bigcup_{i\in I}L_{i}$, which implies that there is a unique $i_{\alpha}\in I$ such that $g(\beta_{\alpha})\in L_{i_{\alpha}}$. Clearly, the function $h:\kappa\rightarrow I$ defined by $h(\alpha)=i_{\alpha}$ is non-decreasing and, since $|I|<\kappa=\bigcup_{i\in I}h^{-1}[\{i\}]$, by the regularity of $\kappa$ there is $i^{*}\in I$ such that $h^{-1}[\{i^{*}\}]\in[\kappa]^{\kappa}$. Let $h^{-1}[\{i^{*}\}]=\{\alpha_{\xi}:\xi<\ \kappa\}$ be an increasing enumeration. Now, if $\alpha\in\kappa\setminus\alpha_{0}$ and $\xi\in\kappa$ where $\alpha<\alpha_{\xi}$, then $i^{*}=h(\alpha_{0})\leq h(\alpha)\leq h(\alpha_{\xi})=i^{*}$ thus for each $\alpha\in[\alpha_{0},\kappa)$ we have $h(\alpha)=i^{*}$, that is $g(\beta_{\alpha})\in L_{i^{*}}$. So, if $\beta_{\alpha_{0}}\leq\theta<\omega^{\delta}$, then, by the cofinality of the sequence $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ in $\omega^{\delta}$, there is $\alpha\in\kappa$ such that $\theta\leq\beta_{\alpha}$, which implies that $g(\beta_{\alpha_{0}})\leq g(\theta)\leq g(\beta_{\alpha})$. So, since $g(\beta_{\alpha_{0}}),g(\beta_{\alpha})\in L_{i^{*}}$ and $L_{i^{*}}$ is a convex suborder of $L$, we have $g(\theta)\in L_{i^{*}}$, for all $\theta\in[\beta_{\alpha_{0}},\omega^{\delta})$. By Fact 2.8(b) we have $\omega^{\delta}\cong[\beta_{\alpha_{0}},\omega^{\delta})\cong g[[\beta_{\alpha_{0}},\omega^{\delta})]\subset L_{i^{*}}$, that is $\omega^{\delta}\hookrightarrow L_{i^{*}}$. $\Box$
Proposition 4.6
Let $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$, where $\delta_{\alpha}\leq\delta_{\beta}$, whenever $\alpha<\beta<\kappa$. If $L:=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, for $\alpha<\kappa$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, for $\alpha\neq\beta$, then
(a) Each $C\in{\mathbb{P}}(L)$ intersects $\kappa$-many summands $L_{\alpha}$;
(b) If $C\subset L$, then $C\in{\mathbb{P}}(L)$ iff
$\forall\alpha,\eta<\kappa\;\;\exists K\in[\kappa\setminus\eta]^{<\kappa}\;\;L_{\alpha}\hookrightarrow\bigcup_{\beta\in K}L_{\beta}\cap C;$ (10)
(c) If, in addition, $\mathop{\rm cf}\nolimits(\omega^{\delta_{\alpha}})\geq\kappa$, for all $\alpha<\kappa$,
$C\in{\mathbb{P}}(L)\Leftrightarrow\forall\alpha<\kappa\;\;S^{\alpha}_{C}:=\{\beta<\kappa:\omega^{\delta_{\alpha}}\hookrightarrow L_{\beta}\cap C\}\in[\kappa]^{\kappa}.$ (11)
Proof. (a) Let $C\in{\mathbb{P}}(L)$ and suppose that $C\subset\bigcup_{\alpha<\eta}L_{\alpha}$, for some $\eta<\kappa$. By Fact 2.8(b) we have $L\cong\bigcup_{\alpha\geq\eta}L_{\alpha}$ and, hence, $L+L\hookrightarrow L$, which is, by Fact 2.7(a), impossible.
(b) Let $C\in{\mathbb{P}}(L)$, let $f:L\hookrightarrow L$, where $C=f[L]=\sum_{\alpha\in\kappa}f[L_{\alpha}]$ and let $\alpha_{0},\eta_{0}\in\kappa$. By (a) there are $\alpha_{1}<\kappa$ and $\eta_{1}\geq\eta_{0}$ such that $f[L_{\alpha_{1}}]\cap L_{\eta_{1}}\neq\emptyset$ and, hence, $C_{1}:=\bigcup_{\alpha>\alpha_{1}}f[L_{\alpha}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. Since $C_{1}$ is a final part of $C$ and $C\cong L$, by Fact 2.8(b) there is $g:L\hookrightarrow\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$ and, hence, $g[L_{\alpha_{0}}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. If $\beta_{0}\geq\eta_{0}$, where $g[L_{\alpha_{0}+1}]\cap L_{\beta_{0}}\neq\emptyset$, then, since $g[L_{\alpha_{0}}]<g[L_{\alpha_{0}+1}]$, we have $g[L_{\alpha_{0}}]\subset\bigcup_{\eta_{0}\leq\beta\leq\beta_{0}}L_{\beta}\cap C$. Thus $L_{\alpha_{0}}\hookrightarrow\bigcup_{\beta\in K}L_{\beta}\cap C$, where $K=[\eta_{0},\beta_{0}]\in[\kappa\setminus\eta_{0}]^{<\kappa}$, and (10) is true.
Assuming that (10) holds, by recursion we define sequences $\langle K_{\alpha}:\alpha\in\kappa\rangle$ and $\langle f_{\alpha}:\alpha\in\kappa\rangle$ such that for each $\alpha,\alpha^{\prime}\in\kappa$ we have:
(i) $K_{\alpha}\in[\kappa]^{<\kappa}$,
(ii) $\alpha<\alpha^{\prime}\Rightarrow K_{\alpha}<K_{\alpha^{\prime}}$ and
(iii) $f_{\alpha}:L_{\alpha}\hookrightarrow\bigcup_{\beta\in K_{\alpha}}L_{\beta}\cap C$.
By (10) for $\alpha=\eta=0$ there are $K_{0}\in[\kappa]^{<\kappa}$ and $f_{0}:L_{0}\hookrightarrow\bigcup_{\beta\in K_{0}}L_{\beta}\cap C$.
Let $\alpha^{\prime}<\kappa$ and let $K_{\alpha}$ and $f_{\alpha}$ be defined for $\alpha<\alpha^{\prime}$ such that (i)-(iii) is true. By (i) and the regularity of $\kappa$ we have $\eta:=\sup(\bigcup_{\alpha<\alpha^{\prime}}K_{\alpha})+1<\kappa$ and, by (10), there are $K_{\alpha^{\prime}}\in[\kappa\setminus\eta]^{<\kappa}$ and $f_{\alpha^{\prime}}:L_{\alpha^{\prime}}\hookrightarrow\bigcup_{\beta\in K_{\alpha^{\prime}}}L_{\beta}\cap C$; the recursion works.
Let $f=\bigcup_{\alpha\in\kappa}f_{\alpha}$. By (ii) and (iii), $\alpha<\alpha^{\prime}$ implies $f_{\alpha}[L_{\alpha}]<f_{\alpha^{\prime}}[L_{\alpha^{\prime}}]$ and, hence, $f:L\hookrightarrow C$. Thus $C^{\prime}=f[L]\in{\mathbb{P}}(L)$, $C^{\prime}\subset C$ and, by Fact 4.3, $C\in{\mathbb{P}}(L)$.
(c) By (b) and Fact 4.5 we have $C\in{\mathbb{P}}(L)$ iff for each $\alpha<\kappa$, for each $\eta<\kappa$ there is $\beta\geq\eta$ such that $\omega^{\delta_{\alpha}}\hookrightarrow L_{\beta}\cap C$. $\Box$
Complete embeddings of ${\mathcal{P}}_{\kappa}$
Theorem 4.7
If $\delta$ is an ordinal, $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa\geq\omega$, $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$, where $\delta_{\xi}\leq\delta_{\zeta}$, whenever $\xi<\zeta<\kappa$, and $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})\geq\kappa$, for all $\xi<\kappa$, then ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Proof. Let ${\mathbb{L}}_{\xi}\cong\omega^{\delta_{\xi}}$, for $\xi<\kappa$, where $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Then $\omega^{\delta}\cong{\mathbb{L}}:=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$ and, by Fact 2.1(d) and Fact 4.3, $\langle P(\omega^{\delta})\setminus{\mathcal{I}}_{\omega^{\delta}},\subset_{\omega^{\delta}}\rangle=\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}({\mathbb{L}})=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$. For $A\subset L$ by Proposition 4.6(c) we have
$A\in P(L)\setminus{\mathcal{I}}\Leftrightarrow\forall\zeta<\kappa\;\;S^{\zeta}_{A}:=\{\xi<\kappa:\omega^{\delta_{\zeta}}\hookrightarrow L_{\xi}\cap A\}\in[\kappa]^{\kappa},$ (12)
and, since $S^{\zeta}_{A}\subset\mathop{\rm supp}\nolimits(A):=\{\xi<\kappa:L_{\xi}\cap A\neq\emptyset\}$, for $\zeta<\kappa$,
$|\mathop{\rm supp}\nolimits(A)|<\kappa\Rightarrow A\in{\mathcal{I}}.$ (13)
By Fact 2.2 it is sufficient to show that $f:\langle[\kappa]^{\kappa},\subset^{*}\rangle\hookrightarrow_{c}\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where $f(S)=\bigcup_{\xi\in S}L_{\xi}$, for $S\in[\kappa]^{\kappa}$. First, if $S\in[\kappa]^{\kappa}$ and $\zeta<\kappa$, then $|S\setminus\zeta|=\kappa$ and for $\xi\in S\setminus\zeta$ we have $\delta_{\zeta}\leq\delta_{\xi}$, which gives $\omega^{\delta_{\zeta}}\hookrightarrow\omega^{\delta_{\xi}}\cong L_{\xi}=L_{\xi}\cap f(S)$ and, hence, $\xi\in S^{\zeta}_{f(S)}$. Thus $[\kappa]^{\kappa}\ni S\setminus\zeta\subset S^{\zeta}_{f(S)}$, for each $\zeta<\kappa$, and by (12) we have $f(S)\in P(L)\setminus{\mathcal{I}}$. So, $f$ maps $[\kappa]^{\kappa}$ to $P(L)\setminus{\mathcal{I}}$. Let $S,T\in[\kappa]^{\kappa}$.
(ce1) If $S\subset^{*}T$, then $|\mathop{\rm supp}\nolimits(f(S)\setminus f(T))|=|\mathop{\rm supp}\nolimits(\bigcup_{\xi\in S\setminus T}L_{\xi})|=|S\setminus T|<\kappa$ and by (13) $f(S)\setminus f(T)\in{\mathcal{I}}$, that is $f(S)\subset_{\mathcal{I}}f(T)$.
(ce2) If $S\perp T$, then we have $|\mathop{\rm supp}\nolimits(f(S)\cap f(T))|=|\mathop{\rm supp}\nolimits(\bigcup_{\xi\in S\cap T}L_{\xi})|=|S\cap T|<\kappa$ and by (13) $f(S)\cap f(T)\in{\mathcal{I}}$, that is $f(S)\perp_{\subset_{\mathcal{I}}}f(T)$. If $S\not\perp T$, then $S\cap T\in[\kappa]^{\kappa}$ and $f(S)\cap f(T)=\bigcup_{\xi\in S\cap T}L_{\xi}=f(S\cap T)\in P(L)\setminus{\mathcal{I}}$ and, hence, $f(S)\not\perp_{\subset_{\mathcal{I}}}f(T)$.
(ce3) First we show that a set $S\in[\kappa]^{\kappa}$ is a reduction of $A\in P(L)\setminus{\mathcal{I}}$ to $\langle[\kappa]^{\kappa},\subset^{*}\rangle$, if $S$ is a pseudointersection of the family $\{S^{\zeta}_{A}:\zeta<\kappa\}$, that is
$(\forall\zeta<\kappa\;\;S\subset^{*}S^{\zeta}_{A})\;\Rightarrow\;\forall T\in[\kappa]^{\kappa}\;\;(T\subset^{*}S\Rightarrow f(T)\cap A\in P(L)\setminus{\mathcal{I}}).$ (14)
So, let $[\kappa]^{\kappa}\ni S\subset^{*}S^{\zeta}_{A}$, for all $\zeta<\kappa$, and $[\kappa]^{\kappa}\ni T\subset^{*}S$. Then for each $\zeta<\kappa$ we have $T\subset^{*}S^{\zeta}_{A}$, which implies that
$\forall\zeta<\kappa\;\;|T\cap S^{\zeta}_{A}|=\kappa.$ (15)
By recursion we construct a sequence $\langle\alpha_{\xi}:\xi<\kappa\rangle$ such that for each $\xi<\kappa$
(i) $\alpha_{\xi}\in T$,
(ii) $\xi^{\prime}<\xi\Rightarrow\alpha_{\xi^{\prime}}<\alpha_{\xi}$ and
(iii) $\omega^{\delta_{\xi}}\hookrightarrow L_{\alpha_{\xi}}\cap A$.
By (15) there is $\alpha_{0}\in T\cap S^{0}_{A}$ and, hence, $\omega^{\delta_{0}}\hookrightarrow L_{\alpha_{0}}\cap A$. Let $\zeta<\kappa$ and let $\langle\alpha_{\xi}:\xi<\zeta\rangle$ be a sequence satisfying (i)–(iii). By (15) we have $|T\cap S^{\zeta}_{A}|=\kappa$ so there is $\alpha_{\zeta}=\min(T\cap S^{\zeta}_{A}\setminus(\sup\{\alpha_{\xi}:\xi<\zeta\}+1))$. Now we have $\alpha_{\zeta}\in T$ and $\alpha_{\zeta}>\alpha_{\xi}$, for $\xi<\zeta$, so (i) and (ii) are true. Since $\alpha_{\zeta}\in S^{\zeta}_{A}$ we have $\omega^{\delta_{\zeta}}\hookrightarrow L_{\alpha_{\zeta}}\cap A$ so (iii) is true and the recursion works.
Clearly $B:=\bigcup_{\xi<\kappa}L_{\alpha_{\xi}}\cap A\subset A$ and by (i) we have $B\subset\bigcup_{\xi<\kappa}L_{\alpha_{\xi}}\subset\bigcup_{\zeta\in T}L_{\zeta}=f(T)$. If $\zeta<\kappa$, then for $\xi\geq\zeta$ we have $L_{\alpha_{\xi}}\cap B=L_{\alpha_{\xi}}\cap A$ and, by (iii) and since $\delta_{\zeta}\leq\delta_{\xi}$, we have $\omega^{\delta_{\zeta}}\hookrightarrow\omega^{\delta_{\xi}}\hookrightarrow L_{\alpha_{\xi}}\cap A=L_{\alpha_{\xi}}\cap B$, which implies that $\alpha_{\xi}\in S^{\zeta}_{B}$. So for each $\zeta<\kappa$ we have $\{\alpha_{\xi}:\xi\geq\zeta\}\subset S^{\zeta}_{B}$, which by (ii) implies that $S^{\zeta}_{B}\in[\kappa]^{\kappa}$. Thus, by (12), $B\in P(L)\setminus{\mathcal{I}}$ and, since $B\subset f(T)\cap A$, we have $f(T)\cap A\in P(L)\setminus{\mathcal{I}}$. So (14) is true.
Finally we prove (ce3). If $A\in P(L)\setminus{\mathcal{I}}$, then by (12) and since $\delta_{0}\leq\delta_{1}\leq\dots$ we have
$S^{0}_{A}\supset S^{1}_{A}\supset\dots\supset S^{\zeta}_{A}\supset\dots\;\;\mbox{ and }\;\;\forall\zeta<\kappa\;\;|S^{\zeta}_{A}|=\kappa.$ (16)
So, taking $\alpha_{0}=\min S^{0}_{A}$ and $\alpha_{\zeta}=\min(S^{\zeta}_{A}\setminus\{\alpha_{\xi}:\xi<\zeta\})$, for $\zeta<\kappa$, we have $S:=\{\alpha_{\zeta}:\zeta<\kappa\}\in[\kappa]^{\kappa}$ and, by (16), $\{\alpha_{\xi}:\xi\geq\zeta\}\subset S^{\zeta}_{A}$, which implies that $S\subset^{*}S^{\zeta}_{A}$, for all $\zeta<\kappa$. By (14), $S$ is a reduction of $A$ to $\langle[\kappa]^{\kappa},\subset^{*}\rangle$. $\Box$
Theorem 4.8
If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa=\mathop{\rm cf}\nolimits(\delta)>\mathop{\rm cf}\nolimits(\theta)=:\lambda\geq\omega$, then ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Proof. By Lemma 4.1, $\theta=\lim_{\nu\rightarrow\lambda}\delta_{\nu}$, where $\mathop{\rm cf}\nolimits(\delta_{\nu})=\kappa$, for all $\nu<\lambda$. Clearly we have $\omega^{\delta}=\omega^{\theta}\omega^{\kappa}=\sup\{\omega^{\delta_{\nu}}:\nu<\lambda\}\kappa$. Let ${\mathbb{L}}=\langle L,<\rangle=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$, where ${\mathbb{L}}_{\xi}=\langle L_{\xi},<_{\xi}\rangle\cong\omega^{\theta}$, for $\xi<\kappa$, and $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Let =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$ and, for $A\subset L$ and $\nu<\lambda$, let
$S^{\nu}_{A}=\{\xi<\kappa:\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap A\}.$
Claim 4.9
For $A,B\subset L$ we have:
(a) $A\in P(L)\setminus{\mathcal{I}}\;\Leftrightarrow\;\forall\nu<\lambda\;\;|S^{\nu}_{A}|=\kappa$;
(b) $A\subset_{{\mathcal{I}}}B\;\Leftrightarrow\;A\setminus B\in{\mathcal{I}}\;\Leftrightarrow\;\exists\nu<\lambda\;\;|S^{\nu}_{A\setminus B}|<\kappa$.
Proof. (a) ($\Rightarrow$) Since ${\mathbb{L}}\cong\omega^{\theta}\kappa$, by Proposition 4.6(b) we have: $A\in P(L)\setminus{\mathcal{I}}$ iff
$\forall\eta<\kappa\;\;\exists K\in[\kappa\setminus\eta]^{<\kappa}\;\;\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$ (17)
and we prove
$\forall\nu<\lambda\;\;\forall\eta<\kappa\;\;\exists\xi\in\kappa\setminus\eta\;\;\xi\in S^{\nu}_{A}.$ (18)
If $\nu<\lambda$ and $\eta<\kappa$, then by (17) there is $K\in[\kappa\setminus\eta]^{<\kappa}$ such that $\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$ which implies that $\omega^{\delta_{\nu}}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$ and, since $|K|<\kappa=\mathop{\rm cf}\nolimits(\delta_{\nu})=\mathop{\rm cf}\nolimits(\omega^{\delta_{\nu}})$, by Fact 4.5 there is $\xi\in K$ such that $\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap A$ and, hence, $\xi\in S^{\nu}_{A}$. Since $K\subset\kappa\setminus\eta$ we have $\xi\geq\eta$.
($\Leftarrow$) Assuming (18) we prove (17). Let $\eta<\kappa$. By recursion we construct a sequence $\langle\xi_{\nu}:\nu<\lambda\rangle$ such that for all $\nu,\nu^{\prime}<\lambda$ we have:
(i) $\nu<\nu^{\prime}\Rightarrow\eta\leq\xi_{\nu}<\xi_{\nu^{\prime}}<\kappa$,
(ii) $\omega^{\delta_{\nu}}\hookrightarrow A\cap L_{\xi_{\nu}}$.
First, by (18), for $\nu=0$ there is $\xi_{0}\geq\eta$ such that $\xi_{0}\in S^{0}_{A}$ that is $\omega^{\delta_{0}}\hookrightarrow A\cap L_{\xi_{0}}$. If $\nu<\lambda$ and the sequence $\langle\xi_{\nu^{\prime}}:\nu^{\prime}<\nu\rangle$ satisfies (i) and (ii), then by (i) and since $\lambda<\kappa$ we have $\eta_{\nu}:=\sup\{\xi_{\nu^{\prime}}:\nu^{\prime}<\nu\}+1<\kappa$ and by (18) there is $\xi_{\nu}\geq\eta_{\nu}$ such that $\xi_{\nu}\in S^{\nu}_{A}$ that is $\omega^{\delta_{\nu}}\hookrightarrow A\cap L_{\xi_{\nu}}$. Thus the recursion works.
Now $K:=\{\xi_{\nu}:\nu<\lambda\}\in[\kappa\setminus\eta]^{<\kappa}$ and $\sum_{\nu<\lambda}\omega^{\delta_{\nu}}\hookrightarrow\bigcup_{\nu<\lambda}L_{\xi_{\nu}}\cap A$, that is $\omega^{\theta}\hookrightarrow\bigcup_{\xi\in K}L_{\xi}\cap A$. So (17) is true. Statement (b) follows from (a). $\Box$
Let $\{I_{\nu}:\nu<\lambda\}\subset[\kappa]^{\kappa}$ be a partition of $\kappa$. For $\nu<\lambda$ and $\xi\in I_{\nu}$, let $L^{\prime}_{\xi}\subset L_{\xi}$ be the initial part of ${\mathbb{L}}_{\xi}$ such that ${\mathbb{L}}_{\xi}^{\prime}\cong\omega^{\delta_{\nu}}$ and let $C=\bigcup_{\nu<\lambda}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$. Then for each $\nu<\lambda$ the set $S^{\nu}_{C}=\{\xi<\kappa:\omega^{\delta_{\nu}}\hookrightarrow L_{\xi}\cap C\}=\bigcup_{\mu\geq\nu}I_{\mu}$ is of size $\kappa$ and, by Claim 4.9(a), $C\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(C)\cong{\mathbb{P}}(\omega^{\delta})$ and by Fact 2.2 it is sufficient to prove that $f:\langle[\lambda]^{\lambda},\subset^{*}\rangle\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(C)$, where for $S\in[\lambda]^{\lambda}$
$f(S)=C_{S}:=\bigcup_{\nu\in S}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}.$
First, since $S^{\nu}_{C_{S}}=\bigcup_{\mu\in S\setminus\nu}I_{\mu}$, for $\nu<\lambda$, by Claim 4.9(a) we have $C_{S}\in{\mathbb{P}}(\omega^{\delta})$ and, hence, $C_{S}\in{\mathbb{P}}(C)$ as well. Let $S,T\in[\lambda]^{\lambda}$.
(ce1) If $S\subset^{*}T$, then, since $f(S)\setminus f(T)=\bigcup_{\nu\in S\setminus T}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$, $|S\setminus T|<\lambda$ and $\lambda$ is a regular cardinal, for $\nu>\sup(S\setminus T)$ we have $S^{\nu}_{f(S)\setminus f(T)}=\emptyset$, which by Claim 4.9(b) gives $f(S)\subset_{\mathcal{I}}f(T)$.
(ce2) Since $f(S)\cap f(T)=\bigcup_{\nu\in S\cap T}\bigcup_{\xi\in I_{\nu}}L_{\xi}^{\prime}$, by Claim 4.9(a) we have $S\perp T$ iff $|S\cap T|<\lambda$ iff $f(S)\cap f(T)\in{\mathcal{I}}$ iff $f(S)\perp_{\mathcal{I}}f(T)$.
(ce3) For $A\in{\mathbb{P}}(C)$ we will find a set $S\in[\lambda]^{\lambda}$ such that $f(T)\cap A\in{\mathbb{P}}(C)$, whenever $[\lambda]^{\lambda}\ni T\subset^{*}S$. By recursion we construct a sequence $\langle\nu_{\alpha}:\alpha<\lambda\rangle$ in $\lambda$ such that
(i) $\nu_{0}=\min\{\nu\in\lambda:|S^{0}_{A}\cap I_{\nu}|=\kappa\}$,
(ii) $\nu_{\alpha+1}=\min\{\nu\in\lambda\setminus(\nu_{\alpha}+1):|S^{\nu_{\alpha}+1}_{A}\cap I_{\nu}|=\kappa\}$, for $\alpha<\lambda$,
(iii) $\nu_{\beta}=\sup\{\nu_{\alpha}:\alpha<\beta\}$, for $\beta\in\mathop{\rm Lim}\nolimits\cap\lambda$.
Since $A\in{\mathbb{P}}({\mathbb{L}})$, by Claim 4.9(a) we have $[\kappa]^{\kappa}\ni S^{0}_{A}\subset\bigcup_{\nu<\lambda}I_{\nu}$ and, since $\kappa$ is a regular cardinal and $\lambda<\kappa$, the ordinal $\nu_{0}$ is well defined by (i). Let $\beta<\lambda$ and let $\langle\nu_{\alpha}:\alpha<\beta\rangle$ be a sequence satisfying (i)–(iii). If $\beta=\alpha+1$, then for $\nu\leq\nu_{\alpha}$ and $\xi\in I_{\nu}$ we have $\mathop{\mathrm{tp}}\nolimits(L_{\xi}^{\prime})\leq\omega^{\delta_{\nu_{\alpha}}}$ and, since $A\subset C$, we have $\omega^{\delta_{\nu_{\alpha}+1}}\not\hookrightarrow L_{\xi}^{\prime}\cap A=L_{\xi}\cap A$. Thus $S^{\nu_{\alpha}+1}_{A}\subset\bigcup_{\nu\geq\nu_{\alpha}+1}I_{\nu}$ and, as above, there is $\nu\geq\nu_{\alpha}+1$ such that $|S^{\nu_{\alpha}+1}_{A}\cap I_{\nu}|=\kappa$. So $\nu_{\alpha+1}$ is well defined by (ii). If $\beta$ is a limit ordinal, then $\nu_{\beta}$ is well defined by (iii), because $\lambda$ is a regular cardinal and $\nu_{\alpha}<\lambda$, for all $\alpha<\beta$. Thus, the recursion works and by (ii) and (iii) we have $S:=\{\nu_{\alpha}:\alpha<\lambda\}\in[\lambda]^{\lambda}$.
Now, if $T\in[\lambda]^{\lambda}$ and $T\subset^{*}S$, then $|T\setminus S|<\lambda$ and, hence, there is $\beta<\lambda$ such that $\{\nu_{\alpha}:\alpha\geq\beta\}\subset T$. So, we have $\bigcup_{\alpha\geq\beta}\bigcup_{\xi\in I_{\nu_{\alpha}}}L_{\xi}^{\prime}\subset f(T)$, which gives $B:=\bigcup_{\alpha\geq\beta}\bigcup_{\xi\in I_{\nu_{\alpha}}}L_{\xi}^{\prime}\cap A\subset f(T)\cap A$. For $\alpha\geq\beta$ by (ii) we have $|S^{\nu_{\alpha}+1}_{A}\cap I_{\nu_{\alpha+1}}|=\kappa$, which means that there are $\kappa$-many $\xi\in I_{\nu_{\alpha+1}}$ such that $\omega^{\delta_{\nu_{\alpha}+1}}\hookrightarrow L_{\xi}\cap A=L_{\xi}\cap B$. Since the sequence $\langle\nu_{\alpha}+1:\alpha<\lambda\rangle$ is cofinal in $\lambda$, by Claim 4.9(a) we have $B\in{\mathbb{P}}(C)$ and, consequently, $f(T)\cap A\in{\mathbb{P}}(C)$. $\Box$
Factorization of factors ${\mathbb{P}}(\omega^{\delta})$
Theorem 4.10
If $\delta>0$ is an ordinal and $\mathop{\rm cf}\nolimits(\delta)=\kappa$, then for ${\mathbb{P}}(\omega^{\delta})$ we have
  • (A)
    If $\delta$ is a successor ordinal or $\mathop{\rm cf}\nolimits(\delta)=\omega$ (that is, $\kappa\leq\omega$),
    then ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed;
  • (B)
    If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=\omega$
    and $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$,
    then ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed;
  • (C)
    If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\kappa>\mathop{\rm cf}\nolimits(\theta)=:\lambda>\omega$
    and $\theta=\lim_{\xi\rightarrow\,\lambda}\delta_{\xi}$, where $\,\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi\in\lambda$,
    then ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$;
  • (D)
    If $\delta=\theta+\kappa$, where $\mathop{\mathrm{Ord}}\nolimits\ni\theta\geq\mathop{\rm cf}\nolimits(\theta)\geq\kappa>\omega$ or $\theta=0$,
    then ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$;
  • (E)
    If $\delta=\lim_{\xi\rightarrow\kappa}\delta_{\xi}$, where $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa>\omega$, for all $\xi<\kappa$,
    then ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
In cases (A) and (B) we have ${\mathbb{P}}(\omega^{\delta})\equiv_{forc}{\mathcal{P}}_{\omega}\ast\pi$, where $\pi$ is a ${\mathcal{P}}_{\omega}$-name for an $\omega$-distributive forcing. In other cases forcing ${\mathbb{P}}(\omega^{\delta})$ collapses (at least) $\omega_{2}$ to $\omega$.
Proof. (A) Let $\delta=\theta+1$. For $\theta=0$ we have $\omega^{\delta}=\omega$ and $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega)={\mathcal{P}}_{\omega}$. If $\theta>0$, then by Lemma 4.4(b) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{\omega}\omega^{\theta}$. Since $\theta\geq 1$ we have $\mathop{\rm cf}\nolimits(\omega^{\theta})\geq\omega$; so, by Theorem 4.7, ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
If $\mathop{\rm cf}\nolimits\delta=\omega$, let $\delta=\sup\{\delta_{i}:i<\omega\}$, where $\langle\delta_{i}:i<\omega\rangle$ is an increasing sequence of non-zero ordinals. Then by Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}=\sum_{i<\omega}\omega^{\delta_{i}}$. For $i<\omega$ we have $\delta_{i}>0$ and, hence, $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$. Thus, by Theorem 4.7, ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
Now we prove that the preorder $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed. By Lemma 4.4 we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\omega$ and $\omega^{\delta}\cong\sum_{i<\omega}\omega^{\delta_{i}}=\sum_{i<\omega}L_{i}=:L$. Thus $\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})\cong\mathop{\rm sm}\nolimits{\mathbb{P}}(L)=\langle P(L)\setminus{\mathcal{I}},\subset_{\mathcal{I}}\rangle$, where :=${\mathcal{I}}:=\{A\subset L:L\not\hookrightarrow A\}$. For $\delta>1$ as in (a) we conclude that $\mathop{\rm cf}\nolimits(\omega^{\delta_{i}})\geq\omega$, for all $i<\omega$, and, by Proposition 4.6(c), for $A\subset L$ we have
$A\in P(L)\setminus{\mathcal{I}}\;\;\Leftrightarrow\;\forall i\in\omega\;\;S^{i}_{A}:=\{j\in\omega:L_{i}\hookrightarrow L_{j}\cap A\}\in[\omega]^{\omega}.$ (19)
For $A_{n}\in P(L)\setminus{\mathcal{I}}$, $n\in\omega$, where $\dots\subset_{\mathcal{I}}A_{1}\subset_{\mathcal{I}}A_{0}$, we construct $A\in P(L)\setminus{\mathcal{I}}$ such that $A\subset_{\mathcal{I}}A_{n}$, for all $n\in\omega$. Since ${\mathcal{I}}$ is an ideal we have $P(L)\setminus{\mathcal{I}}\ni B_{i}:=\bigcap_{n\leq i}A_{n}\subset A_{i}$, so we can assume that $A_{0}\supset A_{1}\supset A_{2}\supset\dots$. By recursion we define the sequences $\langle j_{i}:i\in\omega\rangle$ and $\langle f_{i}:i\in\omega\rangle$ such that for each $i\in\omega$ we have
(i) $j_{i}\in\omega$,
(ii) $j_{i}<j_{i+1}$,
(iii) $f_{i}:L_{i}\hookrightarrow L_{j_{i}}\cap A_{i}$.
Since $A_{0}\in P(L)\setminus{\mathcal{I}}$, by (19) for $i=0$ there are $j_{0}\in\omega$ and $f_{0}:L_{0}\hookrightarrow L_{j_{0}}\cap C_{0}$.
Let $j_{0},\dots,j_{i}$ and $f_{0},\dots,f_{i}$ be sequences satisfying (i)–(iii). Since $A_{i+1}\in P(L)\setminus{\mathcal{I}}$ by (19) for $i+1$ there is $j_{i+1}\in S^{i+1}_{A_{i+1}}\setminus(j_{i}+1)$ so we have $j_{i}<j_{i+1}$ and there is $f_{i+1}:L_{i+1}\hookrightarrow L_{j_{i+1}}\cap A_{i+1}$. The recursion works.
By (ii) $i_{1}<i_{2}$ implies $j_{i_{1}}<j_{i_{2}}$, and, hence, $L_{j_{i_{1}}}<L_{j_{i_{2}}}$, which by (iii) implies $f_{i_{1}}[L_{i_{1}}]<f_{i_{2}}[L_{i_{2}}]$. Thus $f=\bigcup_{i\in\omega}f_{i}:L\hookrightarrow L$ and
$A:=f[L]=\bigcup_{i\in\omega}f_{i}[L_{i}]\in P(L)\setminus{\mathcal{I}}.$ (20)
For $n^{*}\in\omega$ we show that $A\subset_{{\mathcal{I}}}A_{n^{*}}$, that is $A\setminus A_{n^{*}}\in{\mathcal{I}}$. By (iii) we have $A\subset\bigcup_{i\in\omega}L_{j_{i}}\cap A_{i}$ and, since $A_{i}\subset A_{n^{*}}$, for $i\geq n^{*}$, we have $A\setminus A_{n^{*}}\subset\bigcup_{i\in\omega}L_{j_{i}}\cap A_{i}\setminus A_{n^{*}}=\bigcup_{i<n^{*}}L_{j_{i}}\cap A_{i}\setminus A_{n^{*}}\subset\bigcup_{i<n^{*}}L_{j_{i}}$. Thus $S^{0}_{A\setminus A_{n^{*}}}:=\{j\in\omega:L_{0}\hookrightarrow L_{j}\cap A\setminus A_{n^{*}}\}\subset\{j_{i}:i<n^{*}\}$ and by (19) we have $A\setminus A_{n^{*}}\in{\mathcal{I}}$.
(B) By Theorem 4.8 (for $\lambda=\omega$) we have ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$.
By Lemma 4.1, $\theta=\lim_{n\rightarrow\omega}\delta_{n}$, where $\mathop{\rm cf}\nolimits(\delta_{n})=\kappa$, for all $n\in\omega$. Clearly we have $\omega^{\delta}=\omega^{\theta}\omega^{\kappa}=\sup\{\omega^{\delta_{n}}:n\in\omega\}\kappa$. Let ${\mathbb{L}}=\langle L,<\rangle=\sum_{\xi<\kappa}{\mathbb{L}}_{\xi}$, where ${\mathbb{L}}_{\xi}=\langle L_{\xi},<_{\xi}\rangle\cong\omega^{\theta}$, for $\xi<\kappa$, and $L_{\xi}\cap L_{\zeta}=\emptyset$, for $\xi\neq\zeta$. Let =${\mathcal{I}}=\{A\subset L:{\mathbb{L}}\not\hookrightarrow A\}$ and, for $A\subset L$ and $n\in\omega$, let
$S^{n}_{A}=\{\xi<\kappa:\omega^{\delta_{n}}\hookrightarrow L_{\xi}\cap A\}.$
Then by Claim 4.9 (where $\lambda=\omega$) for $A,B\subset L$ we have:
$A\in P(L)\setminus{\mathcal{I}}$$\Leftrightarrow$ $\forall n\in\omega\;\;|S^{n}_{A}|=\kappa,$ (21)
$A\subset_{{\mathcal{I}}}B$$\Leftrightarrow$ $\exists n\in\omega\;\;|S^{n}_{A\setminus B}|<\kappa.$ (22)
We prove that $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\sigma$-closed. Let $A_{k}\in{\mathbb{P}}(L)$, for $k\in\omega$, where $\dots\subset_{\mathcal{I}}A_{1}\subset_{\mathcal{I}}A_{0}$. W.l.o.g. we can suppose that
$A_{0}\supset A_{1}\supset\dots,$ (23)
because ${\mathcal{I}}$ is an ideal and, hence, $B_{k}:=\bigcap_{i\leq k}A_{k}\in{\mathbb{P}}(L)$ and $B_{0}\supset B_{1}\supset\dots$; so a lower bound for $B_{k}$’s will be a lower bound for $A_{k}$’s.
Also w.l.o.g. we suppose that for each $k\in\omega$ we have $A_{k}\not\subset_{\mathcal{I}}A_{k+1}$, so, by (22), $S^{k}_{A_{k}\setminus A_{k+1}}\in[\kappa]^{\kappa}$. For $\xi\in S^{k}_{A_{k}\setminus A_{k+1}}$ we have $\omega^{\delta_{k}}\hookrightarrow L_{\xi}\cap A_{k}\setminus A_{k+1}$, and we choose sets $A^{k}_{\xi}$ such that
$\omega^{\delta_{k}}\cong A^{k}_{\xi}\subset L_{\xi}\cap A_{k}\setminus A_{k+1},\mbox{ for }\xi\in S^{k}_{A_{k}\setminus A_{k+1}}\mbox{ and }k\in\omega,$ (24)
and define
$A=\bigcup_{k\in\omega}\bigcup_{\xi\in S^{k}_{A_{k}\setminus A_{k+1}}}A^{k}_{\xi}.$ (25)
By (24) and (25), for $\xi\in S^{k}_{A_{k}\setminus A_{k+1}}$ we have $\omega^{\delta_{k}}\hookrightarrow L_{\xi}\cap A$, that is $\xi\in S^{k}_{A}$; so $S^{k}_{A_{k}\setminus A_{k+1}}\subset S^{k}_{A}$ and, hence, $|S^{k}_{A}|=\kappa$, for all $k\in\omega$, which by (21) gives $A\in{\mathbb{P}}(L)$.
We show that $A\subset_{\mathcal{I}}A_{k}$, for $k\in\omega$. By (24) we have $A^{k}_{\xi}\subset A_{k}$, for all $\xi\in S^{k}_{A_{k}\setminus A_{k+1}}$ so, by (23) and (25),
$A\setminus A_{k}\subset\bigcup_{l<k}\bigcup_{\xi\in S^{l}_{A_{l}\setminus A_{l+1}}}A^{l}_{\xi}.$ (26)
Suppose that $\omega^{\delta_{k}}\hookrightarrow L_{\xi_{0}}\cap A\setminus A_{k}$, for some $\xi_{0}<\kappa$. Then by (26) and since $A^{l}_{\xi}\subset L_{\xi}$ we have $\omega^{\delta_{k}}\hookrightarrow\bigcup_{l<k\;\land\;\xi_{0}\in S^{l}_{A_{l}\setminus A_{l+1}}}A^{l}_{\xi_{0}}=:B$. But $B$ is a union of $\leq k$ many summands $A^{l}_{\xi_{0}}\cong\omega^{\delta_{l}}$, where $l\leq k-1$ and, hence, $\mathop{\mbox{type}}\nolimits(B)\leq\omega^{\delta_{k-1}}k<\omega^{\delta_{k-1}}\omega=\omega^{\delta_{k-1}+1}\leq\omega^{\delta_{k}}$ and $\omega^{\delta_{k}}\hookrightarrow B$ is impossible. Thus for all $\xi<\kappa$ we have $\omega^{\delta_{k}}\not\hookrightarrow L_{\xi}\cap A\setminus A_{k}$, that is $S^{k}_{A\setminus A_{k}}=\emptyset$ and, by (22), $A\subset_{\mathcal{I}}A_{k}$.
(C) The statement follows from Theorem 4.8 (for $\lambda>\omega$).
(D) Clearly we have $\mathop{\rm cf}\nolimits(\delta)=\kappa$ so, by Lemma 4.4(a), $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$. In addition we have $\omega^{\delta}=\omega^{\theta+\kappa}=\omega^{\theta}\kappa=\sum_{\kappa}\omega^{\theta}$ and, by Lemma 4.4(a) again, $\mathop{\rm cf}\nolimits(\omega^{\theta})=\mathop{\rm cf}\nolimits(\theta)\geq\kappa$. Thus, by Theorem 4.7, ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$.
(E) By Lemma 4.4(a) we have $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\xi<\kappa}\omega^{\delta_{\xi}}$. By the assumption and Lemma 4.4(a) again, for each $\xi<\kappa$ we have $\mathop{\rm cf}\nolimits(\omega^{\delta_{\xi}})=\mathop{\rm cf}\nolimits(\delta_{\xi})\geq\kappa$; thus, by Theorem 4.7, ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sm}\nolimits{\mathbb{P}}(\omega^{\delta})$. $\Box$
The partial orders ${\mathbb{P}}(\alpha)$. ZFC and consistency results
Theorem 4.11
[…]
(a) If $\delta_{i}$ satisfies (A) or (B), for each $i\leq n$, then the partial order $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is $\sigma$-closed and completely embeds ${{\mathcal{P}}_{\omega}}^{k}$, where $k={\sum_{i=1}^{n}s_{i}}$;
(b) Otherwise, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ completely embeds ${\mathcal{P}}_{\lambda}$, for some uncountable regular $\lambda$ and, hence, collapses at least $\omega_{2}$ to $\omega$. If, in addition, $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})=(2^{|\alpha|})^{+}$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\alpha|})$.
Proof. Let $\langle\delta^{\prime}_{j}:1\leq j\leq k\rangle=\langle\delta_{n},\dots,\delta_{n},\dots,\delta_{1},\dots,\delta_{1}\rangle$, where $\delta_{i}$ appears $s_{i}$ many times, for $i\leq n$. Then $\alpha\cong\sum_{1\leq j\leq k}L_{j}:=L$, where $L_{j}\cong\omega^{\delta^{\prime}_{j}}$, for $j\leq k$, and by Theorem 3.2 we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)\cong\prod_{1\leq j\leq k}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$.
(a) Under the assumptions, by Theorem 4.10 for each $j\leq k$ the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$ is $\sigma$-closed and ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta^{\prime}_{j}})$. By Fact 2.1(b) the poset $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ is $\sigma$-closed and by Fact 2.3 ${{\mathcal{P}}_{\omega}}^{k}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$.
(b) Otherwise, by Lemma 4.1 there are a regular $\lambda\geq\omega_{1}$ and $j\leq k$, such that ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta^{\prime}_{j}}))$ and, since $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta^{\prime}_{j}}))\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ we have ${\mathcal{P}}_{\lambda}\hookrightarrow_{c}\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$. If $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\lambda})=(2^{|\alpha|})^{+}$, then by Fact 2.6(e) forcing by $\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)$ collapses $2^{|\alpha|}$ to $\omega$ and, hence, $|\mathop{\rm sq}\nolimits{\mathbb{P}}(\alpha)|=2^{|\alpha|}$. So, by Fact 2.5, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\alpha|})$. $\Box$
Theorem 4.12
If ${\mathfrak{h}}<{\mathfrak{c}}=\omega_{2}=2^{\omega_{1}}$, then $\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$ and
(a) For each non-zero ordinal $\delta<{\mathfrak{c}}$ we have
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\left\{\begin{array}[]{cl}\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}}),&\mbox{ if (A) or (B) holds or }\delta<\omega_{1},\\[2.84526pt] \mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}}),&\mbox{ if (D) or (E) holds;}\end{array}\right.$
(b) $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$ or $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha)))\cong\mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}})$, for $\alpha\in[\omega,{\mathfrak{c}})$.
Proof. First, by Facts 2.6(b) and 2.5(a) we have $\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$.
(a) For $\delta<\omega_{1}$ by Theorem 1.1(b) we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$. If $\delta\in[\omega_{1},{\mathfrak{c}})$, then regarding Lemma 4.1 Case (C) impossible. If $\delta$ satisfies (A) or (B), then by Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\omega_{1}$-closed and ${\mathcal{P}}_{\omega}\hookrightarrow_{c}\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$, so, by Fact 2.6(b), $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses ${\mathfrak{c}}$ to $\omega_{1}$. Since ${\mathfrak{c}}^{<\omega_{1}}={\mathfrak{c}}$, by Fact 2.5 we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})$. If (D) or (E) holds for $\delta$, then by Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses ${\mathfrak{c}}$ to $\omega$ and, by Fact 2.5, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,{\mathfrak{c}})$.
(b) If $\alpha\in[\omega_{1},{\mathfrak{c}})$ and $\alpha=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}+m$, then by Fact 2.9(c) we have $|\alpha|=|\delta_{n}|=\omega_{1}$ and by Fact 4.3(b) $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))|=|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta_{n}}))|\leq 2^{\omega_{1}}={\mathfrak{c}}$. By Theorem 4.11, either the poset $\mathop{\rm sq}\nolimits({\mathbb{P}}(\alpha))$ is $\sigma$-closed and collapses ${\mathfrak{c}}$ to $\omega_{1}$, or collapses ${\mathfrak{c}}$ to $\omega$; thus, the statement follows from Fact 2.5. $\Box$
E.g., regarding Theorem 4.12(a) we obtain $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$, whenever $\mathop{\rm cf}\nolimits(\delta)\leq\omega$, but also for $\delta=\omega_{1}\omega+\omega_{1}$. We note that it is consistent that ${\mathfrak{c}}=\omega_{2}=2^{\omega_{1}}$ and ${\mathbb{P}}(\omega^{2})$ preserves cardinals (see [10]).
5 The posets ${\mathbb{P}}(\omega^{\delta})$ and collapsing algebras
Our intention is to understand how the posets ${\mathbb{P}}(\alpha)$ look like. By Theorem 3.2 one step in that direction is to regard their separative quotients and to classify the corresponding factors of the form $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))=(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$ up to isomorphism of their Boolean completions. By Theorem 1.1 it is consistent that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},{\mathfrak{c}})\cong\mathop{\rm ro}\nolimits(P(\omega)/\mathop{\rm Fin}\nolimits)$ for each non-zero ordinal $\delta<\omega_{1}$ and Theorem 4.12(a) gives an extension for uncountable ordinals.
Collapse to $\omega$
Here we detect some classes of ordinals $\delta$ such that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$. More generally, we have
Fact 5.1
If $\lambda\geq\omega$ is a regular cardinal, $\delta\geq\omega_{1}$ an ordinal, $2^{|\delta|}>\lambda$, and the forcing $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\lambda$-closed and collapses $2^{|\delta|}$ to $\lambda$, then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$.
Proof. Defining ${\mathbb{B}}:=P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$, by Fact 4.3 we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathbb{B}}^{+}$ and $|{\mathbb{B}}^{+}|\leq 2^{|\delta|}$. Since ${\mathbb{B}}^{+}$ collapses $2^{|\delta|}>\lambda$ to $\lambda$ we have $|{\mathbb{B}}^{+}|=2^{|\delta|}$ and $(2^{|\delta|})^{<\lambda}=2^{|\delta|}$; so, by Fact 2.5, $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$. Clearly, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathbb{B}}^{+}$ implies that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))=\mathop{\rm ro}\nolimits({\mathbb{B}})$. $\Box$
Theorem 5.2
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$, if the ordinal $\delta$ satisfies (D) or (E), $2^{\mathop{\rm cf}\nolimits(\delta)}=2^{|\delta|}$ and ($2^{<\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)$ or $2^{\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)^{+}$).
Proof. By Theorem 4.10 we have $\kappa:=\mathop{\rm cf}\nolimits(\delta)>\omega$ and ${\mathcal{P}}_{\kappa}\hookrightarrow_{c}\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$. If $2^{<\kappa}=\kappa$, then $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}$ (since the tree ${}^{<\kappa}2$ has $2^{\kappa}$ almost disjoint branches); by Fact 2.6(a) the equality $2^{\kappa}=\kappa^{+}$ implies $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}$ as well. Thus, by our assumption, $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}=(2^{|\delta|})^{+}$ and by Fact 2.6(e) the poset ${\mathcal{P}}_{\kappa}^{+}$ and, hence, $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$, collapses $2^{|\delta|}$ to $\omega$. So, by Fact 5.1 $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})$. $\Box$
Example 5.3
We regard Theorem 5.2 in the particular case when $|\delta|=\mathop{\rm cf}\nolimits(\delta)=:\kappa>\omega$ and $2^{\mathop{\rm cf}\nolimits(\delta)}=\mathop{\rm cf}\nolimits(\delta)^{+}$. It is easy to check (using the Cantor normal form (4) for $\delta$) that if (D) holds, then $\delta$ satisfies one of the following conditions:
  • (i)
    $\delta=\kappa$
  • (ii)
    $\delta=\theta+\kappa(\gamma+m)$,
    where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma<\kappa$ and $m\geq 2$ and ($\theta=0$ or $\kappa^{2}\leq\theta<\kappa^{+}$);
  • (iii)
    $\delta=\theta+\kappa^{\gamma_{2}+m_{2}}(\gamma_{1}+m_{1})+\kappa$,
    where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma_{1}<\kappa$ and $\mathop{\rm Lim}\nolimits_{0}\ni\gamma_{2}<\kappa^{+}$ and $m_{1},m_{2}\in{\mathbb{N}}$ and ($\theta=0$ or $\kappa^{\gamma_{2}+m_{2}+1}\leq\theta<\kappa^{+}$);
  • (iv)
    $\delta=\theta+\kappa^{\gamma_{2}}(\gamma_{1}+m_{1})+\kappa$,
    where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma_{1}<\kappa$ and $m_{1}\in{\mathbb{N}}$ and $\mathop{\rm cf}\nolimits(\gamma_{2})=\kappa\leq\gamma_{2}<\kappa^{+}$ and ($\theta=0$ or $\kappa^{\gamma_{2}+1}\leq\theta<\kappa^{+}$).
On the other hand, if (E) holds, then $\delta$ satisfies one of the following conditions:
  • (i)
    $\delta=\theta+\kappa^{\gamma+m}$,
    where $\mathop{\rm Lim}\nolimits_{0}\ni\gamma<\kappa^{+}$ and $m\in{\mathbb{N}}$ and $\gamma+m\leq 2$ and ($\theta=0$ or $\kappa^{\gamma+m}\leq\theta<\kappa^{+}$);
  • (ii)
    $\delta=\theta+\kappa^{\gamma}$,
    where $\kappa=\mathop{\rm cf}\nolimits(\gamma)\leq\gamma<\kappa^{+}$ and ($\theta=0$ or $\kappa^{\gamma}\leq\theta<\kappa^{+}$).
If $\delta=\kappa>\omega$ is a regular cardinal, then $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))={\mathcal{P}}_{\kappa}$. If $2^{<\kappa}=\kappa$ or $2^{\kappa}=\kappa^{+}$, then by Theorem 5.2 $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$.
Example 5.4
It is consistent that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$, for a regular cardinal $\kappa>\omega$. Baumgartner [2] constructed a model of ZFC in which $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}<2^{\omega_{1}}$. Thus ${\mathcal{P}}_{\omega_{1}}$ collapses $\omega_{2}$ to $\omega$ and the cardinals $>\omega_{2}$ are preserved in extensions by ${\mathcal{P}}_{\omega_{1}}$. So, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\omega_{1}})))\not\cong\mathop{\rm Col}\nolimits(\omega,2^{\omega_{1}})$. In fact, if $\kappa>\omega$ is a regular cardinal, then by Fact 2.6(e), $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$ iff $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})=(2^{\kappa})^{+}$.
Theorem 5.5
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$, if $\kappa$ is a cardinal, $\kappa>2^{\mathop{\rm cf}\nolimits(\kappa)}>\mathop{\rm cf}\nolimits(\kappa)>\omega$ and $2^{\kappa}=\kappa^{+}$.
Proof. Since $\omega^{\kappa}=\kappa$ we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))={\mathcal{P}}_{\kappa}^{+}$ and by Fact 2.6(d) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa}))$ collapses $2^{\kappa}=\kappa^{+}$ to $\omega$. So, by Fact 5.1, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\kappa})))\cong\mathop{\rm Col}\nolimits(\omega,2^{\kappa})$. $\Box$
Collapse to $\omega_{1}$
Here we are looking for the ordinals $\delta$ such that $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. First we show that if ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$ (examples of such $\delta$’s can be found in the previous paragraph), then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$, for all $n\in{\mathbb{N}}$, namely, the Boolean completion of ${\mathbb{P}}(\omega^{\theta})$ is a “constant function” on the interval $(\delta,\delta+\omega)$. We recall some facts from [10] and [13].
Fact 5.6
Let $\delta$ denote an ordinal, $\kappa$ a cardinal and ${\mathbb{B}}\,$ a Boolean algebra.
(a) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n}))\cong(\mathop{\rm rp}\nolimits^{n}(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}))^{+}$, for each $n\in{\mathbb{N}}$, whenever $\delta\geq 1$.
(b) $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega,\kappa)))\cong\mathop{\rm Col}\nolimits(\omega_{1},\kappa)$, for each $n\in{\mathbb{N}}$, whenever $\kappa^{\omega}=\kappa\geq{\mathfrak{c}}$.
(c) $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega_{1},\kappa)))\cong\mathop{\rm Col}\nolimits(\omega_{1},\kappa^{\omega})$, for each $n\in{\mathbb{N}}$, whenever $\kappa\geq 2$.
(d) $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm ro}\nolimits({\mathbb{B}})))\cong\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))$, for each $n\in{\mathbb{N}}$.
Proof. Statement (a) is a straightforward generalization of Lemma 3.2(d) from [10] (proved there for countable $\delta$). Statements (b), (c) and (d) follow from Theorems 6.9 and 5.4 and Fact 2.6 from [13]. $\Box$
Theorem 5.7
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$, for all $n\in{\mathbb{N}}$, whenever $\delta\geq\omega_{1}$ and (a) or (b) holds, where
(a) ${\mathbb{P}}(\omega^{\delta})$ collapses $2^{|\delta|}$ to $\omega$,
(b) $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))$ is $\sigma$-closed and collapses $2^{|\delta|}$ to $\omega_{1}$.
Proof. Defining $\lambda=\omega$, if (a) holds, and $\lambda=\omega_{1}$, if (b) holds, by Fact 5.1 we have $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$. Defining ${\mathbb{B}}:=P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}}$, by Fact 4.3 we have $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))={\mathbb{B}}^{+}$ and, hence, $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta})))=\mathop{\rm ro}\nolimits({\mathbb{B}})$; thus, $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$. By Fact 5.6(a) we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta+n})\cong(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))^{+}$; so,
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta+n}))$$\cong$ $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}({\mathbb{B}}))\quad\mbox{ (since $\mathop{\rm rp}\nolimits^{n}({\mathbb{B}})$ is a Boolean algebra)}$)
[…]
$\cong$ $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})))\quad\mbox{ (since $\mathop{\rm ro}\nolimits({\mathbb{B}})\cong\mathop{\rm Col}\nolimits(\lambda,2^{|\delta|})$)}.$
Now, if (a) holds, then $\lambda=\omega$ and by Fact 5.6(b) we have $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega,2^{|\delta|})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. If (b) holds, then $\lambda=\omega_{1}$ and, by Fact 5.6(c), $\mathop{\rm ro}\nolimits(\mathop{\rm rp}\nolimits^{n}(\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{|\delta|})$. $\Box$
Example 5.8
It is consistent that ${\mathbb{P}}(\omega^{\kappa+1})$ collapses $2^{\kappa}$ to $\omega_{1}$, although $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})<(2^{\kappa})^{+}$ and ${\mathbb{P}}(\omega^{\kappa})$ does not collapse $2^{\kappa}$ to $\omega$. First we note that the equality ${\mathfrak{h}}=\omega_{1}$ holds in Cohen extensions of models of GCH (see [3]). If $V^{\prime}$ is a Cohen model for ${\mathfrak{c}}=\omega_{5}$ obtained from a model $V$ of GCH, then in $V^{\prime}$ we have ${\mathfrak{h}}=\omega_{1}$ and, by Fact 2.4, $2^{\omega_{1}}=(\max(\omega_{5},2)^{\omega_{1}})^{V}=(\omega_{5}^{\omega_{1}})^{V}=\omega_{5}$. By Theorem 4.10(A) and Fact 2.6(b) we have $1_{{\mathbb{P}}(\omega^{\omega_{1}+1})}\Vdash|2^{\omega_{1}}|=\omega_{1}$ although (see [2]) $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}<\omega_{6}=(2^{\omega_{1}})^{+}$ and forcing by ${\mathbb{P}}(\omega^{\omega_{1}})$ does not collapse $2^{\omega_{1}}$ to $\omega$.
Theorem 5.9
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\mu})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$, if $\mu$ is a cardinal, $\mu>\mathop{\rm cf}\nolimits(\mu)=\omega$ and some of the following conditions hold
(a) $2^{\mu}=\mu^{+}$,
(b) $2^{<\mu}=\mu$,
(c) $\mathrm{MA}+\mu<{\mathfrak{c}}$,
(d) $V$ is a Cohen model for ${\mathfrak{c}}=\kappa$ obtained from a model $V^{\prime}$ of GCH in which $\kappa$ is regular and $\mu<\kappa$.
Then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\mu+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$, for all $n\in{\mathbb{N}}$.
Proof. By Fact 2.9 we have $\omega^{\mu}=\mu$; thus ${\mathbb{P}}(\omega^{\mu})={\mathbb{P}}(\mu)=[\mu]^{\mu}$ and by Fact 2.1(d) we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\mu})={\mathcal{P}}_{\mu}$. By Theorem 4.10 the poset ${\mathcal{P}}_{\mu}$ is $\omega_{1}$-closed and by Fact 2.6(c) collapses $\mu^{\omega}$ to $\omega_{1}$. So, by Fact 5.1 it is sufficient to prove that each of conditions implies that $\mu^{\omega}=2^{\mu}$.
If $2^{\mu}=\mu^{+}$, then, since by König’s theorem $\mu^{+}\leq\mu^{\omega}\leq 2^{\mu}$, we have $\mu^{\omega}=2^{\mu}$. If $2^{<\mu}=\mu$, then, since for a singular cardinal $\mu$ we have $2^{\mu}\in\{2^{<\mu}\mu^{\mathop{\rm cf}\nolimits(\mu)},(2^{<\mu})^{\mathop{\rm cf}\nolimits(2^{<\mu})}\}$, we obtain $\mu^{\omega}=2^{\mu}$. If $V\models\mathrm{MA}+\mu<{\mathfrak{c}}$, then ${\mathfrak{c}}\leq\mu^{\omega}\leq 2^{\mu}={\mathfrak{c}}$ and we have $\mu^{\omega}=2^{\mu}$. Finally, if $V=V^{\prime}_{\mathop{\mathrm{Fn}}\nolimits(\kappa,2)}[G]$, then by Fact 2.4 in $V$ we have $2^{\mu}=(\max(\kappa,2)^{\mu})^{V^{\prime}}=(\kappa^{\mu})^{V^{\prime}}=\kappa$ and $\mu^{\omega}=(\max(\kappa,\mu)^{\omega})^{V^{\prime}}=(\kappa^{\omega})^{V^{\prime}}=\kappa$; so, $\mu^{\omega}=2^{\mu}$ again. The second statement follows from the first and Theorem 5.7(b). $\Box$
Acknowledgement
This research was supported by the Science Fund of the Republic of Serbia, Program IDEAS, Grant No. 7750027: Set-theoretic, model-theoretic and Ramsey-theoretic phenomena in mathematical structures: similarity and diversity–SMART.